### v1

# A Theory of Recommendations

[…]
[linecolor=black,backgroundcolor=gray!10,linewidth=1pt]mybox
A Theory of Recommendations
Jean-Michel Benkert and Armin Schmutzler Benkert: Department of Economics, University of Bern, jean-michel.benkert@unibe.ch. Schmutzler: Department of Economics, University of Zurich, and CEPR, armin.schmutzler@econ.uzh.ch. We are grateful for very helpful discussions with Piotr Dworczak, Igor Letina, Marc Möller, Nick Netzer, Volker Nocke, Egor Starkov, Jakub Steiner and seminar participants in Bern, Fribourg, Zurich, at SAET (2023), Swiss Society of Economics of Statistics (2024), and Swiss IO Day (2024).
(August 21, 2024)
This paper investigates the value of recommendations for disseminating economic information, with a focus on frictions resulting from preference heterogeneity. We consider Bayesian expected-payoff maximizers who receive non-strategic recommendations by other consumers. The paper provides conditions under which different consumer types accept these recommendations. Moreover, we assess the overall value of a recommendation system and the determinants of that value. Our analysis highlights the importance of disentangling objective information from subjective preferences when designing value-maximizing recommendation systems.
Keywords: recommendations, preference heterogeneity, optimal design
JEL Classification: D02, D47, D83
1 Introduction
Recommendations play an essential role for the diffusion of economic information. Job applicants frequently rely on assessments of potential employers by their friends. Conversely, job candidates often have to provide letters of recommendation from former employers. Consumers may base their choices of travel destinations, restaurants, or doctors on what they have heard from their acquaintances. In the digital economy, product ratings, typically reflecting anonymous consumer reports, are omnipresent.
Despite their relevance, recommendations are unlikely to function smoothly. For instance, too few recommendations may be given (Kremer et al., 2014; Che and Hörner, 2018), recommenders may be positively selected (Acemoglu et al., 2022), and interested parties might interfere with the process by choosing which recommendations to publish (Bolton et al., 2013; Tadelis, 2016). In the literature, it is common to make the simplifying assumption that product qualities can be objectively ranked. [^1] See also Che and Hörner (2018) and Hui et al. (2023). Instead, our paper assumes that preferences are heterogeneous, in which case product qualities cannot be objectively ranked and recommendations may be hard to interpret, even without strategic biases. We ask whether and how recommendations can be useful in spite of preference heterogeneity. Even well-intended recommendations by non-strategic players may reflect the senders’ preferences rather than the objective truth. A receiver who does not question the informational content of a recommendation may therefore make biased decisions.
To motivate our approach, consider a consumer’s choice between medical doctors. This is a complex problem, as the perceived quality of medical care will typically depend on many aspects. How carefully does a doctor listen to the patient? Is the technical equipment up to date? Is the staff friendly? How crowded is the waiting room? Moreover, different people may weigh different aspects differently. Despite such preference heterogeneity, rather than randomly choosing a doctor, people often base their decision on recommendations. Importantly, senders have limited stakes in the process and therefore lack incentives to think through the effects of their recommendation carefully. By contrast, adequately using the information can be crucial for the receiver. Finally, even if the sender is willing to think through the recommendation carefully, it will typically not be possible to convey all relevant information. For all these reasons, it is not obvious why or when recommendation can be useful.
This example illustrates three key features of our theory of recommendations:
Heterogeneity of Products and Preferences: Products come in different versions and consumers value them differently.
Asymmetry of Stakes: Making sure that the recommendation is not misunderstood is more important for the receiver than for the sender.
Coarseness of Recommendations: The sender is unable to provide enough detail for all relevant aspects of the service to become fully transparent to the receiver.
These features will be even more salient for ratings provided on internet platforms, as such ratings usually do not come from close acquaintances. Hence, the receivers should consider that senders have preferences which differ from their own and are likely to give an incomplete account of their consumption experience without exerting much effort. Backing out the implications of the recommendation for the receiver’s own decision is therefore a complex task. Relevant considerations for the interpretation are these: To what extent does the recommendation contain objective truth? And when it does not, how likely is it that the preferences of the sender and receiver are aligned? The answers to these questions will determine whether it is in the receiver’s interest to follow the recommendation and how valuable it is.
Our model captures the above three features of recommendations in the simplest possible way. An agent must decide between several options, referred to as products for ease of exposition. Products are two-dimensional objects, differing only in whether they have high or low quality in each dimension. Thus, four product versions exist. All agents agree that a product is worth choosing if it has high quality in both dimensions and is thus (objectively) good. Conversely, they agree that a product is not worth choosing if it has low quality in both dimensions and is thus (objectively) bad. Accordingly, we call the products uncontroversial in both of these cases. By contrast, we refer to products as controversial when they have high quality in one dimension and low quality in the other. An agent prefers a controversial product that has high quality in the dimension that is more relevant for them, and is thus subjectively good, to a product that has high quality only in the less-relevant dimension (thus subjectively bad). Moreover, even agents who agree about which controversial product they prefer may differ in the intensity of these preferences. This heterogeneity is captured in a continuous type distribution.
We assume that products are experience goods so that each agent only has stochastic knowledge of the quality distribution before consumption: They know the probability of each product version. The agent obtains a recommendation from a previous consumer of one of the products. The recommendation reveals whether the sender had a consumption experience that was sufficiently good (a buy recommendation), with payoffs above an exogenous threshold level $R$, or not (a don’t-buy recommendation). Independent of $R$, the sender (she) always gives a buy recommendation for a good product and a don’t-buy recommendation for a bad product. In addition, she sends a buy recommendation for a controversial product if her preferences for the product are strong enough. In line with her low stakes in the issue, the sender thus only provides a coarse description of her own consumption experience. By contrast, the receiver (he) of the recommendation, with his high stakes in the issue, will carefully evaluate the informational content of the recommendation using Bayesian inference, taking into account the distribution of product qualities, his own preferences, and those of the population.
In this framework, we address a variety of questions: How do recommendations create value for receivers, how can we interpret this value, and when is it maximal? If one could influence the payoff threshold beyond which a sender gives a buy recommendation, how high should it be? Are recommendations valuable even when uncontroversial products are unlikely? How does the value depend on how polarized the population is?
To answer these questions, we need to derive intermediate results of independent interest. Most importantly, we investigate under which conditions receivers accept a recommendation. We find that for a Bayesian expected-payoff maximizer, two key quantities measuring the objective and subjective content of the recommendation are decisive for receiver behavior. These quantities are simple functions of the model primitives. Intuitively, the objective informational content of a recommendation is high if it makes the product sufficiently more likely to be good and less likely to be bad. The subjective content captures the effect of the recommendation on the relative probability of the two controversial product versions. The evaluation of the subjective content thus depends on the receiver’s type. When the recommendation bears little or no subjective content, all types accept it, thus buying after a buy recommendation and choosing the outside option (one of the other alternatives) after a don’t-buy recommendation. As the subjective content increases, revealing that one of the two controversial product versions becomes more likely, only receivers without strong preferences for the other version accept the recommendation. When the recommendation threshold $R$ approaches the maximal payoff, a buy recommendation essentially becomes objective, and similarly for a don’t-buy recommendation when $R$ approaches the minimal payoff. As receivers are aware of this, they accept the recommendation in each of these polar cases since the subjective component vanishes.
Based on this analysis, we then study the value of recommendation systems. We think of this value as the expected-payoff increase of a randomly chosen receiver resulting from a recommendation given by a randomly chosen sender. When the recommendation is objective, or the subjective content is small, the analysis is comparatively simple: As all receivers accept the recommendation, they choose the outside option after a don’t-buy recommendation, just as if they had not received any recommendation at all. Therefore, only buy recommendations create value—they do so by inducing the receiver to purchase the product if it generates a higher expected payoff for him than the outside option. When the recommendation contains more subjective content, only a subset of the receivers accept it, rendering the analysis more complex. The value of the recommendation system then has two sources. First, receivers who accept buy recommendations benefit because they expect the recommended product to be better than the outside option. Second, receivers who do not accept the recommendation benefit from buying the product in spite of a don’t-buy recommendation: These receivers have preferences that are not well aligned with the general population so that the don’t-buy recommendation provides good news about the chances of obtaining a subjectively good product rather than one they like less.
Next, we turn to the design of recommendation systems. Motivated by the practice of digital platforms, we consider the recommendation threshold $R$ as a design variable. When the type distribution is symmetric, so that there is no preference bias on average, the value of the recommendation system is a monotone function of the threshold. If the value is increasing, the product optimally receives a buy recommendation only if it is good; if the value is decreasing, the product optimally receives a don’t-buy recommendation only if it is bad. The former case arises if good products are more likely ex ante than bad products are, and conversely for the latter case. Either way, the content of the recommendation essentially becomes objective. In contrast, with an asymmetrically distributed population, value maximization does not necessarily require an extreme threshold. Rather than revealing only whether a product is good or bad, it may be optimal for the sender to give a buy recommendation even if the product is controversial but sufficiently aligned with her preferences. Intuitively, giving up the exclusive focus on objective content makes sense when both controversial product versions are common but there is broad agreement on which of them is preferable. By contrast, extreme recommendation thresholds that only reveal objective content are still desirable if the product is much more likely to be objectively good than objectively bad or vice versa.
We then discuss further determinants of the value of the recommendation system, assuming that the population is symmetric. First, we find that an increase in the probability of controversial products lowers the value only if the probabilities of the two controversial versions are similar. Second, whether a mean-preserving spread in the population distribution (corresponding to a more polarized population) increases the value of a given recommendation system depends on how strict the recommendation threshold is and on how common good products are relative to bad ones.
Finally, we show the robustness of our results. With mild adjustments, the main insights continue to hold when sender and receiver distributions are distinct—a situation that could arise because, for instance, recommendations stem from early adopters who do not represent the population at large. Second, we allow for more than two types of recommendations. Again, extreme recommendation thresholds increase the probability that a recommendation has objective content and thus increase the value of the system when the population is symmetrically distributed. Third, we show that having access to many recommendations does not necessarily change receiver behavior compared to the case of a single recommendation with an optimal threshold. Then, having access to multiple recommendations does not increase the value.
Section 2 relates the paper to the literature. Section 3 introduces the model and discusses its key assumptions. In Section 4, we describe the conditions under which different receiver types accept recommendations. Then, in Section 5 we first characterize the value of a recommendation system before investigating the determinants of the value and the design of value-maximizing systems. Section 6 discusses our results and considers extensions. Section 7 concludes. All proofs are relegated to the appendix.
2 Relation to the Literature
Our paper investigates frictions in recommendation systems arising from preference heterogeneity, emphasizing the behavior of receivers. By contrast, the literature has mainly focused on frictions on the sender side.
Public Goods Problems
Research on recommendation systems often deals with the exploration-exploitation dilemma and its implications for providing recommendations, interpreted as the provision of a public good. Several authors argue that consumers usually neglect the positive externalities generated by their reviews, leading to insufficient exploration and reviewing, in particular for lesser-known niche products (e.g., Kremer et al. (2014) and Che and Hörner (2018)). Expanding on these papers, Vellodi (2022) considers the impact on market structure, showing how unbiased reviews might disadvantage entrants due to the “cold start” problem, thereby reinforcing the monopoly position of incumbents. These studies view the design of the recommendation system as the solution of a public goods problem by appropriate information management. [^2] For example, Kremer et al. (2014) and Che and Hörner (2018) suggest reducing transparency or endorsing yet-to-be-proven products, while Vellodi (2022) supports the idea of withholding reviews from incumbents to encourage new participants.
Sender Selection
The selection of recommenders is another source of frictions in recommendation systems. In their analysis of the learning dynamics resulting from recommendations, Acemoglu et al. (2022) argue that consumers with strong initial preferences for a product are more likely to purchase it. [^3] In contrast to our paper, their question is essentially whether and how fast consumers find out if the product is good or not, an assessment with which all consumers agree. Beyond that, we model preference heterogeneity differently. Recommendations thus inevitably carry an inherent bias reflecting the predispositions of the recommender pool. [^4] In Section 6.1, we allow for distinct sender and receiver populations and thus for the possibility of a bias among the recommenders relative to the overall population. Receivers should recognize this bias rather than accepting recommendations at face value. Our paper intersects with theirs by highlighting the significance of preference heterogeneity for interpreting recommendations. However, contrary to Acemoglu et al. (2022), we focus on the determinants of the value of recommendation systems and the design of those systems rather than the possibility and speed of learning. [^5] In Section 6.3, we introduce multiple recommendations, as do Acemoglu et al. (2022), but, in contrast to them, we ask under which conditions multiple recommendations enhance system value.
Non-truthful Recommendations and Strategic Manipulations
Strategic actions of recommenders, firms, and platforms may further reduce the value of recommendations for consumers. Bolton et al. (2013) have shown that eBay users often refrain from negative feedback to avoid retaliation. [^6] Tadelis (2016) discusses the design implications of these issues—using the example of eBay deciding to refrain from allowing seller evaluations of buyers—and also touches on strategic motives in recommendations more generally. Like us, Chakraborty and Harbaugh (2010) assume that the sender possesses multidimensional information. They show how this helps her influence the behavior of the receiver in spite of a common bias. Instead, we emphasize the effects of preference heterogeneity without strategic behavior, assuming that senders mechanically provide truthful recommendations. In contrast with our model with consumer recommendations, in Peitz and Sobolev (forthcoming) and Johnen and Ng (2024) firms engage in sophisticated strategies to inflate their ratings. [^7] Peitz and Sobolev (forthcoming) investigate firm-driven recommendations and the choice of sales channel: targeted (with recommendation) or non-targeted. They identify motives behind “inflated recommendations,” where the firm also recommends welfare-reducing bad product matches. Johnen and Ng (2024) analyze a related phenomenon, where firms initially lower prices to attract positive ratings, later leveraging these ratings to justify price increases—effectively reducing the reliability of ratings as a quality indicator. More closely related to our paper, Bourreau and Gaudin (2022) consider incentives for firms to manipulate recommendations in a setting with heterogeneous preferences.
3 Model
We begin by introducing the assumptions of our benchmark model in Section 3.1. Then, in Section 3.2, we discuss and motivate these assumptions.
3.1 Assumptions
Products and Payoffs
We consider a consumer’s choice of a product from a large set of alternatives. We assume that each product is fully characterized by a two-dimensional quality vector, where $Q_{d}=1$ and $Q_{d}=0$ capture high and low quality, respectively, in dimension $d\in\{1,2\}$. Accordingly, four different product versions exist. We assume that the products are experience goods. Thus, the consumer cannot observe product quality before consumption, so the alternatives are identical for him ex ante. However, he knows the prior distribution of the quality vectors, referred to as quality distribution for brevity. Consumers have heterogeneous preferences fully characterized by their type $i\in[-1/2,1/2]$. A consumer of type $i$ receives a payoff of
$v(Q_{1},Q_{2},i)=(1/2+i)Q_{1}+(1/2-i)Q_{2}$
from consuming $(Q_{1},Q_{2})$. [^8] For convenience, we identify a product version with the corresponding quality vector. Thus, possible payoffs range from 0 to 1, with all consumers obtaining the highest payoff, $1$, from consuming version $(1,1)$ and the lowest payoff, 00, from consuming $(0,0)$. Therefore, consumers agree that product versions $(1,1)$ and $(0,0)$ are (objectively) good and bad, respectively. By contrast, the payoff from consuming $(1,0)$ or $(0,1)$ depends on a consumer’s type $i$ so that the assessment of $(1,0)$ and $(0,1)$ is subjective. We thus refer to products with these quality vectors as controversial. Consumers of type $i>0$ prefer $(1,0)$ to $(0,1)$, and vice versa for $i<0$. Those of type $i=0$ are indifferent between these two versions.
The distribution of consumer types is denoted by $F$. For now, we leave the distribution unspecified but for simplicity assume continuity and full support. Finally, the quality distribution is given as $\mathbf{q}=(q_{H},q_{1},q_{2},q_{L})$ where
$\Pr[(Q_{1},Q_{2})=(1,1)]$ $=q_{H},$ $\Pr[(Q_{1},Q_{2})=(1,0)]$ $=q_{1},$
$\Pr[(Q_{1},Q_{2})=(0,0)]$ $=q_{L},$ $\Pr[(Q_{1},Q_{2})=(0,1)]$ $=q_{2}.$
We assume that $\mathbf{q}>0$ unless mentioned otherwise.
Recommendations and Updating
We assume that one of the products comes with a recommendation. The receiver has to choose between this product and an alternative without a recommendation. We distinguish between buy recommendations ($r=B$) and don’t-buy recommendations ($r=D$). Moreover, $r=0$ corresponds to the case without recommendation. In the benchmark model, we assume that the senders (she) and receivers (he) of recommendations are both randomly drawn from $F$.
To eliminate any scope for strategic behavior of the sender, we assume that she mechanically gives a buy recommendation if she obtained a payoff of at least $R\in(0,1)$ and a don’t-buy recommendation otherwise. [^9] We exclude the threshold $R=0$ as this would render the recommendation uninformative. For symmetry, we have also excluded $R=1$, but all results would essentially go through if we allowed $R=1$. As good products (version $(1,1)$) yield a payoff of 1, they always result in a buy recommendation, whereas bad products (version $(0,0)$) always yield a don’t-buy recommendation. The sender’s type determines the recommendation for a controversial product: After having consumed $(1,0)$, a sender of type $i$ gives a buy recommendation if and only if $i\geq R-1/2$, whereas only senders with $i\leq 1/2-R$ give a buy recommendation after consuming $(0,1)$. All told, a recommendation can reflect objective and subjective considerations.
Figure 1: Illustration of Sender Behavior
1/212-1/2- 1 / 2001/4141/41 / 41/2121/21 / 2Product version $(1,0)$1/212-1/2- 1 / 21/414-1/4- 1 / 4001/2121/21 / 2Product version $(0,1)$
We assume that receivers are Bayesian, updating beliefs about product quality after having obtained a recommendation while taking into account the uncertainty about its informational content. We define $\pi^{B}(R)$ and $\pi^{D}(R)$ as the probability that a randomly chosen sender gives a buy or don’t-buy recommendation, respectively. Clearly,
$\pi^{B}(R)$ $:=q_{H}+q_{1}\phi_{1}(R)+q_{2}\phi_{2}(R),$ (1)
$\pi^{D}(R)$ $:=q_{1}(1-\phi_{1}(R))+q_{2}(1-\phi_{2}(R))+q_{L},$ (2)
where $\pi^{D}(R)=1-\pi^{B}(R)$, and $\phi_{1}(R):=1-F(R-1/2)$ and $\phi_{2}(R):=F(1/2-R)$ denote the probability of a buy recommendation conditional on the product version being $(1,0)$ or $(0,1)$, respectively. In words, the probability of receiving a buy recommendation from a randomly chosen sender is given by the sum of the probability that a product is objectively good and the probability that it is subjectively good enough from a randomly chosen sender’s perspective. Thus, following a buy recommendation the receiver’s posterior $\mathbf{p}^{B}(R):=(p_{H}^{B}(R),p_{1}^{B}(R),p_{2}^{B}(R),p_{L}^{B}(R))$ becomes
$p_{H}^{B}(R):=\frac{q_{H}}{q_{H}+q_{1}\phi_{1}(R)+q_{2}\phi_{2}(R)},\quad\quad$ $p_{2}^{B}(R):=\frac{q_{2}\phi_{2}(R)}{q_{H}+q_{1}\phi_{1}(R)+q_{2}\phi_{2}(R)},$
$p_{1}^{B}(R):=\frac{q_{1}\phi_{1}(R)}{q_{H}+q_{1}\phi_{1}(R)+q_{2}\phi_{2}(R)},\quad\quad$ $p_{L}^{B}(R):=0.$
A don’t-buy recommendation gives rise to $\mathbf{p}^{D}(R):=(p_{H}^{D}(R),p_{1}^{D}(R),p_{2}^{D}(R),p_{L}^{D}(R))$, where
$p_{H}^{D}(R):=0,\,$ $p_{2}^{D}(R):=\frac{q_{2}(1-\phi_{2}(R))}{q_{1}(1-\phi_{1}(R))+q_{2}(1-\phi_{2}(R))+q_{L}},$
$p_{1}^{D}(R):=\frac{q_{1}(1-\phi_{1}(R))}{q_{1}(1-\phi_{1}(R))+q_{2}(1-\phi_{2}(R))+q_{L}},\,$ $p_{L}^{D}(R):=\frac{q_{L}}{q_{1}(1-\phi_{1}(R))+q_{2}(1-\phi_{2}(R)))+q_{L}}.$
In the above expressions $\pi^{B}(R)$, $p^{B}(R)$, etc., we will typically drop the dependence on $R$ to simplify notation. Finally, we use the following summary terminology:
Definition 1.
$\mathcal{E}:=(\textbf{q},F)$ defines a decision environment. A recommendation system $\mathcal{R}$ consists of a decision environment $\mathcal{E}$ and a recommendation threshold $R\in(0,1)$.
3.2 Discussion of the Framework
Our theory of recommendations incorporates three key features: heterogeneity of products and preferences, coarseness of recommendations, and asymmetry of stakes between senders and receivers. We now discuss how our model addresses these and further points.
Heterogeneity of Products and Preferences
Without controversial products or preference heterogeneity, the analysis would become trivial. The case without controversial products would correspond to $q_{1}=q_{2}=0$. Then, for arbitrary $R\in(0,1),$ buy recommendations would reveal that the product is good, while don’t-buy recommendations would reveal that it is bad. Thus, there would be no friction in the recommendation system and no role for design considerations. With controversial products but absent preference heterogeneity, all consumers would have the same type $i^{*}$. Assuming w.l.o.g. that $i^{*}\geq 0$, senders would send a buy recommendation (i) only after consuming a good product if $R>\frac{1}{2}+i^{*}$, (ii) after consuming a good product or a controversial product of version $(1,0)$ if $\frac{1}{2}+i^{*}\geq R>\frac{1}{2}-i^{*}$, and (iii) after consuming the good product or any of the controversial products if $\frac{1}{2}-i^{*}\geq R$. The first-best could be induced by the appropriate choice of the recommendation threshold, taking into account how the payoffs from controversial products compare to the expected payoff from the outside option, In sum, a meaningful analysis requires controversial products as well as preference heterogeneity.
Coarse Information Transmission
With a single threshold above and below which senders give buy and don’t-buy recommendations, respectively, it is impossible to always reveal the product version. Thus, as desired, recommendations provide coarse information. If, in contrast, the sender could give four different ratings depending on the consumption experience, full revelation would be possible in our parsimonious model with only four product versions. However, in practice the greater complexity of most products would typically render such full revelation infeasible.
Asymmetry of Stakes
To capture the sender’s limited stakes in the process, we assume that she behaves truthfully, mechanically reporting whether her consumption experience was good enough or not. More sophisticated and potentially strategic recommendations seem at odds with the lack of incentives to invest in this reporting activity, which results in the underprovision of recommendations (see the literature review).
Recommendation Threshold
Our preferred way of thinking about the recommendation threshold is as a design parameter. To motivate this, we rely on Decker (2022) who argues that Airbnb advises its customers to give a five-star rating to every host that provides reasonably good services, thus reserving lower ratings exclusively to very bad hosts. [^10] In our setting, this approach corresponds to the use of a coarse recommendation system, with the vast majority of hosts receiving a buy recommendation and a small subset receiving a don’t-buy recommendation. The recommendation threshold would then be close to $R=0$ so that products with a don’t-buy recommendation are typically objectively bad. In this spirit, we think of the threshold as reflecting the guidance of a platform such as Airbnb or as a social norm that is being actively shaped.
Other Recommendation Systems
Assuming that senders mechanically provide a binary recommendation is adequate in view of their limited willingness to invest more time into giving a recommendation. Nonetheless, one could imagine more sophisticated recommendation systems. For instance, the sender might recommend a product when it is either good or when it is controversial but preferred by most of the population. This, however, presupposes a sender with detailed information about the receiver’s decision environment. Instead, the sender could give a buy recommendation when the average payoff she expects receivers to obtain from consuming the product in question is higher than for the outside option. In that case, the expected value of the outside option would determine the recommendation threshold, which would then no longer be a design parameter.
Product Ranking
When $q_{H}>0$ and $q_{L}>0$, all types agree on which product version is the best and which is the worst because their utility is fully determined by the qualities in each dimension and is increasing in each of them. In other environments, no products are objectively the best or the worst. For instance, there would be no such products in a modified version of our setting where $q_{H}=0$ and $q_{L}=0$, as the valuations for $(1,0)$ and $(0,1)$ are perfectly negatively correlated. We nonetheless keep the specific feature of the model because we believe it covers a wide range of plausible scenarios.
4 Optimal Receiver Behavior
In this section, we ask under which circumstances a receiver will accept recommendations, depending on his type $i$ and the recommendation system $\mathcal{R}$. In Section 4.1, we discuss through which channels recommendations affect beliefs, distinguishing between the objective and subjective content of the recommendations. Section 4.2 then characterizes the optimal behavior of the receiver by identifying thresholds which divide the type set into receivers that accept the recommendation and those that do not.
4.1 The Effects of Recommendations on Beliefs
Recommendations contain coarse information about the product under consideration, leading the receiver to update his beliefs about its quality. We first focus on buy recommendations, which have the following intuitive effects: [^11] In Appendix A.1, we provide a more formal analysis, decomposing the difference between posteriors and priors into three corresponding components.
  1. (I)
    The probability of a bad product falls to zero.
  2. (II)
    A good product becomes more likely relative to the controversial products.
  3. (III)
    The odds between the controversial products change.
The first two effects in this belief-updating decomposition are objectively positive: They correspond to increases of the expected payoff for every receiver type, albeit of different sizes. By contrast, the third effect is fully type-dependent and thus subjective, resulting in an increase in the payoff for some receivers and a decrease for others. The following distinction is useful:
Definition 2.
A recommendation $r\in\{B,D\}$ that induces posteriors $\mathbf{p}^{r}$ given prior $\mathbf{q}$
  1. (i)
    shifts probability from $(0,1)$ to $(1,0)$ if $p_{2}^{r}-q_{2}<p_{1}^{r}-q_{1}$;
  2. (ii)
    shifts probability from $(1,0)$ to $(0,1)$ if $p_{2}^{r}-q_{2}>p_{1}^{r}-q_{1}$;
  3. (iii)
    does not affect the probability between $(0,1)$ and $(1,0)$ otherwise.
As we will lay out below, whether a receiver considers the subjective content of a recommendation as positive or negative depends on his type $i$ and on the direction in which the recommendation shifts probability.
The effects of don’t-buy recommendations are analogous to those of buy recommendations: First, the probability of a good product falls to zero; second, bad products become more likely relative to controversial products, and third, the odds between the controversial products change. Thus, again, the recommendation has two objective effects (but these are both bad now) and a subjective effect (which can be good or bad, depending on the receiver’s type).
4.2 Characterizing Optimal Receiver Behavior
We are now in a position to understand whether it is optimal for the receiver to accept a recommendation. To this end, denote by $U_{i}^{B}$ and $U_{i}^{D}$ player $i$’s expected payoff of buying a product with a buy and don’t-buy recommendation, respectively, and by $U_{i}^{0}$ the expected payoff of buying the alternative product with no recommendation. We have
$U_{i}^{r}$ $:=p_{H}^{r}+(1/2+i)p_{1}^{r}+(1/2-i)p_{2}^{r},\quad r\in\{B,D\};$ (3)
$U_{i}^{0}$ $:=q_{H}+(1/2+i)q_{1}+(1/2-i)q_{2}.$ (4)
For an expected-payoff maximizer, it is optimal to accept a buy recommendation if and only if $U_{i}^{B}-U_{i}^{0}\geq 0$, that is, the expected payoff from buying the recommended product is at least as high as the expected payoff from buying one of the alternative products (that have no recommendation). [^12] We use the tie-breaking rule that indifferent agents accept the recommendation. This condition can equivalently be written as
$p_{H}^{B}-q_{H}+\frac{p_{1}^{B}-q_{1}}{2}+\frac{p_{2}^{B}-q_{2}}{2}\geq i[(p_{2}^{B}-q_{2})-(p_{1}^{B}-q_{1})].$ (5)
To formulate this in compact form, we introduce the following terminology.
Definition 3.
  1. (i)
    The objective effect of a recommendation $r$ is defined as
    $\Delta_{O}^{r}:=p_{H}^{r}-q_{H}+\frac{p_{1}^{r}-q_{1}}{2}+\frac{p_{2}^{r}-q_{2}}{2}.$ (6)
  2. (ii)
    The subjective effect of a recommendation is defined as
    $\Delta_{S}^{r}:=(p_{2}^{r}-q_{2})-(p_{1}^{r}-q_{1}).$ (7)
We can now restate condition (5) as
$\Delta_{O}^{B}\geq\Delta_{S}^{B}i.$ (8)
We use this formulation to justify our terminology. For the unbiased receiver $i=0$, equation (8) simplifies to $\Delta_{O}^{B}\geq 0$. As this receiver does not care about the shift in probabilities between the two controversial versions of the products, $\Delta_{O}^{B}\geq 0$ fully reflects the two objective effects (I) and (II) in the belief-updating decomposition in Section 4.1. For $i\neq 0$, the issue is more subtle. As long as $q_{1}=q_{2}$, the expected payoff conditional on the information that a product is controversial is $1/2$ for all $i$ and thus type-independent. Therefore, all types gain the same amount from (I) learning that a product is good or controversial instead of bad; the same applies to (II) learning that a product is good instead of controversial. For $q_{1}=q_{2}$, $\Delta_{O}^{B}$ therefore exactly captures these two type-independent (and thus objective) effects of a buy recommendation. [^13] The value of learning that the recommended product is more likely to be good is $p_{H}^{r}-q_{H}$, and the value of the change in the probability that it is controversial is $\frac{p_{1}^{r}+p_{2}^{r}-(q_{1}+q_{2})}{2}$. For $i\neq 0$ and $q_{1}\neq q_{2}$, the issue is complicated further. Though the effects captured by (I) and (II) in Section 4.1 are positive for all types, their size depends on the receiver’s type. Intuitively, if one of the controversial product versions (say (1,0))(1,0))( 1 , 0 ) ) is more likely than the other ex ante, learning that a product is controversial rather than bad is more valuable for types who prefer $(1,0)$ over $(0,1)$. Similarly, learning that the product is good rather than controversial is less valuable for those types. Even in this case, the interpretation of $\Delta_{O}^{B}\geq 0$ as an objective effect is justified because of the general agreement on its sign. We can now state our first result.
Proposition 1.
For a given recommendation system $\mathcal{R}$, a receiver of type $i$
  1. (i)
    accepts a buy recommendation if and only if he accepts a don’t-buy recommendation;
  2. (ii)
    accepts a recommendation if and only if $\Delta_{O}^{B}\geq i\Delta_{S}^{B}$.
Intuitively, a receiver accepts a recommendation if he thinks that it has sufficient objective content and/or that his preferences are sufficiently aligned with the sender’s. Whether the recommendation is positive or negative does not matter for this assessment. Rather, as laid out in the decomposition in Section 4.1, the objective effect $\Delta_{O}^{B}$ of a buy recommendation is positive, as the latter makes a good product more likely relative to the controversial products and rules out a bad one. By contrast, the sign of the subjective effect $\Delta_{S}^{B}$ can be positive or negative. In short, a receiver accepts a buy recommendation unless the subjective effect is sufficiently large and his preferences are not sufficiently well aligned with those of the rest of the population.
Denoting the type who is indifferent between accepting and not accepting recommendations as
$\tilde{i}:=$ $\frac{\Delta_{O}^{B}}{\Delta_{S}^{B}}\text{ if }\Delta_{S}^{B}\neq 0\text{,}$ (9)
we can summarize optimal receiver behavior as follows.
Corollary 1.
  1. (i)
    If $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$, then all $i\in[-1/2,1/2]$ accept the recommendation.
  2. (ii)
    If $\Delta_{S}^{B}<-2\Delta_{O}^{B}$, then all $i\geq\tilde{i}$ accept the recommendation.
  3. (iii)
    If $\Delta_{S}^{B}>2\Delta_{O}^{B}$, then all $i\leq\tilde{i}$ accept the recommendation.
The intuition for Corollary 1, the proof of which is omitted, is straightforward. For $\Delta_{S}^{B}=0$, the recommendation does not shift probability between the controversial product versions. Accordingly, accepting the recommendation is optimal for all players because of its positive objective effect $\Delta_{O}^{B}\geq 0$. For $\Delta_{S}^{B}\neq 0$, the conclusion is the same as long as $|\Delta_{S}^{B}|$ remains small enough—as required by (i). In that case, even the receivers who are most adversely affected by the subjective content of the recommendation (those at $i=1/2$ or $i=-1/2$) are convinced by its relatively objective character. As $|\Delta_{S}^{B}|$ increases, the role of the probability shift becomes apparent. As types $i>0$ prefer (1,0) over (0,1), a shift of the probability from the latter to the former induced by a recommendation ($\Delta_{S}^{B}<0$) increases such receivers’ expected payoff, reinforcing the objective effect. By contrast, such a shift reduces the expected payoff of types $i<0$. The condition in (ii) always holds for $i>0$, as the objective and subjective effects reinforce each other. It also holds for $i<0$ as long as $|i|<|\tilde{i}|$ so that the adverse subjective effect does not dominate the objective effect. Similarly, when the recommendation shifts the probability from (1,0) to (0,1) (when $\Delta_{S}^{B}>0$), all types $i<0$ accept. By contrast, types $i>0$ experience a reduction in their value, which can outweigh the positive objective effect. The condition in (iii) makes sure that the latter case does not arise. [^14] The effect on the expected payoffs in cases (ii) and (iii) is similar to the effect of a demand rotation as discussed in Johnson and Myatt (2006): For the types who accept the recommendation, the expected payoffs increase; for those who do not, they fall.
Finally, we analyze receiver behavior for extreme recommendation thresholds.
Proposition 2.
Suppose $R\rightarrow 1$ or $R\rightarrow 0$. Then, all types accept the recommendation irrespective of the product distribution.
In the limits, both types of recommendations essentially contain only objective information. For instance, when $R\rightarrow 1$, buy recommendations require that the product is objectively good, while don’t-buy recommendations require that it is not. Either way, the recommendation does not reveal anything about the relative probabilities of the two controversial product versions. Instead, it provides objective information that is equally valuable for all types, irrespective of their preferences for controversial products. Thus, all types accept the recommendation.
5 The Value of the Recommendation System
We now deal with the value of recommendation systems. In Section 5.1, we provide a characterization result. Section 5.2 asks how to use the recommendation threshold to influence the value. Finally, we investigate the effects of the preference and quality distributions on the value in Section 5.3.
5.1 Characterizing the Value of the Recommendation System
Before defining the value of a recommendation system, we describe the expected payoffs of a fixed type $i$ receiver, depending on whether he has access to a recommendation system and whether he accepts its recommendations. Without a recommendation system, all products are identical ex ante, with an expected payoff based on the prior distribution of product qualities. Hence, the expected payoff of a receiver $i$ is
$V_{0}(i):=U_{i}^{0}.$ (10)
Now suppose one of the products comes with a buy or don’t-buy recommendation, so the receiver has to update his belief about the quality distribution. A receiver who accepts the recommendation chooses this product (rather than one of the alternatives) only if it comes with a buy recommendation. With a recommendation system, the expected payoff of a receiver $i$ who accepts a recommendation is thus
$V_{A}(i):=\pi^{B}U_{i}^{B}+(1-\pi^{B})U_{i}^{0}.$ (11)
The first term is the probability of receiving a buy recommendation multiplied by the expected payoff from accepting it (and thus buying the recommended product). The second term is the probability of a don’t-buy recommendation multiplied by the expected payoff from buying the product without any recommendation, that is, based on priors. [^15] The notation $V_{A}(i)$ emphasizes the dependence of this expected payoff on the type of the receiver. Clearly, $V_{A}$ also depends on $R$ and on the decision environment; similarly for the expressions below.
A receiver who does not accept a recommendation will buy the recommended product if it comes with a don’t-buy recommendation. Otherwise, he will buy one of the other products. Hence, the expected payoff of such a receiver is
$V_{N}(i):=$ $\pi^{B}U_{i}^{0}+(1-\pi^{B})U_{i}^{D}.$ (12)
A buy recommendation does not create value for receivers who do not follow it, as such types choose the alternative for which no recommendation is available. Instead, the don’t-buy recommendation, which is given with probability $(1-\pi^{B})$, does influence the behavior of the types who do not accept it: Instead of buying an alternative based on prior probabilities, updating leads them to choose the product for which they received the explicit recommendation not to buy.
We now define the value of a recommendation system. [^16] This definition is in line with the definition of a valuable signal in Kamenica and Gentzkow (2011).
Definition 4.
$V(\mathcal{R})$, the value of a recommendation system $\mathcal{R}$, is the expected increase in payoffs resulting from the existence of recommendations, where the expectation is taken over all pairs of independently drawn senders and receivers.
Using Definition 4, we can characterize $V(\mathcal{R})$ as follows:
vertical space
[…]
vertical space
The simple structure of $V(\mathcal{R})$ for $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$ in (i) results because, by Corollary 1, all types accept a recommendation. The difference between the expressions for $\Delta_{S}^{B}<-2\Delta_{O}^{B}$ in (ii) and $\Delta_{S}^{B}>2\Delta_{O}^{B}$ in (iii) arises because only types $i>\tilde{i}$ accept the recommendation in the former case, whereas only types $i<\tilde{i}$ do so in the latter.
We now express the value $V(\mathcal{R})$ in terms of its objective and subjective content.
Proposition 3.
The value $V(\mathcal{R})$ of the recommendation system $\mathcal{R}$ is:
vertical space
(i) $\pi^{B}[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i]]$ if $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$; (ii) $(1-\pi^{B})F(\tilde{i})\left[\Delta_{O}^{D}-\Delta_{S}^{D}\mathbb{E}[i\mid i\leq\tilde{i}]\right]+\pi^{B}(1-F(\tilde{i}))\left[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i\mid i\geq\tilde{i}]\right]$ if $\Delta_{S}^{B}<-2\Delta_{O}^{B}$; (iii) $\pi^{B}F(\tilde{i})\left[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i\mid i\leq\tilde{i}]\right]+(1-\pi^{B})(1-F(\tilde{i}))\left[\Delta_{O}^{D}-\Delta_{S}^{D}\mathbb{E}[i\mid i\geq\tilde{i}]\right]$ if $\Delta_{S}^{B}>2\Delta_{O}^{B}$.
Intuitively, when $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$, Corollary 1 implies that all receivers accept the recommendation as the objective effect dominates. Hence, by equation (11), the recommendation system creates value only in case of a buy recommendation, as a don’t-buy recommendation leads to the same expected payoff as without a recommendation system. The positive payoff effect of the recommendation reflects the type-independent objective part $\Delta_{O}^{B}$ and the subjective part $\Delta_{S}^{B}$, which is weighted by the average type. Finally, all of that only materializes with the probability of a buy recommendation, $\pi_{B}$.
As the absolute value of $\Delta_{S}^{B}$ increases, the analysis becomes more complex because some types do not follow the recommendation. We consider the case $\Delta_{S}^{B}>2\Delta_{O}^{B}$ and reproduce the expression for $V(\mathcal{R})$ as follows: [^17] The argument for $\Delta_{S}^{B}<-2\Delta_{O}^{B}$ is analogous.
$\underbrace{\pi^{B}F(\tilde{i})\left[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i\mid i\leq\tilde{i}]\right]}_{\text{types accept recommendation}}+\underbrace{(1-\pi^{B})(1-F(\tilde{i}))\left[\Delta_{O}^{D}-\Delta_{S}^{D}\mathbb{E}[i\mid i\geq\tilde{i}]\right]}_{\text{types reject recommendation}}.$ (13)
In this case, the buy recommendation shifts probability from $(1,0)$ to $(0,1)$, which is unfavorable for receivers $i>0$, reducing their gains from accepting and possibly making them negative. Thus, for $\Delta_{S}^{B}>2\Delta_{O}^{B}$, only types below $\tilde{i}$ (a fraction $F(\tilde{i})$ of the receivers) accept the buy recommendation, which comes with probability $\pi^{B}$. In addition to the objective value of the recommendation ($\Delta_{O}^{B}$), these receivers obtain a subjective value. In (13), the subjective value is reflected in $-\Delta_{S}^{B}\mathbb{E}[i\mid i\leq\tilde{i}]$, the average effect of the shift in the relative probabilities of $(1,0)$ and $(0,1)$ on the value of accepting the recommendation for these buyers. The sign of this term depends on the properties of $F$ captured in $\mathbb{E}[i\mid i\leq\tilde{i}]$. It is positive if the negative contribution of types $i<0$ dominates the positive contribution of types $i>0$. The second component in (13), corresponding to the types who do not accept the recommendation, is arguably more surprising. Given their lack of alignment with the rest of the population, these types treat the recommendation not to buy as good news about the product and thus do not accept the recommendation. [^18] In detail, a don’t-buy recommendation (which happens with probability $1-\pi^{B}$) carries an adverse objective content (summarized by $\Delta_{O}^{D}$ in the term in squared brackets). However, for receivers with $i>0$ it also carries the good news that the product is more likely to be of the preferred version $(1,0)$ rather than $(0,1)$. Those with $i>\tilde{i}(>0)$ then buy the product despite the adverse objective content, valuing the beneficial subjective effect with $-\Delta_{S}^{D}\mathbb{E}[i\mid i\geq\tilde{i}]>0$ on average. Such receivers exist only if $\Delta_{S}^{B}$ is sufficiently large (and hence $\Delta_{S}^{D}$ is sufficiently negative relative to $\Delta_{O}^{D}$). Nonetheless, its informational content creates value for them.
In the following, we shall often invoke the following symmetry assumption on the preference distribution.
Assumption 1.
$F(-i)=1-F(i)$ for all $i\in[-1/2,1/2]$.
With this assumption in place, the quantity
$\beta:=F(1/2-R)$ (14)
has a simple interpretation. Given a threshold $R$, the parameter $\beta$ captures the fraction of the population that is willing to recommend product versions $(1,0)$ or, equivalently (given the symmetry of the distribution), $(0,1)$. For a fixed type distribution $F$, the variables $\beta$ and $R$ are inversely related. We then obtain the following result:
Proposition 4.
Suppose Assumption 1 holds. Then, all types accept the recommendation and
$V(\beta)=\pi^{B}\Delta_{O}^{B}\text{.}$
To understand why all types accept a recommendation, the belief-updating decomposition in Section 4.1 is helpful. Its third component captures the potential change in the odds of the controversial products, which induces a purely type-dependent payoff effect. Given the symmetry assumption, both controversial product versions lead to a buy recommendation with probability $\beta$. Hence, a recommendation does not change the odds of the controversial products. As the other two effects of a buy recommendation on payoffs are positive for all types, the buy recommendation increases expected payoffs for any type, yielding the result. [^19] An analogous argument applies for don’t-buy recommendations. Proposition 3 thus implies $V(R)=\pi^{B}\left[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i]\right]$. Further, because of the symmetry of the population, $\mathbb{E}[i]=0$, so $V(R)=\pi^{B}\Delta_{O}^{B}$ as required.
5.2 Optimal Design of Recommendation Systems
We now use our characterization of the value of a recommendation system to address design issues, focusing on the optimal choice of the threshold. We start with the case of a symmetric population, then we treat asymmetric populations.
5.2.1 Symmetric Type Distribution
In the symmetric setting, we obtain a simple expression for $V(R)$, yielding transparent comparative statics results. [^20] As we focus on $R$ as the design parameter, we often write $V(R)$ instead of $V(\mathcal{R})$. To this end, a reparametrization is helpful:
Definition 5.
The prevalence of controversial products is given as $Q:=\frac{q_{1}+q_{2}}{2}$. The odds of a good product are $\sigma:=q_{H}/q_{L}$.
Using these parameters, we immediately obtain:
Lemma 1.
Suppose Assumption 1 holds. Then,
$\pi^{B}$ $=\frac{(1-2Q)\sigma}{1+\sigma}+2Q\beta;$
$\Delta_{O}^{B}$ $=\frac{\beta Q(\sigma+1)-2Q\sigma+\sigma}{2\beta Q(\sigma+1)-2Q\sigma+\sigma}-\frac{(1-2Q)\sigma}{1+\sigma}-Q;$
$V(\beta)$ $=\pi^{B}\Delta_{O}^{B}=(1-2Q)\frac{\sigma+Q\left(\beta-\sigma+\sigma^{2}(1-\beta)\right)}{(\sigma+1)^{2}}.$
The value of the recommendation system depends on $\beta$ as well as on the prevalence of controversial products $Q$ and on the odds ratio $\sigma$ between good and bad products. Recall that, for each of the two controversial product versions, $\beta$ measures the fraction of the population willing to recommend it. Therefore, the parameter captures the influence of the distribution and the recommendation threshold on the value. [^21] Observe that by varying $R$ we can obtain any $\beta\in(0,1)$. We can interpret changes of $\beta$ as either changes of the threshold $R$ for a given distribution $F$ or as changes of the distribution $F$ for a given threshold $R$. The following result abstracts from this distinction, allowing for either interpretation. It shows that countervailing effects of $\beta$ on the probability of a buy recommendation and the objective value $\Delta_{O}^{B}$ of accepting it result in a non-trivial effect on $V(\beta)=\pi^{B}\Delta_{O}^{B}$.
Lemma 2.
Suppose Assumption 1 holds. Then,
  • (i)
    $\pi^{B}$ is strictly increasing and $\Delta_{O}^{B}$ strictly decreasing in $\beta$;
  • (ii)
    $\frac{\partial^{2}V}{\partial\beta\partial\sigma}<0$ and $V^{\prime}(\beta)=\begin{cases}>0&\text{if }\sigma<1;\\ =0&\text{if }\sigma=1;\\ <0&\text{if }\sigma>1.\end{cases}$
The first statement in (i) holds because as $\beta$ increases a greater share of the population is willing to recommend a controversial product. Moreover, conditional on receiving a buy recommendation, the recommended product is thus more likely to be controversial. Therefore, the objective payoff component $\Delta_{O}^{B}$ decreases as $\beta$ increases, implying the second statement in (i). The effect of $\beta$ on $V(\beta)$ in (ii) reflects the effect on its two components. As $\sigma$ increases (good products become relatively more likely than bad ones), a buy recommendation becomes more likely. Hence, a high $\Delta_{O}^{B}$ is particularly valuable when $\sigma$ is high so that increasing $\sigma$ and lowering $\beta$ are complementary ($\frac{\partial^{2}V}{\partial\beta\partial\sigma}<0$). [^22] More technically, repeated application of the product rule shows that $\frac{\partial^{2}V}{\partial\beta\partial\sigma}=\frac{\partial\pi^{B}}{\partial\beta}\frac{\partial\Delta_{O}^{B}}{\partial\sigma}+\frac{\partial\pi^{B}}{\partial\sigma}\frac{\partial\Delta_{O}^{B}}{\partial\beta}+\frac{\partial^{2}\pi^{B}}{\partial\beta\partial\sigma}\Delta_{O}^{B}+\frac{\partial^{2}\Delta_{O}^{B}}{\partial\beta\partial\sigma}\pi^{B}$. The (incomplete) intuition just given corresponds to the first term being positive. The value-maximizing level of $\beta$ must thus be weakly decreasing in $\sigma$. The symmetry assumption turns out to render the problem linear in $\beta$, so the second result in (ii) follows. Using $\beta=F(1/2-R)$, Lemma 2 directly implies the following result. [^23] Result (iii) follows because $V(\beta)$ is independent of $\beta$ in this case.
Proposition 5.
Suppose Assumption 1 holds. Then, $V(R)$ is
  1. (i)
    decreasing in $R$ if $\sigma<1$;
  2. (ii)
    increasing in $R$ if $\sigma>1$;
  3. (iii)
    constant in $R$ if $\sigma=1$.
By Lemma 2, an increase in $R$ reduces the probability of a buy recommendation, while it increases $\Delta_{O}^{B}$, the expected average payoff conditional on a buy recommendation. These effects exactly offset each other if and only if $\sigma=1$. Otherwise, one effect dominates the other. Based purely on uncontroversial products, a buy recommendation is more likely than a don’t-buy recommendation when the good product is relatively more likely than the bad one ($\sigma>1$). It is thus worthwhile to increase the expected average payoff conditional on a buy recommendation by increasing $R$ at the expense of the probability of receiving such a recommendation. Put differently, the designer chooses between revealing whether the product is good or whether it is bad. He moves the threshold in the direction which yields full revelation with higher probability, which hinges only on the odds ratio of the uncontroversial product versions.
5.2.2 Asymmetric Type Distribution
For a symmetric population, the subjective component cancels out in the value of the recommendation system, allowing for a focus on the role of the objective content. Moreover, symmetry simplifies the analysis by ensuring that all types accept the recommendation. However, the assumption of a symmetric population may prove restrictive in some contexts, and it does not allow us to study the role of the subjective component for the optimal design of the recommendation system. To make headway in that direction, we introduce the following assumption.
Assumption 2.
(i) $F(i)=(i+1/2)^{a}$ for $a>0$; (ii) $q_{1}=q_{2}=Q$.
We thus capture asymmetry by means of a simple but flexible parameterization. By assuming that the controversial products are equally likely a priori, we ensure that any effect of a recommendation exclusively reflects the population asymmetry. Corollary 1 implies the following result:
Corollary 2.
Suppose Assumption 2 holds. Then, we have:
[…]
Thus, whether and how a recommendation shifts the probabilities of the controversial product versions depends entirely on the type distribution. When it is symmetric ($a=1$), there is no shift. When it is convex ($a>1$), so that more people prefer version $(1,0)$ to $(0,1)$, a buy recommendation shifts probability from $(0,1)$ to $(1,0)$, as it is more likely to come from a person who prefers $(1,0)$ to $(0,1)$. Similarly, when the type distribution is concave ($a<1$), the probability is shifted to $(0,1)$.
We can now analyze the extent to which the result in Proposition 5 that, generically, extreme thresholds are desirable is an artifact of the symmetry assumption. This is challenging because, contrary to the symmetric case, the set of types that accept a recommendation depends on the threshold $R$. Reflecting this complication, there is no direct counterpart to Proposition 5. Nevertheless, the following two results provide some insights into the optimal recommendation system in the case of an asymmetrically distributed population.
Proposition 6.
Suppose Assumption 2 holds. An optimal threshold $R^{\ast}\in(0,1)$ exists if $Q>1-\max\{\frac{a}{a+1},\frac{1}{a+1}\}$ or if (i) or (ii) holds:
[…]
Figure 2: Illustration of Propositions 6 and 7
As illustrated in Figure 2, Proposition 6 provides sufficient conditions for interior solutions. It shows that an interior optimum exists when (i) controversial products are likely ($Q$ is large) and the population is very asymmetric ($a$ is far away from $1$) or otherwise when (ii) the uncontroversial product versions are similarly likely ($\sigma$ is close to $1$). The result starkly contrasts the symmetric case, where an interior solution is only possible in the knife-edge case $\sigma=1$. In that case, recommendations do not reveal any information about the relative likelihood of the controversial products beyond the prior, so the subjective payoff effect cancels out and the optimal threshold depends only on the odds ratio $\sigma$ of the uncontroversial products. In contrast, the average subjective payoff effect does not cancel out in the asymmetric case. Thus, the choice of the optimal threshold no longer only depends on $\sigma$ but also on the asymmetry parameter $a$ and the prevalence of controversial products $Q$. For a wide range of parameters, the conditions given in Proposition 6 are necessary as well as sufficient:
Proposition 7.
Suppose Assumption 2 holds and fix $\sigma$. Then, as $Q\rightarrow 0$, $V(R)$ is
  1. (i)
    decreasing in $R$ if $\sigma\leq\min\left\{\frac{1-Q(a+1)}{a-Q(a+1)},\frac{a-Q(a+1)}{1-Q(a+1)}\right\}$;
  2. (ii)
    increasing in $R$ if $\sigma\geq\max\left\{\frac{1-Q(a+1)}{a-Q(a+1)},\frac{a-Q(a+1)}{1-Q(a+1)}\right\}$;
  3. (iii)
    maximized for some $R^{\ast}\in(0,1)$ otherwise.
The proof essentially proceeds in two steps. First, all types accept the recommendation when $Q$ is sufficiently small, since then the subjective content in the recommendation vanishes.242424 $Q\rightarrow 0$ implies that $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$, ensuring that all types follow by Corollary 1. The key step is that $\Delta_{S}^{B}\rightarrow 0$ as $Q\rightarrow 0$, while $\Delta_{O}^{B}\rightarrow 1-q_{H}>0$ when $Q\rightarrow 0$. The proof shows that, even for larger values of $Q$, $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$ often holds, explaining why the conclusion of Proposition 7 holds not only in the limit. This simplifies $V(R)$ by eliminating the case distinctions. Leveraging this, we can prove the quasiconcavity of the function, ensuring that the sufficient conditions in Proposition 6 are necessary, too.
The requirement that $Q\rightarrow 0$ is much stronger than necessary for this result. Even for higher values of $Q$, Figure 2 quite accurately separates the parameter range under which an interior solution exists from the area where it does not. [^25] The figure is drawn for $Q=0.1$, in which case the if-and-only-if statement is correct for $\sigma\in[1/60,60]$ and $a\in[1/10000,10000]$. In a nutshell, the result shows that an interior solution exists if (i) controversial products are sufficiently prevalent, (ii) the population asymmetry is substantial, and (iii) the odds ratio of the uncontroversial products is sufficiently small or large. To understand this intuitively and to contrast it with the symmetric case, suppose that $\sigma>1$. Hence, the value of the recommendation system is increasing in $R$ in the symmetric case, as the gain in the expected payoff conditional on receiving a buy recommendation outweighs the loss from the lower probability of a buy recommendation. This follows because the subjective effect in $V(R)=\pi^{B}[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i]]$ cancels out with symmetry. With an asymmetric population, the term $-\Delta_{S}^{B}\mathbb{E}[i]$ does not vanish. To illustrate this, consider a very asymmetric preference distribution such as $F(x)=(x+1/2)^{10}$, for which the probability of a type $i\leq 0$ is less than 0.1%. Then, $\mathbb{E}[i]>0$ and, by Corollary 2, $\Delta_{S}^{B}<0$ so that $-\Delta_{S}^{B}\mathbb{E}[i]>0$. Hence, the subjective effect is positive on average. Intuitively, though this population is distributed very asymmetrically around zero, it is essentially homogeneous in the sense that almost all types agree on the ranking of the controversial products: If such a product results in a buy recommendation, it almost surely has quality vector (1,0), as almost all types prefer this to $(0,1)$. As such, these products are not that controversial anymore, and even the subjective effect is “fairly objective.” Now, as $R$ increases, the probability of a buy recommendation still falls, and the “purely objective” payoff part $\Delta_{O}^{B}$ increases. However, for $R$ high enough, this “fairly objective” payoff part $-\Delta_{S}^{B}\mathbb{E}[i]$ also decreases, as an extreme threshold $R$ impedes the ability to distinguish between a good product and a widely preferred controversial product. While consumers are homogeneous regarding which controversial product version they prefer, they are still heterogeneous with respect to the intensity with which they do so. Hence, disentangling the good and the preferred controversial product is worthwhile. Therefore, an interior optimal threshold results.
5.3 Preference Heterogeneity and Controversial Products
As discussed after Proposition 7, preference heterogeneity only matters when controversial products exist. Moreover, the less homogeneous the population, the more likely it is that extreme thresholds are desirable so that receivers learn solely about uncontroversial products. Hence, one might conjecture that greater preference heterogeneity and a higher prevalence of controversial products reduce the value of recommendation systems. In the following, we address both issues in turn, demonstrating that these conjectures do not hold without qualification. For tractability, we focus on symmetric populations.
5.3.1 Preference Heterogeneity
Under Assumption 1, we can analyze increases in the polarization of the distribution of preferences by keeping the threshold $R$ fixed and varying $\beta$ [^26] For $R=1/2$, we must have $\beta=1/2$ because the symmetry of $F$ requires $F(0)=1/2$, so this exercise is only meaningful for thresholds $R\neq 1/2$. Further, for $R>1/2$ we must have $\beta\leq 1/2$ because $\beta=F(1/2-R)\leq F(0)=1/2$. Accordingly, for $R<1/2$ we must have $\beta\geq 1/2$. This corresponds to a mean-preserving spread of $F$ (see Figure 3), coinciding with an increase in $\beta$ for $R>1/2$ and a decrease for $R<1/2$. We obtain the following result.
Proposition 8.
Suppose Assumption 1 holds.
  1. 1.
    Suppose $\sigma<1$. Then, a mean-preserving spread of the type distribution increases the value of the recommendation system for $R>1/2$ and decreases it for $R<1/2$.
  2. 2.
    Suppose $\sigma>1$. Then, a mean-preserving spread of the type distribution decreases the value of the recommendation system for $R>1/2$ and increases it for $R<1/2$.
To gain some intuition, suppose $\sigma<1$ and $R>1/2$. Figure 1 shows that with $R>1/2$, only extreme types give buy recommendations for controversial products. [^27] Conversely, for $R<1/2$ only relatively extreme types give don’t-buy recommendations for controversial products. This can be seen by flipping the colors in Figure 1 so that it corresponds to $R=1/4$. Increasing $\beta$ corresponds to a mean-preserving spread of $F$ given $R$. Consequently, there are more such extreme senders and, hence, buy recommendations are increasingly likely to result from controversial products. This increases the probability of buy recommendations while lowering their objective value. The resolution of this trade-off, and thus whether the value of the recommendation system goes up or down, is determined by the odds ratio $\sigma$ (see Proposition 5). For instance, if $\sigma<1$, good products are rare compared to bad ones, so the value of the recommendation system is increasing in $\beta$. Therefore, the positive effect of revealing bad products by means of don’t-buy recommendations outweighs the negative effect of a reduced value conditional on receiving a buy recommendation. Thus, given a high recommendation threshold ($R>1/2$), a more polarized population benefits more from the recommendation system, as this means only bad products generate don’t-buy recommendations. Conversely, when the recommendation threshold is relatively low ($R<1/2$), a less polarized population centered around the middle gives don’t-buy recommendations only to bad products.
Figure 3: Population polarization
[…]
5.3.2 Prevalence of Controversial Products
Without controversial products, preference heterogeneity would not matter: All receivers would accept recommendations, as they would reveal whether the product is good or bad. Hence, the value of the recommendation system would be independent of the preference distribution. More broadly, we now ask how an increasing prevalence of controversial products (an increase in $Q$) affects the value of a recommendation system. A natural conjecture is that recommendations become less valuable when controversial products are more likely. We find that this conjecture needs to be qualified.
Corollary 3.
Suppose Assumption 1 holds.
  • (i)
    If $03\sigma-\beta-\sigma^{2}+\sigma^{2}\beta>$ 0, then the value of the recommendation system is decreasing in $Q$. In particular, this is the case when $\sigma=1$.
  • (ii)
    If $03\sigma-\beta-\sigma^{2}+\sigma^{2}\beta<$ 0, then $Q^{*}:=\frac{3\sigma-\beta-\sigma^{2}+\sigma^{2}\beta}{4\sigma-4\beta-4\sigma^{2}+4\sigma^{2}\beta}\in(0,\frac{1}{2})$ maximizes $V(R)$.
Figure 4: Illustration of Corollary 3
(a) Panel A
$\sigma$$\beta$111133335555777799990.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81.01.01.01.0
(b) Panel B
$\sigma$$\beta$111133335555777799990.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81.01.01.01.0
(c) Panel C
$\sigma$$\beta$111133335555777799990.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81.01.01.01.0
Figure 4 illustrates the result. In the white area in Panel A, $V(\beta)$ is decreasing in $\beta$; in the shaded area, an interior maximizer exists. When good and bad products are similarly likely ($\sigma$ is close to $1$), a higher prevalence of controversial products indeed reduces $V(\beta)$. For very high or low values of $\sigma$, however, this need not be true. This surprising result reflects the multiplicative structure of $V(\beta)$. Under Assumption 1, $V(\beta)$ is the product of the buying probability $\pi^{B}$ and the objective value of the recommendation per buyer, $\Delta_{O}^{B}$. The effect of $Q$ on $\pi^{B}$ is negative only if $\sigma$ is sufficiently large (see Panel B). Intuitively, as $Q$ increases, good products are less likely, reducing the probability of buy recommendations. However, there are now more buy recommendations stemming from controversial products, as these have become more likely. $\beta$ and $\sigma$ determine which of these effects dominates. When $\beta$ is high, the increase induced by controversial products leads to more buy recommendations, strengthening the positive effect. When $\sigma$ is high, the lower share of good products dominates, strengthening the negative effect. Similarly, one can show that the effect of $Q$ on $\Delta_{O}^{B}$ becomes negative only for small $\sigma$ (see Panel C). Again, this reflects countervailing forces. The negative effect results from the increased probability that a product with a buy recommendation is controversial. The positive effect reflects a change in the value of the outside option. The values of $\sigma$ and $\beta$ determine which effect dominates. [^28] When $\beta$ is small, the negative effect is weak, as only few types give buy recommendations for controversial products. When $\sigma$ is small, the outside option becomes more valuable, as bad products become less likely, weakening the positive effect.
6 Extensions
Our benchmark model relies on several assumptions. First, the preference distribution of senders and receivers is identical. Second, there are only two levels of recommendations (“buy” and “don’t-buy”). Third, each receiver only has access to a single recommendation. We now relax each of these assumptions in turn.
6.1 Distinct Sender and Receiver Distributions
We modify the benchmark model by assuming that the sender and the receiver are drawn from different distributions $F$ and $G$, respectively. For instance, this enables us to take into account that the “early adopters” of a new product may have different preferences than the general population (similar to Acemoglu et al. (2022)). We show that the results of the benchmark case easily generalize to this setting.
The sender distribution $F$ fully determines the probabilities of buy and don’t-buy recommendations, the posterior probabilities, and thus the objective and subjective effects and the expected payoffs, $V_{A}$ and $V_{N}$. Hence, all results on receiver behavior in Section 4 apply verbatim if $F$ is taken to be the sender distribution, as the receiver’s behavior is independent of the receiver distribution. The value of the recommendation system, however, depends on the receiver distribution $G$ as well, as it involves integration over this distribution. The next result generalizes Proposition 3.292929The notation for the quantities in Proposition 9 reflects their dependence on $F$ and $G$, respectively.
Proposition 9.
Consider the model with distinct preference distributions of senders and receivers. Suppose the sender distribution $F$ satisfies Assumption 1. Then, all types accept the recommendation. Further, $V(R)=\pi^{B}_{F}[\Delta_{O,F}^{B}-\Delta_{S,F}^{B}\mathbb{E}_{G}[i]]$ is
  1. (i)
    decreasing in $R$ if $\sigma<\frac{Q-(q_{1}-q_{2})\mathbb{E}_{G}[i]}{Q-(q_{2}-q_{1})\mathbb{E}_{G}[i]}$;
  2. (ii)
    increasing in $R$ if $\sigma>\frac{Q-(q_{1}-q_{2})\mathbb{E}_{G}[i]}{Q-(q_{2}-q_{1})\mathbb{E}_{G}[i]}$;
  3. (iii)
    constant in $R$ if $\sigma=\frac{Q-(q_{1}-q_{2})\mathbb{E}_{G}[i]}{Q-(q_{2}-q_{1})\mathbb{E}_{G}[i]}$.
Recall that in the benchmark model symmetry of $F$ implies that all receivers follow the recommendation and the subjective effect of the recommendation cancels out. This is still true if $F$ is distinct from $G$ as long as $G$ is also symmetric and hence $\mathbb{E}_{G}[i]=0$. For more general distributions $G$, all receivers still accept the recommendation for the same reason as above. However, as the receiver distribution is not symmetric, the subjective effect does not cancel out anymore. Even so, the result qualitatively mirrors Proposition 5, as changing the threshold still leads to the same trade-off: Reducing the threshold increases the probability of a buy recommendation but decreases its value conditional on that recommendation. However, in the latter effect, the value is no longer purely objective but consists of the objective part $\Delta_{O}^{B}$ and a subjective part given by $-\Delta_{S}^{B}\mathbb{E}_{G}[i]$, reflecting the receiver distribution $G$. Further, as the subjective effect now influences the value of the system, so does the prevalence of controversial products by affecting the tipping point at which the designer opts for a high or a low threshold. In any case, when $\mathbb{E}_{G}[i]=0$, we obtain the result from the baseline model as a special case. Similarly, the other results in the benchmark model remain (qualitatively) valid. With asymmetrically distributed populations, Propositions 6 and 7 hold for arbitrary receiver distributions as long as we impose Assumption 2 on the sender distribution.
6.2 Multiple Recommendation Thresholds
We now analyze the effects of reducing the coarseness of the recommendation system. We maintain all assumptions of the benchmark case, except that there are two threshold levels $R_{1},R_{2}\in(0,1)$ with $R_{2}\geq R_{1}$ and three types of recommendations $r\in\{D,N,B\}$ which we interpret as “don’t-buy,” “neither buy nor don’t-buy,” and “buy.” The sender gives a buy recommendation if the payoff is at least $R_{2}$, a don’t-buy recommendation if it is below $R_{1}$, and neither if it is in $[R_{1},R_{2})$. We maintain Assumption 1.
The proof of Proposition 10 in Appendix A.3 provides the receiver’s posteriors. For buy and don’t-buy recommendations, these posteriors are analogous to the benchmark, except that the different thresholds in each case have to be taken into account. As in the benchmark model, the highest and lowest recommendations rule out the bad and good product versions, respectively. The intermediate recommendation $r=N$, however, rules out both of these versions, so the receiver knows that the recommended product is controversial.
We say that a receiver accepts a recommendation if he buys if $r=B$ and never buys if $r=D$, allowing for either behavior if $r=N$. Assumption 1 again implies that the receiver always buys a product with $r=B$ and never buys if $r=D$. As noted above, $r=N$ rules out both uncontroversial products. Given the symmetric population, $r=N$ leaves the relative likelihood of the controversial products unchanged. Hence, whether a type $i$ wants to buy a product with $r=N$ depends on (i) how likely a good product is and (ii) how much he (dis-)likes the more likely controversial product. As in the benchmark model, we can derive the indifferent type so that, depending on parameters, either all, only sufficiently aligned, or no types buy the product with an intermediate recommendation.
$V(R)$ now consists of two components. First, as in the benchmark, all types accept $r=B$, resulting in the component $\pi^{B}\Delta_{O}^{B}$. [^30] As before, following don’t-buy recommendations does not create value. Moreover, $r=N$ can create value for the types that buy following that neutral recommendation. As noted above, parameter values determine which types do this. Despite this ambiguity, we can formulate the following result.
Proposition 10.
Consider the model with multiple recommendation thresholds. Suppose Assumption 1 holds. $V(R)$ is (i) decreasing in $R_{1}$ if $\sigma<1$ and (ii) increasing in $R_{2}$ if $\sigma>1$.
The result provides an incomplete characterization of the optimal recommendation system: It establishes that one of the two thresholds should be extreme so that one recommendation essentially becomes objective. Thus, as in Proposition 5, the designer maximizes the probability of an objective recommendation, considering which uncontroversial product version is more likely a priori. The optimal value of the threshold that is not pinned down by Proposition 10 depends on the model parameters unless $q_{1}=q_{2}$. In this degenerate case, any value for the other threshold is optimal, but it is also possible that both the high and the low recommendations are essentially turned objective. [^31] This arises for instance for $F(i)=1/2+i$ with parameters $q_{1}=0.7,\ q_{2}=0.1,$ and $q_{H}=0.03.$
6.3 Multiple Recommendations
We now show that the value of a recommendation system does not always increase if receivers obtain more than one recommendation: The additional recommendations do not add value if the ex ante probabilities of the good and bad products are similar while those of the two controversial product versions are not.
We assume that receivers have access to a number $b$ of buy recommendations and a number $d$ of don’t-buy recommendations. As in the benchmark model, recommendations come from randomly drawn senders. We add the assumption that they are drawn independently. [^32] In Appendix A.4, we provide the posteriors, depending on $b$ and $d$. Just as before, a single buy recommendation is enough to rule out a bad product, and a single don’t-buy recommendation is enough to rule out a good product. Consequently, when the receiver obtains mixed recommendations, that is, both buy and don’t-buy recommendations, he rules out both uncontroversial product versions. [^33] In Sun (2011), a higher variance in the recommendations corresponds to a niche product that some consumers like and some don’t. Similarly, in our paper, mixed recommendations (and thus a higher variance) result from controversial products that, indeed, some consumers like more than others.
To make our main point that multiple recommendations need not increase the value of the recommendation system as starkly as possible, we focus on the case of “infinite learning” where the receiver has access to an unlimited number of recommendations.
Proposition 11.
Consider the model with multiple recommendations. Suppose Assumption 1 holds. Let $\lambda:=q_{1}/q_{2}$. Compared with the optimal recommendation system in the benchmark model, infinite learning does not increase the value of the recommendation system if and only if either (i), (ii), or (iii) holds:
[…]
To grasp the result intuitively, recall the receiver’s behavior in the single-recommendation system. With a threshold of $R\rightarrow 0$, the receiver buys the recommended product unless it is bad. For $R\rightarrow 1$, he does so only if it is good. In the benchmark model, V(R) is decreasing in $R$ if $\sigma<1$ and increasing if $\sigma>1$. Lemma 4 in Appendix A.4 describes optimal receiver behavior with infinite learning. It shows that behavior coincides with the optimally designed single-recommendation system if the product is uncontroversial: Good products are bought but bad ones are not. Thus, the infinite-learning system potentially differs from the benchmark only for controversial products, which generate mixed recommendations. When $\sigma$ is very low, bad products are relatively likely, so the outside option looks relatively bleak. A consumer facing a product with mixed recommendations will thus buy this (controversial) product. In contrast, when $\sigma$ is very high, any such consumer will buy the outside option, as it is likely to be good. Thus, when the probabilities of the uncontroversial product versions are not too similar, the behavior with a single recommendation and with multiple recommendations coincides, and multiple recommendations do not increase the value of the recommendation system. However, multiple recommendations do affect behavior if the probabilities of the good and bad products are similar. For instance, when $\lambda>1$, low types who value the less likely controversial product version $(0,1)$ more strongly than product version $(1,0)$ will opt for the outside option rather than the controversial product. Then, the system’s value with multiple recommendations is higher than that of the single-recommendation system, as it provides more granular information. Inspecting the conditions on $\sigma$ in (ii) and (iii), we find that the more extreme (further away from 1) the odds ratio of the controversial products is, the more extreme the odds ratio of the uncontroversial product versions needs to be for there to be no value in additional recommendations.
7 Conclusion
We have set out to study the role of preference heterogeneity in recommendations, asking questions such as: Under what circumstances do people follow a recommendation? What determines the value of a recommendation system and when is it maximized? Our analysis highlights the importance of disentangling objective information from subjective preferences. We have seen that, depending on the preference distribution, the subjective content can reinforce the positive news of a buy recommendation or reverse it so that a buy recommendation is treated as a bad signal. Conversely, the subjective content can turn a don’t-buy recommendation into good news for some people. The design of a recommendation system, as summarized in the threshold $R$, should thus take into account how it influences the relative objectivity of recommendations.
We find that extreme recommendation systems may be valuable, as they convey objective information even when preferences are heterogeneous. However, when preferences are not distributed symmetrically and controversial products are sufficiently prevalent, with enough consensus on their ranking, the optimal recommendation threshold is set at an intermediate level. This allows receivers to not only draw inferences about the uncontroversial products but also about the controversial ones. For asymmetrically distributed populations, subjective content in the recommendation is thus valuable when controversial products are common and not too controversial.
Throughout the paper, we have analyzed the value of the recommendation system from a consumer perspective. Arguably, online recommendation systems employed on an e-commerce platform are more likely to be geared towards the interests of the platform. Depending on the details of the fee system that it employs, the platform could, for instance, be interested in maximizing the probability of purchases. It is possible to look at such questions in the context of our model, too. A naïve conjecture would be that it is then optimal to set as low a recommendation threshold as possible so that all products except the bad one yield buy recommendations. However, recall that it is not necessarily the case that all receivers accept recommendations. The analysis of this problem is then non-trivial. [^34] In the case of a symmetric population, in which all recommendations are accepted, the problem is simpler, and a sales-maximizing recommendation system indeed sets as low a threshold as possible. Moreover, our results show that this may even be optimal for consumers. This latter part stands in contrast to, for instance, the welfare-reducing inflated ratings in Peitz and Sobolev (forthcoming).
Finally, we have maintained a number of classical assumptions such as risk neutrality, linear probability weights, and Bayesian updating. Preliminary results suggest that a promising avenue for future research is to relax these assumptions so as to study their impact on the receiver’s tendency to accept a recommendation or not and, thus, on the value of the recommendation system.
Appendix A Formal Details and Proofs
A.1 Proofs and Technical Details for the Benchmark Model
The Effects of Recommendations on Beliefs
To understand how beliefs are updated after a recommendation, it is useful to decompose the difference between beliefs with and without a recommendation into several components. We illustrate this for buy recommendations. Let $\textbf{p}(r)=(p_{H}^{r},p_{1}^{r},p_{2}^{r},p_{L}^{r})$ denote the vector of posterior probabilities following a recommendation $r\in\{B,D\}$. The total effect of the buy recommendation $B$ on the probability vectors is the sum of three parts,
$\textbf{p}(B)-\textbf{q}=(\textbf{p}(B)-\textbf{q}^{\prime\prime})+(\textbf{q}^{\prime\prime}-\textbf{q}^{\prime})+(\textbf{q}^{\prime}-\textbf{q})\text{,}$ (15)
where the three components of the right-hand side will be explained in Steps 1 to 3 below.
Step 1
The first effect of a buy recommendation is that the probability of an objectively bad product falls to zero. We isolate this effect by assuming that the odds ratios between the remaining events remain unchanged so that the effect corresponds to a change from the prior probability q to $\textbf{q}^{\prime}:=(\frac{q_{H}}{1-q_{L}},\frac{q_{1}}{1-q_{L}},\frac{q_{2}}{1-q_{L}},0)$. This effect of the buy recommendation states that all three quality profiles other than $(0,0)$ become more likely at the expense of quality profile $(0,0)$.
Step 2
Next, the buy recommendation means that the objectively good product becomes more likely relative to the controversial products. We isolate this effect by assuming that (i) the probability of $(1,1)$ increases to $p_{H}$ and (ii) the odds ratios between the remaining events remain unchanged so that the effect corresponds to a change from ${\textbf{q}}^{\prime}:=(\frac{q_{H}}{1-q_{L}},\frac{q_{1}}{1-q_{L}},\frac{q_{2}}{1-q_{L}},0)$ to ${\textbf{q}}^{\prime\prime}:=(p_{H},k\frac{q_{1}}{1-q_{L}},k\frac{q_{2}}{1-q_{L}},0)$, where $k=\frac{(1-p_{H})(1-q_{L})}{q_{1}+q_{2}}$ to guarantee that ${\textbf{q}}^{\prime\prime}$ is a probability vector. Clearly, $p_{H}>\frac{q_{H}}{1-q_{L}}$.
Step 3
Finally, the buy recommendation may change the odds between the controversial product versions. We isolate this effect as a change in the probability vector from ${\textbf{q}}^{\prime\prime}:=(p_{H},k\frac{q_{1}}{1-q_{L}},k\frac{q_{2}}{1-q_{L}},0)$ to $\textbf{p}(B)$.
Proof of Proposition 1
Part (i): Accepting the don’t-buy recommendation is optimal if and only if $U(0)-U(D)\geq 0$ or, equivalently,
$q_{H}+q_{1}(1/2+i)+q_{2}(1/2-i)\geq p_{1}^{D}(1/2+i)+p_{2}^{D}(1/2-i)\text{.}$ (16)
Similarly, accepting the buy recommendation is optimal if and only if $U(B)-U(0)\geq 0$ or, equivalently,
$p_{H}^{B}+p_{1}^{B}(1/2+i)+p_{2}^{B}(1/2-i)\geq q_{H}+q_{1}(1/2+i)+q_{2}(1/2-i)\text{.}$ (17)
We shall show that conditions (16) and (17) are equivalent by proving the following two steps.
Step 1: (16) is equivalent with
$1-p_{1}^{B}-p_{2}^{D}\geq(1/2+i)[p_{2}^{B}-p_{2}^{D}+p_{1}^{D}-p_{1}^{B}]\text{.}$ (18)
Step 2: (17) is equivalent with (18).
The proofs of both steps will rely on the identities
$q_{s}=\pi^{B}p_{s}^{B}+\pi^{D}p_{s}^{D}=(1-\pi^{D})p_{s}^{B}+\pi^{D}p_{s}^{D}\text{ for }s\in\{H,1,2\}$ (19)
and
$p_{H}^{B}=1-p_{1}^{B}-p_{2}^{B}\text{,}$ (20)
Proof of Step 1: Replacing $q_{H}$, $q_{1}$ and $q_{2}$ in (16) using (19), we obtain
$\left(1-\pi^{D}\right)p_{H}^{B}+\left(\left(1-\pi^{D}\right)p_{1}^{B}+\pi^{D}p_{1}^{D}\right)\left(\frac{1}{2}+i\right)$
$+\left(\left(1-\pi^{D}\right)p_{2}^{B}+\pi^{D}p_{2}^{D}\right)\left(\frac{1}{2}-i\right)\geq p_{1}^{D}\left(\frac{1}{2}+i\right)+p_{2}^{D}\left(\frac{1}{2}-i\right)\text{.}$
This simplifies to
$p_{H}^{B}+p_{1}^{B}\left(\frac{1}{2}+i\right)+p_{2}^{B}\left(\frac{1}{2}-i\right)\geq\left(p_{1}^{D}\left(\frac{1}{2}+i\right)+p_{2}^{D}\left(\frac{1}{2}-i\right)\right)\text{.}$
Using (20) and then rearranging further gives
$1-p_{1}^{B}-p_{2}^{B}+p_{1}^{B}\left(\frac{1}{2}+i\right)+p_{2}^{B}\left(\frac{1}{2}-i\right)\geq p_{1}^{D}\left(\frac{1}{2}+i\right)+p_{2}^{D}\left(\frac{1}{2}-i\right)\Leftrightarrow$
$1-p_{1}^{B}-p_{2}^{D}-p_{2}^{B}+p_{1}^{B}\left(\frac{1}{2}+i\right)+p_{2}^{B}\left(\frac{1}{2}-i\right)\geq p_{1}^{D}\left(\frac{1}{2}+i\right)-p_{2}^{D}\left(\frac{1}{2}+i\right)\Leftrightarrow$
$1-p_{1}^{B}-p_{2}^{D}+p_{1}^{B}\left(\frac{1}{2}+i\right)+p_{2}^{B}\left(-\frac{1}{2}-i\right)\geq\left(p_{1}^{D}-p_{2}^{D}\right)\left(\frac{1}{2}+i\right)\Leftrightarrow$
$1-p_{1}^{B}-p_{2}^{D}\geq(1/2+i)[p_{2}^{B}-p_{2}^{D}+p_{1}^{D}-p_{1}^{B}]\text{.}$
Proof of Step 2: Replacing $q_{H}$ and $q_{1}$ and $q_{2}$ in (17) using (19) gives
$p_{H}^{B}+p_{1}^{B}\left(\frac{1}{2}+i\right)+p_{2}^{B}\left(\frac{1}{2}-i\right)\geq$
$\left(1-\pi^{D}\right)p_{H}^{B}+\left(\left(1-\pi^{D}\right)p_{1}^{B}+\pi^{D}p_{1}^{D}\right)\left(\frac{1}{2}+i\right)+\left(\left(1-\pi^{D}\right)p_{2}^{B}+\pi^{D}p_{2}^{D}\right)\left(\frac{1}{2}-i\right)\text{.}$
Rearranging this expression gives
$p_{H}^{B}+p_{1}^{B}\left(\frac{1}{2}+i\right)+p_{2}^{B}\left(\frac{1}{2}-k\right)-p_{1}^{D}\left(\frac{1}{2}+i\right)-p_{2}^{D}\left(\frac{1}{2}-i\right)\geq 0$
Adding $-p_{1}^{B}-p_{2}^{D}+p_{2}^{D}+p_{1}^{B}$ on the left-hand side gives:
$p_{H}^{B}-p_{1}^{B}-p_{2}^{D}+p_{2}^{D}+p_{1}^{B}+p_{1}^{B}\left(\frac{1}{2}+i\right)+p_{2}^{B}\left(\frac{1}{2}-i\right)-p_{1}^{D}\left(\frac{1}{2}+i\right)-p_{2}^{D}\left(\frac{1}{2}-i\right)\geq 0$
and thus
$p_{H}^{B}-p_{1}^{B}-p_{2}^{D}\geq-p_{2}^{D}-p_{1}^{B}-p_{1}^{B}\left(\frac{1}{2}+i\right)-p_{2}^{B}\left(\frac{1}{2}-i\right)+p_{1}^{D}\left(\frac{1}{2}+i\right)+p_{2}^{D}\left(\frac{1}{2}-i\right)\text{.}$
Using (20) gives
$1-p_{1}^{B}-p_{2}^{B}-p_{1}^{B}-p_{2}^{D}$
$\geq-p_{2}^{D}-p_{1}^{B}-p_{1}^{B}\left(\frac{1}{2}+i\right)-p_{2}^{B}\left(\frac{1}{2}-i\right)+p_{1}^{D}\left(\frac{1}{2}+i\right)+p_{2}^{D}\left(\frac{1}{2}-i\right)$
or equivalently
$1-p_{1}^{B}-p_{2}^{B}\geq-p_{1}^{B}\left(\frac{1}{2}+i\right)-p_{2}^{B}\left(\frac{1}{2}-i\right)+p_{1}^{D}\left(\frac{1}{2}+i\right)+p_{2}^{D}\left(\frac{1}{2}-i\right)$
Adding $-p_{2}^{D}+p_{2}^{B}$ on both sides gives:
$1-p_{1}^{B}-p_{2}^{D}-p_{2}^{B}+p_{2}^{B}$
$\geq-p_{1}^{B}\left(\frac{1}{2}+i\right)-p_{2}^{B}\left(\frac{1}{2}-i\right)+p_{1}^{D}\left(\frac{1}{2}+i\right)+p_{2}^{D}\left(\frac{1}{2}-i\right)-p_{2}^{D}+p_{2}^{B}\text{,}$
which is equivalent with (18).
Part (ii): This follows directly from plugging in $\Delta^{r}_{O}$ and $\Delta^{r}_{S}$ into equation (5) for the buy recommendation the analogous equation for a don’t-buy recommendation.
Proof of Proposition 2
Recall from equation (5) the optimal acceptance condition
$(p_{H}^{B}-q_{H})+\frac{p_{1}^{B}-q_{1}}{2}+\frac{p_{2}^{B}-q_{2}}{2}\geq i[(p_{2}^{B}-q_{2})-(p_{1}^{B}-q_{1})].$
First, observe that
$\lim_{R\rightarrow 1}p_{1}^{B}=\lim_{R\rightarrow 1}p_{2}^{B}=0$
and thus $\lim_{R\rightarrow 1}p_{H}^{B}=1$. Hence, in the case $R\rightarrow 1$ the optimality condition simplifies to
$1-q_{H}-\frac{q_{1}+q_{2}}{2}\geq i(q_{1}-q_{2}).$ (21)
Equation (21) always holds for $i\geq 0$. Further, for $q_{1}>q_{2}$ the condition is most stringent for $i=1/2$. In this case, it is equivalent to $1\geq q_{1}+q_{H}$, which always holds. The case $i<0$ is analogous. In the case $R\rightarrow 0$ we get
$\lim_{R\rightarrow 0}p_{s}(B)=\frac{q_{s}}{1-q_{L}}$
for $s\in\{H,1,2\}$ and thus the optimality condition simplifies to
$\frac{q_{H}q_{L}}{1-q_{L}}+\frac{q_{1}q_{L}}{2(1-q_{L})}+\frac{q_{2}q_{L}}{2(1-q_{L})}\geq i[\frac{q_{2}q_{L}}{1-q_{L}}-\frac{q_{1}q_{L}}{1-q_{L}}]$
$\Leftrightarrow 2q_{H}+q_{1}+q_{2}\geq 2i[q_{2}-q_{1}].$
For $i\geq 0$, the last condition always holds if $q_{1}\geq q_{2}$. Further, for $q_{2}>q_{1}$ the condition is most stringent for $i=1/2$. In this case, it is equivalent to $q_{H}+q_{1}\geq 0$, which is always satisfied. The case $i<0$ is analogous.
Proof of Proposition 3
(i) If $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$, inserting (10) and (11) yields
[…]
$=\pi^{B}\int_{-1/2}^{1/2}\bigl{(}p_{H}^{B}+(1/2+i)p_{1}^{B}+(1/2-i)p_{2}^{B}\bigr{)}-(q_{H}+(1/2+i)q_{1}+(1/2-i)q_{2})dF(i)$
[…]
$=\pi^{B}[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i]]$
(ii) If $\Delta_{S}^{B}<-2\Delta_{O}^{B}$, inserting (10), (11) and (12) yields
[…]
$=\int_{-1/2}^{\tilde{i}}\pi^{B}\left(q_{H}+q_{1}(\frac{1}{2}+i)+q_{2}(\frac{1}{2}-i)\right)+(1-\pi^{B})\left(p_{H}^{D}+p_{1}^{D}(\frac{1}{2}+i)+p_{2}^{D}(\frac{1}{2}-i)\right)dF(i)$
$+\int_{\tilde{i}}^{1/2}\pi^{B}\left(p_{H}^{B}+p_{1}^{B}(\frac{1}{2}+i)+p_{2}^{B}(\frac{1}{2}-i)\right)+(1-\pi^{B})\left(q_{H}+q_{1}(\frac{1}{2}+i)+q_{2}(\frac{1}{2}-i)\right)dF(i)$
$-\int_{-1/2}^{1/2}q_{H}+q_{1}(\frac{1}{2}+i)+q_{2}(\frac{1}{2}-i)dF(i)$
[…]
$=\int_{-1/2}^{\tilde{i}}\pi^{B}\left(q_{H}+q_{1}(\frac{1}{2}+i)+q_{2}(\frac{1}{2}-i)\right)+(1-\pi^{B})\left(p_{H}^{D}+p_{1}^{D}(\frac{1}{2}+i)+p_{2}^{D}(\frac{1}{2}-i)\right)dF(i)$
$+\int_{\tilde{i}}^{1/2}\pi^{B}\left(p_{H}^{B}+p_{1}^{B}(\frac{1}{2}+i)+p_{2}^{B}(\frac{1}{2}-i)\right)+(1-\pi^{B})\left(q_{H}+q_{1}(\frac{1}{2}+i)+q_{2}(\frac{1}{2}-i)\right)dF(i)$
$-\int_{-1/2}^{\tilde{i}}q_{H}+q_{1}(\frac{1}{2}+i)+q_{2}(\frac{1}{2}-i)dF(i)-\int_{\tilde{i}}^{1/2}q_{H}+q_{1}(\frac{1}{2}+i)+q_{2}(\frac{1}{2}-i)dF(i)$
$=(1-\pi^{B})\int_{-1/2}^{\tilde{i}}\underbrace{(p_{H}^{D}-q_{H})+\frac{p_{1}^{D}-q_{1}}{2}+\frac{p_{2}^{D}-q_{2}}{2}}_{=\Delta_{O}^{D}}-i(\underbrace{[(p_{2}^{D}-q_{2})-(p_{1}^{D}-q_{1})]}_{=\Delta_{S}^{D}})dF(i)$
$+\pi^{B}\int_{\tilde{i}}^{1/2}\underbrace{(p_{H}^{B}-q_{H})+\frac{p_{1}^{B}-q_{1}}{2}+\frac{p_{2}^{B}-q_{2}}{2}}_{=\Delta_{O}^{B}}-i(\underbrace{[(p_{2}^{B}-q_{2})-(p_{1}^{B}-q_{1})]}_{=\Delta_{S}^{B}})dF(i)$
$=(1-\pi^{B})F(\tilde{i})\left[\Delta_{O}^{D}-\Delta_{S}^{D}\mathbb{E}[i\mid i\leq\tilde{i}]\right]+\pi^{B}(1-F(\tilde{i}))\left[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i\mid i\geq\tilde{i}]\right].$
[…]
Proof of Proposition 4
According to Corollary 1, all types accept the recommendation if
$\min\{\Delta_{O}^{B}+\frac{\Delta_{S}^{B}}{2},\Delta_{O}^{B}-\frac{\Delta_{S}^{B}}{2}\}\geq 0.$ (22)
To see that this condition is fulfilled, first observe that
$\Delta_{O}^{B}+\frac{\Delta_{S}^{B}}{2}$ $=p_{H}^{B}-q_{H}+p_{2}^{B}-q_{2}$
$=\frac{q_{H}}{q_{H}+\beta(q_{1}+q_{2})}-q_{H}+\frac{\beta q_{2}}{q_{H}+\beta(q_{1}+q_{2})}-q_{2}$
$=\frac{q_{H}+\beta q_{2}-(q_{H}+q_{2})(q_{H}+\beta(q_{1}+q_{2}))}{q_{H}+\beta(q_{1}+q_{2})}$
$=\frac{\beta(q_{2}-(q_{H}+q_{2})(q_{1}+q_{2}))+q_{H}(1-q_{H}-q_{2})}{q_{H}+\beta(q_{1}+q_{2})}$
This term is positive if $q_{2}-(q_{H}+q_{2})(q_{1}+q_{2})\geq 0$. If $q_{2}-(q_{H}+q_{2})(q_{1}+q_{2})<0$, then
$\frac{\beta(q_{2}-(q_{H}+q_{2})(q_{1}+q_{2}))+q_{H}(1-q_{H}-q_{2})}{q_{H}+\beta(q_{1}+q_{2})}$
$\geq\frac{(q_{2}-(q_{H}+q_{2})(q_{1}+q_{2}))+q_{H}(1-q_{H}-q_{2})}{q_{H}+\beta(q_{1}+q_{2})}=\frac{(q_{H}+q_{2})(1-q_{1}-q_{2}-q_{H})}{q_{H}+\beta(q_{1}+q_{2})}\geq 0\text{.}$
Thus, $\Delta_{O}^{B}+\frac{\Delta_{S}^{B}}{2}\geq 0$ even in this case. Proceeding analogously, one obtains $\Delta_{O}^{B}-\frac{\Delta_{S}^{B}}{2}=p_{H}^{B}-q_{H}+p_{1}^{B}-q_{1}\geq 0.$ Corollary 1 thus implies that all types accept the recommendation. Assumption 1 implies $\mathbb{E}[i]=0$, so that the result follows from Proposition 3.
Proof of Lemma 1
The result on $\pi^{B}$ follows from inserting $q_{H}=(1-2Q)\sigma/(1+\sigma)$, $\phi_{1}(R)=\phi_{2}(R)=\beta$ and $q_{1}+q_{2}=2Q$ into $\pi^{B}=q_{H}+q_{1}\phi_{1}(R)+q_{2}\phi_{2}(R)$. Next, equation (6) gives
$\Delta_{O}^{B}=\frac{\beta Q(\sigma+1)-2Q\sigma+\sigma}{2\beta Q(\sigma+1)-2Q\sigma+\sigma}-\frac{(1-2Q)\sigma}{1+\sigma}-Q$ (23)
The result on $V(\beta)$ then follows immediately from Proposition 4.
Proof of Lemma 2
(i) Note that $\pi^{B}=(1-2Q)\sigma/(1+\sigma)+2Q\beta$, so that $\pi_{B}$ is strictly increasing. Moreover,
$\frac{\partial{\Delta_{O}^{B}}}{\partial{\beta}}=-\frac{(Q(1-2Q)\sigma(1+\sigma))}{(\sigma-2Q\sigma+2\beta Q(1+\sigma))^{2}}<0.$
(ii) Lemma 1 implies
$V^{\prime}(\beta)=\frac{Q(1-2Q)(1-\sigma)}{1+\sigma},$
and hence
$\frac{\partial{{}^{2}V}}{\partial{\beta\partial\sigma}}=-\frac{2Q(1-2Q)}{(1+\sigma)^{2}}<0,$
$V^{\prime}(\beta)$ is negative for $\sigma>1$, positive for $\sigma<1$ and zero for $\sigma=1$.
Proof of Corollary 2
Using (7), applying the $q_{H}$ notation for ease of exposition and imposing Assumption 2 and $q_{1}=q_{2}$ we obtain
$\Delta_{S}^{B}=-\frac{Q(1-R^{a}-(1-R)^{a})}{q_{H}+Q(1-R^{a}+(1-R)^{a})}.$ (24)
For $a=1$, this term simplifies to $\Delta_{S}^{B}=0$. Hence, it suffices to show that $\Delta_{S}^{B}$ is strictly increasing in $a$. To this end, note that the derivative of $\Delta_{S}^{B}$ with respect to $a$ is given by
$\frac{\partial{\Delta_{S}^{B}}}{\partial{a}}$ $=\frac{Q\left[\log(R)R^{a}\left(2Q(1-R)^{a}+q_{H}\right)+\log(1-R)(1-R)^{a}\left(2Q\left(1-R^{a}\right)+q_{H}\right)\right]}{\left(Q\left(1-R^{a}+(1-R)^{a}\right)+q_{H}\right)^{2}}$
Since $R\in(0,1)$ this expression is negative. Further, given the assumptions, the indifferent type can be expressed as (for $a\neq 1$)
$\tilde{i}$ $=\frac{2Q^{2}\left(1-R^{a}+(1-R)^{a}\right)-2(1-q_{H})q_{H}}{2Q\left(1-R^{a}-(1-R)^{a}\right)}$
$-\frac{Q\left(1-R^{a}+(1-R)^{a}-2q_{H}\left(2-R^{a}+(1-R)^{a}\right)\right)}{2Q\left(1-R^{a}-(1-R)^{a}\right)}$
where
$\frac{\partial{\tilde{i}}}{\partial{Q}}=\frac{Q^{2}\left(1-R^{a}+(1-R)^{a}\right)+(1-q_{H})q_{H}}{Q^{2}\left(1-R^{a}-(1-R)^{a}\right)}$
so that $\tilde{i}$ is increasing in $Q$ when $a>1$ and decreasing in $Q$ when $a<1$.
Proof of Proposition 6
[…]
$V(R)=\begin{cases}\pi^{B}\left[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i]\right]&\text{if }\tilde{i}>1/2\\ \pi^{B}F(\tilde{i})\left[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i\mid i\leq\tilde{i}]\right]&\text{if }\tilde{i}\leq 1/2\\ +(1-\pi^{B})(1-F(\tilde{i}))\left[\Delta_{O}^{D}-\Delta_{S}^{D}\mathbb{E}[i\mid i\geq\tilde{i}]\right]&\\ \end{cases}$
Proposition 2 implies $\tilde{i}>1/2$ if $R\rightarrow 0$ or $R\rightarrow 1$. Thus, it suffices to consider
$V(R)=$ $(q_{H}+Q(1-R^{a}+(1-R)^{a}))\left(\frac{2q_{H}+Q(1-R^{a}+(1-R)^{a})}{2(q_{H}+Q(1-R^{a}+(1-R)^{a}))}-q_{H}-Q\right)$
$+(q_{H}+Q(1-R^{a}+(1-R)^{a}))\frac{Q(1-R^{a}-(1-R)^{a})}{q_{H}+Q(1-R^{a}+(1-R)^{a})}\frac{a-1}{2(a+1)}$
$=\left(\frac{2q_{H}+Q(1-R^{a}+(1-R)^{a})}{2}-(q_{H}+Q(1-R^{a}+(1-R)^{a}))(q_{H}+Q)\right)$
$+Q(1-R^{a}-(1-R)^{a})\frac{a-1}{2(a+1)}$
for $R$ sufficiently close to 0 and 1. Using $q_{H}=\frac{(1-2Q)\sigma}{1+\sigma}$ and rearranging gives
$V(R)$ $=\frac{(1-2Q)\sigma}{1+\sigma}+\frac{aQ(1-R^{a})+Q(1-R)^{a}}{a+1}$
$-\left(\frac{(1-2Q)\sigma}{1+\sigma}+Q(1-R^{a})+Q(1-R)^{a}\right)\left(\frac{(1-2Q)\sigma}{1+\sigma}+Q\right)$
We then obtain the corresponding derivative
$V^{\prime}(R)=\frac{aQ\left((a+1)\left(R^{a-1}+(1-R)^{a-1}\right)(Q+\frac{(1-2Q)\sigma}{1+\sigma})-aR^{a-1}-(1-R)^{a-1}\right)}{a+1}$
We can then write
$V^{\prime}(R)>0$
$\Leftrightarrow(a+1)\left(R^{a-1}+(1-R)^{a-1}\right)(Q+\frac{(1-2Q)\sigma}{1+\sigma})-aR^{a-1}-(1-R)^{a-1}>0$
$\Leftrightarrow(a+1)\left(Q+\frac{(1-2Q)\sigma}{1+\sigma}\right)>\frac{aR^{a-1}+(1-R)^{a-1}}{R^{a-1}+(1-R)^{a-1}}=\frac{a(1-R)^{1-a}+R^{1-a}}{(1-R)^{1-a}+R^{1-a}},$
so that
$\lim_{R\rightarrow 0}V^{\prime}(R)>0\Leftrightarrow(a+1)\left(Q+\frac{(1-2Q)\sigma}{1+\sigma}\right)>a$
and analogously
$\lim_{R\rightarrow 1}V^{\prime}(R)<0\Leftrightarrow(a+1)\left(Q+\frac{(1-2Q)\sigma}{1+\sigma}\right)<1.$
Thus, a sufficient condition for an interior solution when $a<1$ is given by
$\frac{1}{a+1}>\left(Q+\frac{(1-2Q)\sigma}{1+\sigma}\right)>\frac{a}{a+1}.$ (25)
Rearranging the first inequality, we obtain
$\left(\frac{1}{a+1}-Q\right)(1+\sigma)>(1-2Q)\sigma$
$\Leftrightarrow\frac{1}{a+1}-Q>(1-2Q)\sigma-\left(\frac{1}{a+1}-Q\right)\sigma$
$\Leftrightarrow\frac{1}{a+1}-Q>\sigma\left(\frac{a}{1+a}-Q\right)$
For $a<1$, the condition $Q\geq 1-\max\{\frac{a}{a+1},\frac{1}{a+1}\}$ is equivalent to $Q\geq\frac{a}{1+a}$ so that the right-hand side is negative while the left-hand side is positive. Hence, this inequality is always satisfied. Proceeding analogously for the second inequality, we obtain
$\left(\frac{a}{a+1}-Q\right)(1+\sigma)<(1-2Q)\sigma,$
which is also implied by $Q\geq 1-\max\{\frac{a}{a+1},\frac{1}{a+1}\}$. Hence, the first statement of Proposition 6 for the case $a<1$ follows. For the second statement when $Q<1-\max\{\frac{a}{a+1},\frac{1}{a+1}\}$, we can rewrite the two inequalities in (25) as
$\frac{1-Q(a+1)}{a-Q(a+1)}>\sigma>\frac{a-Q(a+1)}{1-Q(a+1)}$
giving the second condition in the statement.
Proof of Proposition 7
Using (23) and (24), we obtain
$\lim_{Q\rightarrow 0}\Delta_{S}^{B}$ $=\lim_{Q\rightarrow 0}-\frac{Q(1-R^{a}-(1-R)^{a})}{q_{H}+Q(1-R^{a}+(1-R)^{a})}=0,$
and
$\lim_{Q\rightarrow 0}\Delta_{O}^{B}$ $=\lim_{Q\rightarrow 0}\frac{2q_{H}+Q(1-R^{a}+(1-R)^{a})}{2(q_{H}+Q(1-R^{a}+(1-R)^{a}))}-q_{H}-Q=1-q_{H}>0.$
Therefore, $2\Delta_{O}^{B}\geq|\Delta_{S}^{B}|$ so that all types accept a recommendation according to Corollary 1. Thus, Proposition 3 implies
$V(R)$ $=\pi^{B}[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i]]$
Rearranging this equation gives
$V(R)$ $=\frac{(1-2Q)\sigma}{1+\sigma}+\frac{aQ(1-R^{a})+Q(1-R)^{a}}{a+1}$
$-\left(\frac{(1-2Q)\sigma}{1+\sigma}+Q(1-R^{a})+Q(1-R)^{a}\right)\left(\frac{(1-2Q)\sigma}{1+\sigma}+Q\right)$
$=\frac{(1-2Q)\sigma}{1+\sigma}\left(1-\left(\frac{(1-2Q)\sigma}{1+\sigma}+Q\right)\right)$
$+\frac{Q}{a+1}(a(1-R^{a})+(1-R)^{a})$
$-\frac{Q}{a+1}\left((1-R^{a})+(1-R)^{a}\right)(a+1)\left(\frac{(1-2Q)\sigma}{1+\sigma}+Q\right)$
$=\frac{(1-2Q)\sigma(Q(\sigma-1)+1)}{(\sigma+1)^{2}}$
$+\frac{Q}{a+1}(1-R^{a})\left(a-(a+1)\left(\frac{(1-2Q)\sigma}{1+\sigma}+Q\right)\right)$
$+\frac{Q}{a+1}(1-R)^{a}\left(1-(a+1)\left(\frac{(1-2Q)\sigma}{1+\sigma}+Q\right)\right)$
$=c_{0}+\frac{Q}{a+1}\left[c_{1}(a)(1-R^{a})+c_{2}(a)(1-R)^{a}\right]$
with $c_{0}=\frac{(1-2Q)\sigma(Q(\sigma-1)+1)}{(\sigma+1)^{2}}$, $c_{1}(a)=a+(a+1)\frac{Q(\sigma-1)-\sigma}{\sigma+1}$ and $c_{2}(a)=1+(1+a)\frac{Q(\sigma-1)-\sigma}{\sigma+1}$. We thus obtain
$V^{\prime}(R)$ $=-\frac{Qa}{a+1}\left[R^{a-1}c_{1}(a)+(1-R)^{a-1}c_{2}(a)\right],$
$V^{\prime\prime}(R)$ $=-\frac{Qa(a-1)}{a+1}\left[R^{a-2}c_{1}(a)-(1-R)^{a-2}c_{2}(a)\right],$
For $c_{1}(a),c_{2}(a)\geq 0$ and $c_{1}(a),c_{2}(a)\leq 0$ the function $V(R)$ is monotonically decreasing and increasing, respectively. Further, for $a\geq 1$, $c_{1}(a)>0$ and $c_{2}(a)<0$ it is concave. Conversely, for $a<1$ and $c_{1}(a)<0$ and $c_{2}(a)>0$ it is concave. Further,
$c_{1}(a)-c_{2}(a)\begin{cases}<0&a<1\\ =0&a=1\\ >0&a>1,\end{cases}$
Therefore, for $a<1$ either (i) $c_{1}(a),c_{2}(a)\leq 0$, (ii) $c_{1}(a)\leq 0\leq c_{2}(a)$, or (iii) $0\leq c_{1}(a),c_{2}(a)$. Hence, in all three cases $V(R)$ is quasiconcave. Similarly, for $a>1$ either (i) $c_{1}(a),c_{2}(a)\leq 0$, (ii) $c_{2}(a)\leq 0\leq c_{1}(a)$, or (iii) $0\leq c_{1}(a),c_{2}(a)$. Hence, in all three cases $V(R)$ is quasiconcave. Finally, for $a=1$ quasiconcavity follows because $c_{1}(a)=c_{2}(a)$ and thus $V$ is monotone. Putting all of this together, it follows that $V$ is quasiconcave in $R$. Hence, the sufficient conditions in Proposition 6 for an interior solution are also necessary.
Proof of Proposition 8
As argued in the text, a mean-preserving spread corresponds to an increase in $\beta$ for $R>1/2$, while for $R<1/2$ it corresponds to a decrease in $\beta$. From the proof of Lemma 2, we can see that the value of the recommendation system is increasing in $\beta$ for $\sigma<1$ and decreasing in $\beta$ for $\sigma>1$. Thus, the statement follows.
Proof of Corollary 3
We first take the derivative of $V$ with respect to $Q$ to obtain
$\frac{\partial V}{\partial Q}=\frac{1}{\left(\sigma+1\right)^{2}}\left(\beta-3\sigma+4Q\sigma-4Q\beta+\sigma^{2}-4Q\sigma^{2}-\sigma^{2}\beta+4Q\sigma^{2}\beta\right)$
Setting this equal to zero and solving for the candidate solution we obtain
$Q^{\ast}=\frac{3\sigma-\beta-\sigma^{2}+\sigma^{2}\beta}{4\sigma-4\beta-4\sigma^{2}+4\sigma^{2}\beta}$
Next, the SOC reads
$4\sigma-4\beta-4\sigma^{2}+4\sigma^{2}\beta<0,$
where the l.h.s. of the SOC is the denominator of $Q^{\ast}$. Thus, a necessary condition for a positive interior solution is
$3\sigma-\beta-\sigma^{2}+\sigma^{2}\beta<0\text{.}$
This proves (i). To prove (ii), note that $03\sigma-\beta-\sigma^{2}+\sigma^{2}\beta<$ 0 implies $04\sigma-4\beta-4\sigma^{2}+4\sigma^{2}\beta<$ 0. Hence, the candidate solution is positive. It remains to show that $Q^{\ast}<\frac{1}{2}$. Recall that, by the requirements that the SOC holds and $Q^{\ast}>0$, we can focus on the case that both the numerator and the denominator in $Q^{\ast}$ are negative. Accordingly, the requirement that $\frac{3\sigma-\beta-\sigma^{2}+\sigma^{2}\beta}{4\sigma-4\beta-4\sigma^{2}+4\sigma^{2}\beta}<1/2$ becomes
$2\sigma-2\beta-2\sigma^{2}+2\sigma^{2}\beta<3\sigma-\beta-\sigma^{2}+\sigma^{2}\beta.$
Equivalently,
$\left(\sigma+1\right)\left(\sigma(1-\beta)+\beta\right)>0$
which always holds for $\beta\leq 1$.
A.2 Distinct Sender and Receiver Distributions
Proof of Proposition 9
The value of the recommendation system can be written as
[…]
and the result from Proposition 3 translates accordingly so that we can rewrite $V(\mathcal{R})$ as
$\pi^{B}[\Delta_{O,F}^{B}-\Delta_{S,F}^{B}\mathbb{E}_{G}[i]]$ if |$\text{if }|\Delta_{S,F}^{B}|\leq 2\Delta_{O,F}^{B}$
$(1-\pi^{B})G(\tilde{i})\left[\Delta_{O}^{D}-\Delta_{S,F}^{D}\mathbb{E}_{G}[i\mid i\leq\tilde{i}]\right]+\pi^{B}(1-G(\tilde{i}))\left[\Delta_{O,F}^{B}-\Delta_{S,F}^{B}\mathbb{E}_{G}[i\mid i\geq\tilde{i}]\right]$ if $\text{if }\Delta_{S,F}^{B}<-2\Delta_{O,F}^{B}$
$\pi^{B}G(\tilde{i})\left[\Delta_{O,F}^{B}-\Delta_{S,F}^{B}\mathbb{E}_{G}[i\mid i\leq\tilde{i}]\right]+(1-\pi^{B})(1-G(\tilde{i}))\left[\Delta_{O}^{D}-\Delta_{S,F}^{D}\mathbb{E}_{G}[i\mid i\geq\tilde{i}]\right]$ if $\text{if }\Delta_{S,F}^{B}>2\Delta_{O,F}^{B}.$
Since $F$ satisfies Assumption 1, all receivers will accept the recommendation so that we are in the case $|\Delta_{S,F}^{B}|\leq 2\Delta_{O,F}^{B}$. Thus, we have
$V(\beta)$ $=\pi^{B}[\Delta_{O,F}^{B}-\Delta_{S,F}^{B}\mathbb{E}_{G}[i]]$
$=q_{H}+\beta Q-(q_{H}+2\beta Q)(q_{H}+Q)+(q_{2}-q_{1})(q_{H}-\beta(1-2Q))\mathbb{E}_{G}[i]$
and therefore
$V^{\prime}(\beta)=Q-2Q(q_{H}+Q)+(q_{2}-q_{1})(1-2Q)\mathbb{E}_{G}[i]$
Replacing $q_{H}=((1-2Q)\sigma)/(\sigma+1)$ and rearranging the inequalities $V^{\prime}(\beta)>0$ and $V^{\prime}(\beta)<0$ yields the statement in the proposition.
A.3 Multiple Recommendation Levels
Proof of Proposition 10
The posteriors are given as follows. Following a buy recommendation,
$p_{H}^{B}(R)$ $=\frac{q_{H}}{q_{H}+q_{1}(1-F(R_{2}-1/2))+q_{2}F(1/2-R_{2})},$
$p_{1}^{B}(R)$ $=\frac{q_{1}(1-F(R_{2}-1/2))}{q_{H}+q_{1}(1-F(R_{2}-1/2))+q_{2}F(1/2-R_{2})},$
$p_{2}^{B}(R)$ $=\frac{q_{2}F(1/2-R_{2})}{q_{H}+q_{1}(1-F(R_{2}-1/2))+q_{2}F(1/2-R_{2})}$
$p_{L}^{B}(R)$ $=0.$
A don’t-buy recommendation results in a posterior of
$p_{H}^{D}(R)$ $=0,$
$p_{1}^{D}(R)$ $=\frac{q_{1}F(R_{1}-1/2)}{q_{1}F(R_{1}-1/2)+q_{2}(1-F(1/2-R_{1}))+q_{L}},$
$p_{2}^{D}(R)$ $=\frac{q_{2}(1-F(1/2-R_{1}))}{q_{1}F(R_{1}-1/2)+q_{2}(1-F(1/2-R_{1}))+q_{L}}$
$p_{L}^{D}(R)$ $=\frac{q_{L}}{q_{1}F(R_{1}-1/2)+q_{2}(1-F(1/2-R_{1}))+q_{L}}.$
Finally, the recommendation neither to buy nor not to buy yields
$p_{H}^{N}(R)$ $=0$ $\quad p_{2}^{N}(R)$ $=\frac{q_{2}\Gamma_{2}}{q_{1}\Gamma_{1}+q_{2}\Gamma_{2}}$
$p_{1}^{N}(R)$ $=\frac{q_{1}\Gamma_{1}}{q_{1}\Gamma_{1}+q_{2}\Gamma_{2}}$ $\quad p_{L}^{N}(R)$ $=0.$
where $\Gamma_{1}\equiv F(R_{2}-1/2)-F(R_{1}-1/2)$ and $\Gamma_{2}\equiv F(1/2-R_{1})-F(1/2-R_{2})$. For the symmetric case of Assumption 1, setting $F(1/2-R_{2})=1-F(R_{2}-1/2)=\beta_{2}$ and $F(1/2-R_{1})=1-F(R_{1}-1/2)=\beta_{1}$, this reduces to
$p_{H}^{B}$ $=\frac{q_{H}}{q_{H}+2Q\beta_{2}},$ $\quad p_{H}^{N}=0,\quad\quad\quad p_{H}^{D}$ $=0,$
$p_{1}^{B}$ $=\frac{q_{1}\beta_{2}}{q_{H}+2Q\beta_{2}},$ $\quad p_{1}^{N}=\frac{q_{1}}{q_{1}+q_{2}},\quad\quad\quad p_{1}^{D}$ $=\frac{q_{1}(1-\beta_{1})}{1-q_{H}-2Q\beta_{1}},$
$p_{2}^{B}$ $=\frac{q_{2}\beta_{2}}{q_{H}+2Q\beta_{2}},$ $\quad p_{2}^{N}=\frac{q_{2}}{q_{1}+q_{2}},\quad\quad\quad p_{2}^{D}$ $=\frac{q_{2}(1-\beta_{1})}{1-q_{H}-2Q\beta_{1}},$
$p_{L}^{B}$ $=0,$ $\quad p_{L}^{N}=0,\quad\quad\quad p_{L}^{D}$ $=\frac{q_{L}}{1-q_{H}-2Q\beta_{1}}.$
Next, we determine the indifferent type following an intermediate recommendation. Type $i$ will buy the product with the intermediate recommendation whenever
$\frac{q_{1}}{q_{1}+q_{2}}(1/2+i)+\frac{q_{2}}{q_{1}+q_{2}}(1/2-i)\geq q_{H}+q_{1}(1/2+i)+q_{2}(1/2-i)$
$\Leftrightarrow$ $i\left(\frac{(q_{1}-q_{2})(q_{H}+q_{L})}{q_{1}+q_{2}}\right)\geq\frac{1}{2}(q_{H}-q_{L})$
If $q_{1}=q_{2}$, all types buy if $q_{L}\geq q_{H}$ and none buy otherwise. If $q_{1}>q_{2}$, all types
$i\geq\frac{1}{2}\frac{(q_{H}-q_{L})(q_{1}+q_{2})}{(q_{1}-q_{2})(q_{H}+q_{L})}$
buy the product if $r=N$ [^36] The argument for $q_{1}<q_{2}$ is analogous. Taken together, we can define
$\tilde{i}_{M}:=\begin{cases}1/2&q_{1}=q_{2},q_{L}<q_{H}\\ -1/2&q_{1}=q_{2},q_{L}\geq q_{H}\\ \frac{1}{2}\frac{(q_{H}-q_{L})(q_{1}+q_{2})}{(q_{1}-q_{2})(q_{H}+q_{L})}&q_{1}\neq q_{2}.\end{cases}$
The value of the recommendation system therefore reads
$V(\beta_{2},\beta_{1})=$ $(q_{H}+2Q\beta_{2})\left(\frac{q_{H}+Q\beta_{2}}{q_{H}+2Q\beta_{2}}-(q_{H}+Q)\right)$
$+2Q(\beta_{1}-\beta_{2})\int_{\tilde{i}_{M}}^{1/2}1/2(q_{L}-q_{H})+i\frac{(1-2Q)(q_{1}-q_{2})}{2Q}dF(i),$
where the first part corresponds to the value of accepting the highest recommendation and the second part comes from buying following an intermediate recommendation. Depending on parameters, $\tilde{i}_{M}$ changes so that some, all or no types buy following an intermediate recommendation. Taking the respective derivatives we obtain
$\frac{\partial{V}}{\partial{\beta_{2}}}$ $=Q(q_{L}-q_{H})F(\tilde{i}_{M})-(1-2Q)(q_{1}-q_{2})(1-F(\tilde{i}_{M}))\mathbb{E}[i\mid i\geq\tilde{i}_{M}]$
$\frac{\partial{V}}{\partial{\beta_{1}}}$ $=Q(q_{L}-q_{H})(1-F(\tilde{i}_{M}))+(1-2Q)(q_{1}-q_{2})(1-F(\tilde{i}_{M}))\mathbb{E}[i\mid i\geq\tilde{i}_{M}]$
Suppose $\sigma>1$ so that $q_{L}<q_{H}$. Then, $\frac{\partial{V}}{\partial{\beta_{2}}}<0$. Next, consider $\sigma<1$ so that $q_{L}>q_{H}$. Then, $\frac{\partial{V}}{\partial{\beta_{1}}}>0$.
Finally, suppose $q_{1}=q_{2}$. If $q_{L}<q_{H}$, then $\tilde{i}_{M}=1/2$, so that the derivatives read
$\frac{\partial{V}}{\partial{\beta_{2}}}$ $=Q(q_{L}-q_{H})<0$
$\frac{\partial{V}}{\partial{\beta_{1}}}$ $=0\text{.}$
For $q_{L}>q_{H}$, we obtain $\tilde{i}_{M}=-1/2$, so that the derivatives read
$\frac{\partial{V}}{\partial{\beta_{2}}}$ $=0$
$\frac{\partial{V}}{\partial{\beta_{1}}}$ $=Q(q_{L}-q_{H})>0.$
Thus, for $q_{L}<q_{H}$ any $\beta_{1}$ and thus any $R_{1}$ is optimal, whereas for $q_{L}>q_{H}$ any $\beta_{2}$ and thus any $R_{2}$ is optimal.
A.4 Multiple Recommendations
After $b$ buy recommendations and $d$ don’t-buy recommendations, where $b+d>0$, the posterior reads
$p_{H}(b,d)$ =$=\begin{cases}\frac{q_{H}}{q_{H}+q_{1}(\phi_{1}(R))^{b}+q_{2}(\phi_{2}(R))^{b}}&\text{if }d=0\\ 0&\text{if }b=0\\ \end{cases}$
$p_{1}(b,d)$ $=\frac{q_{1}(\phi_{1}(R))^{b}(1-\phi_{1}(R))^{d}}{q_{1}(\phi_{1}(R))^{b}(1-\phi_{1}(R))^{d}+q_{2}(\phi_{2}(R))^{b}(1-\phi_{2}(R))^{d}}$
$p_{2}(b,d)$ $=\frac{q_{2}(\phi_{2}(R))^{b}(1-\phi_{2}(R))^{d}}{q_{1}(\phi_{1}(R))^{b}(1-\phi_{1}(R))^{d}+q_{2}(\phi_{2}(R))^{b}(1-\phi_{2}(R))^{d}}$
$p_{L}(b,d)$ =$=\begin{cases}\frac{q_{L}}{q_{1}(1-\phi_{1}(R))^{d}+q_{2}(1-\phi_{2}(R))^{d}+q_{L}}&\text{if }b=0\\ 0&\text{if }d=0\\ \end{cases}$
Lemma 3.
Consider the model with multiple recommendations. Suppose Assumption 1 holds and consider any threshold $R\in(0,1)$.
  1. (i)
    If the receiver obtains only buy (don’t buy) recommendations, then, as the number of recommendations increases, her posterior belief converges to $p_{H}=1$ ($p_{L}=1$).
  2. (ii)
    If the receiver obtains mixed recommendations, then $p_{1}(b,d)/p_{2}(b,d)=q_{1}/q_{2}$ and $p_{H}(b,d)=p_{L}(b,d)=0$ for any $b,d>0$.
Proof.
(i) Suppose the receiver gets only buy recommendations and that $b\rightarrow\infty$. Then, $\phi_{1}(R)=\phi_{2}(R)\in(0,1)$ if $R\in(0,1)$, so that $\lim_{b\rightarrow\infty}(\phi_{i}(R))^{b}=0$ for $i=1,2$. Hence, $\lim_{b\rightarrow\infty}p_{H}(b,0)=1$. The argument for don’t-buy recommendations is analogous.
(ii) For any $b,d>0$, $p_{H}(b,d)=p_{L}(b,d)=0$ obviously holds. Further, because $\phi_{1}(R)=\phi_{2}(R)$ by Assumption 1,
$p_{i}(b,d)=\frac{q_{i}}{q_{1}+q_{2}}$
for $i=1,2$, yielding statement (ii).
Optimal Receiver Behavior with Infinite Learning
Lemma 4.
Consider the “infinite learning” recommendation system with interior threshold $R\in(0,1)$. Then,
  1. (i)
    any receiver type $i\in[-1/2,1/2]$ buys the product if it is good;
  2. (ii)
    no receiver type $i\in[-1/2,1/2]$ buys the product if it is bad;
  3. (iii)
    if the product is controversial,
    1. (a)
      and $\lambda>1$, all receiver types $i\geq\tilde{i}_{\infty}$ buy the product;
    2. (b)
      and $\lambda<1$, all receiver types $i\leq\tilde{i}_{\infty}$ buy the product.
    3. (c)
      and $\lambda=1$, all receiver types $i\in[-1/2,1/2]$ buy the product if $\sigma\leq 1$ and none buy it if $\sigma>1$.
Proof.
First, Lemma 3 implies that infinite learning fully reveals the good and bad products, whereas the posteriors of the controversial product versions are $p_{i}(b,d)=q_{i}/(q_{1}+q_{2})$ for $b,d>0$ Hence, the expected payoff of buying the recommended good with mixed recommendations reads
$\frac{q_{1}}{q_{1}+q_{2}}(1/2+i)+\frac{q_{2}}{q_{1}+q_{2}}(1/2-i)=\frac{1}{2}+i\frac{q_{1}-q_{2}}{q_{1}+q_{2}}.$
Therefore, receiver $i$ buys the product with mixed recommendations if and only if
$\frac{1}{2}+i\frac{q_{1}-q_{2}}{q_{1}+q_{2}}\geq V_{0}(i)=q_{H}+q_{1}(1/2+i)+q_{2}(1/2-i).$ (26)
The indifferent consumer satisfies
$\frac{1}{2}+\tilde{i}_{\infty}\frac{q_{1}-q_{2}}{q_{1}+q_{2}}=q_{H}+q_{1}(1/2+\tilde{i}_{\infty})+q_{2}(1/2-\tilde{i}_{\infty})\text{ and thus}$
$\tilde{i}_{\infty}=\frac{(q_{1}+q_{2})(q_{1}+q_{2}+2q_{H}-1)}{2(q_{1}-q_{2})(1-(q_{1}+q_{2}))}$
Using $q_{1}=\frac{2Q\lambda}{\lambda+1}$, $q_{2}=\frac{2Q}{\lambda+1}$ and $q_{H}=((1-2Q)\sigma)/(\sigma+1)$ gives
$\tilde{i}_{\infty}=\frac{(\sigma-1)}{2(1+\sigma)}\frac{\lambda+1}{\lambda-1}$
Inserting this into (26) gives the result for part (iii) when $\lambda\neq 1$. For the case $\lambda=1$, the condition to accept the recommendation simplifies to
$\frac{1}{2}\geq q_{H}+Q\Leftrightarrow 1\geq 2q_{H}+(1-q_{H}-q_{L})\Leftrightarrow q_{L}\geq q_{H}\Leftrightarrow\sigma\leq 1.$
Finally, parts (i) and (ii) follow trivially.
Proof of Proposition 11
  1. (i)
    If $\lambda=1$ the value of the infinite learning recommendation system is given by
    $V_{\infty}=q_{H}(1-q_{H}-Q)+2Q\max\left\{\frac{1}{2}-q_{H}-Q,0\right\}.$
    The first term corresponds to the value generated when the product is good, so that the sender always buys it. The second term captures the value generated from a controversial product if it is bought, where the purchasing decision depends on the parameters. Hence, $V_{\infty}$ coincides with the value of the single-recommendation system for the optimal $\beta$, depending on whether $q_{L}$ or $q_{H}$ is bigger.
  2. (ii)
    If $\lambda>1$, only $i\geq\tilde{i}_{\infty}$ buy the product by Lemma 4. Hence, no type buys a controversial product in the infinite-learning recommendation system if $\tilde{i}_{\infty}\geq 1/2$, which is equivalent to $\sigma\geq\lambda$. This inequality implies $\sigma>1$. Hence, in that case all receivers behave in the same way in both the optimal single-recommendation system (Proposition 5) and the infinite-learning system, buying only objectively good products. Conversely, if $\tilde{i}_{\infty}\leq-1/2$ all types buy a controversial product with the infinite-learning recommendation system. This can be rewritten as $\sigma\leq 1/\lambda$, which implies $\sigma<1$, so that all receivers buy all products except those that are objectively bad and therefore behave in the same way in the optimal single-recommendation system and the infinite-learning system.
  3. (iii)
    The proof is analogous to case (ii) and thus omitted.

### v2

@@
 Schmutzler: Department of Economics, University of Zurich, and CEPR, armin.schmutzler@econ.uzh.ch.
-We are grateful for very helpful discussions with Piotr Dworczak, Igor Letina, Marc Möller, Nick Netzer, Volker Nocke, Egor Starkov, Jakub Steiner and seminar participants in Bern, Fribourg, Zurich, at SAET (2023), Swiss Society of Economics of Statistics (2024), and Swiss IO Day (2024).
-(August 21, 2024)
+We are grateful for very helpful discussions with Piotr Dworczak, Johannes Johnen, Igor Letina, Marc Möller, Nick Netzer, Volker Nocke, Egor Starkov, Jakub Steiner and seminar participants in Bern, Fribourg, Zurich, at SAET (2023), Swiss Society of Economics of Statistics (2024), Swiss IO Day (2024) and Verein für Socialpolitik (2024).
+(October 11, 2024)
 This paper investigates the value of recommendations for disseminating economic information, with a focus on frictions resulting from preference heterogeneity.
@@
 And when it does not, how likely is it that the preferences of the sender and receiver are aligned?
-The answers to these questions will determine whether it is in the receiver’s interest to follow the recommendation and how valuable it is.
+The answers to these questions will determine whether it is in the receiver’s interest to accept the recommendation and how valuable it is.
 Our model captures the above three features of recommendations in the simplest possible way.
@@
 Thus, four product versions exist.
-All agents agree that a product is worth choosing if it has high quality in both dimensions and is thus (objectively) good.
+There is a continuum of agent types.
+All types agree that a product is worth choosing if it has high quality in both dimensions and is thus (objectively) good.
 Conversely, they agree that a product is not worth choosing if it has low quality in both dimensions and is thus (objectively) bad.
@@
 An agent prefers a controversial product that has high quality in the dimension that is more relevant for them, and is thus subjectively good, to a product that has high quality only in the less-relevant dimension (thus subjectively bad).
-Moreover, even agents who agree about which controversial product they prefer may differ in the intensity of these preferences.
-This heterogeneity is captured in a continuous type distribution.
+Moreover, even agents who agree about which controversial product they prefer may differ in the intensity of these preferences (as captured by the continuous type distribution).
 We assume that products are experience goods so that each agent only has stochastic knowledge of the quality distribution before consumption: They know the probability of each product version.
@@
 To answer these questions, we need to derive intermediate results of independent interest.
-Most importantly, we investigate under which conditions receivers accept a recommendation.
+Most importantly, we investigate under which conditions receivers accept all recommendations.
 We find that for a Bayesian expected-payoff maximizer, two key quantities measuring the objective and subjective content of the recommendation are decisive for receiver behavior.
 These quantities are simple functions of the model primitives.
-Intuitively, the objective informational content of a recommendation is high if it makes the product sufficiently more likely to be good and less likely to be bad.
+Intuitively, the objective informational content of a recommendation is high if it makes the product sufficiently more likely to be (objectively) good and less likely to be bad.
 The subjective content captures the effect of the recommendation on the relative probability of the two controversial product versions.
 The evaluation of the subjective content thus depends on the receiver’s type.
-When the recommendation bears little or no subjective content, all types accept it, thus buying after a buy recommendation and choosing the outside option (one of the other alternatives) after a don’t-buy recommendation.
-As the subjective content increases, revealing that one of the two controversial product versions becomes more likely, only receivers without strong preferences for the other version accept the recommendation.
+When the recommendation bears little or no subjective content, all types accept all recommendations, thus buying after a buy recommendation and choosing the outside option (one of the other alternatives) after a don’t-buy recommendation.
+As the subjective content increases, revealing that one of the two controversial product versions becomes more likely, only receivers without strong preferences for the other version accept all recommendations.
 When the recommendation threshold $R$ approaches the maximal payoff, a buy recommendation essentially becomes objective, and similarly for a don’t-buy recommendation when $R$ approaches the minimal payoff.
-As receivers are aware of this, they accept the recommendation in each of these polar cases since the subjective component vanishes.
+As receivers are aware of this, they accept all recommendations in each of these polar cases since the subjective component vanishes.
 Based on this analysis, we then study the value of recommendation systems.
 We think of this value as the expected-payoff increase of a randomly chosen receiver resulting from a recommendation given by a randomly chosen sender.
-When the recommendation is objective, or the subjective content is small, the analysis is comparatively simple: As all receivers accept the recommendation, they choose the outside option after a don’t-buy recommendation, just as if they had not received any recommendation at all.
+When the recommendation is objective, or the subjective content is small, the analysis is comparatively simple: As all receivers accept all recommendations, they choose the outside option after a don’t-buy recommendation, just as if they had not received any recommendation at all.
 Therefore, only buy recommendations create value—they do so by inducing the receiver to purchase the product if it generates a higher expected payoff for him than the outside option.
@@
 The value of the recommendation system then has two sources.
-First, receivers who accept buy recommendations benefit because they expect the recommended product to be better than the outside option.
-Second, receivers who do not accept the recommendation benefit from buying the product in spite of a don’t-buy recommendation: These receivers have preferences that are not well aligned with the general population so that the don’t-buy recommendation provides good news about the chances of obtaining a subjectively good product rather than one they like less.
+First, receivers who accept all recommendations benefit because they expect the recommended product to be better than the outside option.
+Second, receivers who never accept recommendations benefit from buying the product in spite of a don’t-buy recommendation: These receivers have preferences that are not well aligned with the general population so that the don’t-buy recommendation provides good news about the chances of obtaining a subjectively good product rather than one they like less.
 Next, we turn to the design of recommendation systems.
@@
 Rather than revealing only whether a product is good or bad, it may be optimal for the sender to give a buy recommendation even if the product is controversial but sufficiently aligned with her preferences.
-Intuitively, giving up the exclusive focus on objective content makes sense when both controversial product versions are common but there is broad agreement on which of them is preferable.
+Intuitively, giving up the exclusive focus on objective content makes sense when both controversial product versions are common and there is broad agreement on which of them is preferable.
 By contrast, extreme recommendation thresholds that only reveal objective content are still desirable if the product is much more likely to be objectively good than objectively bad or vice versa.
@@
 With mild adjustments, the main insights continue to hold when sender and receiver distributions are distinct—a situation that could arise because, for instance, recommendations stem from early adopters who do not represent the population at large.
+This also has non-trivial design implications: Our results on the relation between mean-preserving spreads can be adapted to show how a designer would want to choose the sender distribution, which may be at least partly under his control.
 Second, we allow for more than two types of recommendations.
@@
 Section 3 introduces the model and discusses its key assumptions.
-In Section 4, we describe the conditions under which different receiver types accept recommendations.
+In Section 4, we describe the conditions under which different receiver types accept all recommendations.
 Then, in Section 5 we first characterize the value of a recommendation system before investigating the determinants of the value and the design of value-maximizing systems.
@@
 Products and Payoffs
-We consider a consumer’s choice of a product from a large set of alternatives.
+We consider a consumer’s choice of a product from a set of alternatives $N\in\mathbb{N}$ with $N>1$.
 We assume that each product is fully characterized by a two-dimensional quality vector, where $Q_{d}=1$ and $Q_{d}=0$ capture high and low quality, respectively, in dimension $d\in\{1,2\}$.
@@
 Clearly,
-$\pi^{B}(R)$ $:=q_{H}+q_{1}\phi_{1}(R)+q_{2}\phi_{2}(R),$ (1)
-$\pi^{D}(R)$ $:=q_{1}(1-\phi_{1}(R))+q_{2}(1-\phi_{2}(R))+q_{L},$ (2)
+$\pi^{B}(R)$ $:=q_{H}+q_{1}\phi_{1}(R)+q_{2}\phi_{2}(R),$
+$\pi^{D}(R)$ $:=q_{1}(1-\phi_{1}(R))+q_{2}(1-\phi_{2}(R))+q_{L},$
 where $\pi^{D}(R)=1-\pi^{B}(R)$, and $\phi_{1}(R):=1-F(R-1/2)$ and $\phi_{2}(R):=F(1/2-R)$ denote the probability of a buy recommendation conditional on the product version being $(1,0)$ or $(0,1)$, respectively.
 In words, the probability of receiving a buy recommendation from a randomly chosen sender is given by the sum of the probability that a product is objectively good and the probability that it is subjectively good enough from a randomly chosen sender’s perspective.
-Thus, following a buy recommendation the receiver’s posterior $\mathbf{p}^{B}(R):=(p_{H}^{B}(R),p_{1}^{B}(R),p_{2}^{B}(R),p_{L}^{B}(R))$ becomes
+Thus, accepting a buy recommendation the receiver’s posterior $\mathbf{p}^{B}(R):=(p_{H}^{B}(R),p_{1}^{B}(R),p_{2}^{B}(R),p_{L}^{B}(R))$ becomes
 $p_{H}^{B}(R):=\frac{q_{H}}{q_{H}+q_{1}\phi_{1}(R)+q_{2}\phi_{2}(R)},\quad\quad$ $p_{2}^{B}(R):=\frac{q_{2}\phi_{2}(R)}{q_{H}+q_{1}\phi_{1}(R)+q_{2}\phi_{2}(R)},$
@@
 If, in contrast, the sender could give four different ratings depending on the consumption experience, full revelation would be possible in our parsimonious model with only four product versions.
-However, in practice the greater complexity of most products would typically render such full revelation infeasible.
+In reality, recommendations can convey more content than whether a buyer liked a product or not, for instance, by including verbal descriptions.
+However, given the typical complexity of products, even relatively detailed recommendations will usually only transmit coarse information.
 Asymmetry of Stakes
@@
 4 Optimal Receiver Behavior
-In this section, we ask under which circumstances a receiver will accept recommendations, depending on his type $i$ and the recommendation system $\mathcal{R}$.
+In this section, we ask under which circumstances a receiver will accept all recommendations, depending on his type $i$ and the recommendation system $\mathcal{R}$, as introduced in Definition 1.
 In Section 4.1, we discuss through which channels recommendations affect beliefs, distinguishing between the objective and subjective content of the recommendations.
-Section 4.2 then characterizes the optimal behavior of the receiver by identifying thresholds which divide the type set into receivers that accept the recommendation and those that do not.
+Section 4.2 then characterizes the optimal behavior of the receiver by identifying thresholds which divide the type set into receivers that accept all recommendations and those that never do.
 4.1 The Effects of Recommendations on Beliefs
 Recommendations contain coarse information about the product under consideration, leading the receiver to update his beliefs about its quality.
-We first focus on buy recommendations, which have the following intuitive effects: [^11] In Appendix A.1, we provide a more formal analysis, decomposing the difference between posteriors and priors into three corresponding components.
+We first focus on buy recommendations, the total effect of which can be decomposed as follows: [^11] In Appendix A.1, we provide a more formal analysis, decomposing the difference between posteriors and priors into three corresponding components.
   1. (I)
@@
 By contrast, the third effect is fully type-dependent and thus subjective, resulting in an increase in the payoff for some receivers and a decrease for others.
-The following distinction is useful:
-Definition 2.
-A recommendation $r\in\{B,D\}$ that induces posteriors $\mathbf{p}^{r}$ given prior $\mathbf{q}$
-  1. (i)
-    shifts probability from $(0,1)$ to $(1,0)$ if $p_{2}^{r}-q_{2}<p_{1}^{r}-q_{1}$;
-  2. (ii)
-    shifts probability from $(1,0)$ to $(0,1)$ if $p_{2}^{r}-q_{2}>p_{1}^{r}-q_{1}$;
-  3. (iii)
-    does not affect the probability between $(0,1)$ and $(1,0)$ otherwise.
-As we will lay out below, whether a receiver considers the subjective content of a recommendation as positive or negative depends on his type $i$ and on the direction in which the recommendation shifts probability.
 The effects of don’t-buy recommendations are analogous to those of buy recommendations: First, the probability of a good product falls to zero; second, bad products become more likely relative to controversial products, and third, the odds between the controversial products change.
@@
 4.2 Characterizing Optimal Receiver Behavior
-We are now in a position to understand whether it is optimal for the receiver to accept a recommendation.
-To this end, denote by $U_{i}^{B}$ and $U_{i}^{D}$ player $i$’s expected payoff of buying a product with a buy and don’t-buy recommendation, respectively, and by $U_{i}^{0}$ the expected payoff of buying the alternative product with no recommendation.
+We are now in a position to understand whether it is optimal for the receiver to accept recommendations.
+To this end, note that, after receiving a recommendation, the receiver can either buy the recommended product or not buy it.
+Denote by $U^{B}(i)$ and $U^{D}(i)$ player $i$’s expected payoff of buying a product with a buy and don’t-buy recommendation, respectively, and by $U^{0}(i)$ the expected payoff of not buying, i.e., of buying the alternative product with no recommendation.
 We have
-$U_{i}^{r}$ $:=p_{H}^{r}+(1/2+i)p_{1}^{r}+(1/2-i)p_{2}^{r},\quad r\in\{B,D\};$ (3)
-$U_{i}^{0}$ $:=q_{H}+(1/2+i)q_{1}+(1/2-i)q_{2}.$ (4)
-For an expected-payoff maximizer, it is optimal to accept a buy recommendation if and only if $U_{i}^{B}-U_{i}^{0}\geq 0$, that is, the expected payoff from buying the recommended product is at least as high as the expected payoff from buying one of the alternative products (that have no recommendation).
-[^12] We use the tie-breaking rule that indifferent agents accept the recommendation.
-This condition can equivalently be written as
+$U^{r}(i)$ $:=p_{H}^{r}+(1/2+i)p_{1}^{r}+(1/2-i)p_{2}^{r},\quad r\in\{B,D\};$
+$U^{0}(i)$ $:=q_{H}+(1/2+i)q_{1}+(1/2-i)q_{2}.$
+Now, the difference between $U^{r}(i)$ and $U^{0}(i)$ determines the behavior of an expected-payoff maximizing receiver as follows.
+Definition 2.
+We say that a receiver $i$
+  1. (i)
+    accepts a buy recommendation if $U^{B}(i)\geq U^{0}(i)$; otherwise we say that he does not accept a buy recommendation;
+  2. (ii)
+    accepts a don’t-buy recommendation if $U^{0}(i)\geq U^{D}(i)$; otherwise we say that he does not accept a don’t-buy recommendation
+  3. (iii)
+    accepts all recommendations if he accepts both buy and don’t-buy recommendations;
+  4. (iv)
+    never accepts recommendations if he does not accept either buy and don’t-buy recommendations.
+Hence, a receiver who accepts all recommendations “does as he is told”, whereas a receiver who never accepts recommendations is a “contrarian” who does the opposite of what he is told.
+[^12] The use of the term aligns with Drehmann et al.
+(2005), who document contrarian behavior, i.e., investors going against the behavior of the predecessors, in an financial-markets experiment.
+Note that the above definition does not cover all conceivable types of behavior, as a receiver $i$ could accept buy recommendations and not accept don’t-buy recommendations, which is tantamount to always buying the recommended product and thus ignoring the content of the recommendations.
+Analogously, a receiver $i$ could ignore the content of the recommendations and never buy the recommended product.
+We will return to these cases shortly.
+The condition from Definition 2(i) for receiver $i$ to accept a buy recommendation can equivalently be written as
 $p_{H}^{B}-q_{H}+\frac{p_{1}^{B}-q_{1}}{2}+\frac{p_{2}^{B}-q_{2}}{2}\geq i[(p_{2}^{B}-q_{2})-(p_{1}^{B}-q_{1})].$ (1)
-To formulate this in compact form, we introduce the following terminology.
+For more compact formulations, we introduce the following terminology.
 Definition 3.
@@
     $\Delta_{S}^{r}:=(p_{2}^{r}-q_{2})-(p_{1}^{r}-q_{1}).$ (3)
-We can now restate condition (5) as
+We can now restate condition (1) about accepting a buy recommendation as
 $\Delta_{O}^{B}\geq\Delta_{S}^{B}i.$ (4)
-We use this formulation to justify our terminology.
-For the unbiased receiver $i=0$, equation (8) simplifies to $\Delta_{O}^{B}\geq 0$.
-As this receiver does not care about the shift in probabilities between the two controversial versions of the products, $\Delta_{O}^{B}\geq 0$ fully reflects the two objective effects (I) and (II) in the belief-updating decomposition in Section 4.1.
-For $i\neq 0$, the issue is more subtle.
-As long as $q_{1}=q_{2}$, the expected payoff conditional on the information that a product is controversial is $1/2$ for all $i$ and thus type-independent.
-Therefore, all types gain the same amount from (I) learning that a product is good or controversial instead of bad; the same applies to (II) learning that a product is good instead of controversial.
-For $q_{1}=q_{2}$, $\Delta_{O}^{B}$ therefore exactly captures these two type-independent (and thus objective) effects of a buy recommendation.
-[^13] The value of learning that the recommended product is more likely to be good is $p_{H}^{r}-q_{H}$, and the value of the change in the probability that it is controversial is $\frac{p_{1}^{r}+p_{2}^{r}-(q_{1}+q_{2})}{2}$.
-For $i\neq 0$ and $q_{1}\neq q_{2}$, the issue is complicated further.
-Though the effects captured by (I) and (II) in Section 4.1 are positive for all types, their size depends on the receiver’s type.
-Intuitively, if one of the controversial product versions (say (1,0))(1,0))( 1 , 0 ) ) is more likely than the other ex ante, learning that a product is controversial rather than bad is more valuable for types who prefer $(1,0)$ over $(0,1)$.
-Similarly, learning that the product is good rather than controversial is less valuable for those types.
-Even in this case, the interpretation of $\Delta_{O}^{B}\geq 0$ as an objective effect is justified because of the general agreement on its sign.
-We can now state our first result.
+[…]
 Proposition 1.
@@
   1. (i)
-    accepts a buy recommendation if and only if he accepts a don’t-buy recommendation;
+    either accepts all recommendations or never accepts recommendations;
   2. (ii)
-    accepts a recommendation if and only if $\Delta_{O}^{B}\geq i\Delta_{S}^{B}$.
-Intuitively, a receiver accepts a recommendation if he thinks that it has sufficient objective content and/or that his preferences are sufficiently aligned with the sender’s.
+    accepts all recommendations if and only if $\Delta_{O}^{B}\geq i\Delta_{S}^{B}$.
+The first part of the result states that Definition 2 indeed captures all relevant forms of behavior.
+That is, no type of receiver will ignore the informational content of recommendations by always or never buying the recommended product.
+Intuitively, a receiver accepts all recommendations if he thinks that they have sufficient objective content and/or that his preferences are sufficiently aligned with the sender’s.
+Conversely, a receiver never accepts recommendations if he thinks that they do not have sufficient objective content and/or that his preferences are not sufficiently aligned with the sender’s.
 Whether the recommendation is positive or negative does not matter for this assessment.
@@
 By contrast, the sign of the subjective effect $\Delta_{S}^{B}$ can be positive or negative.
-In short, a receiver accepts a buy recommendation unless the subjective effect is sufficiently large and his preferences are not sufficiently well aligned with those of the rest of the population.
-Denoting the type who is indifferent between accepting and not accepting recommendations as
-$\tilde{i}:=$ $\frac{\Delta_{O}^{B}}{\Delta_{S}^{B}}\text{ if }\Delta_{S}^{B}\neq 0\text{,}$ (9)
+The second part of the result shows which types accept all recommendations, which happens unless the subjective effect is sufficiently large and the receiver’s preferences are not sufficiently well aligned with those of the rest of the population.
+We emphasize that receivers who never accept recommendations still make use of them, by doing the opposite of what they are told.
+Proposition 1 tells us that we can decompose the population into at most two different groups: those who accept all and those who never accept recommendations.
+Denoting the type who is indifferent between always accepting and never accepting recommendations as
+$\tilde{i}:=$ $\frac{\Delta_{O}^{B}}{\Delta_{S}^{B}}\text{ if }\Delta_{S}^{B}\neq 0\text{,}$
 we can summarize optimal receiver behavior as follows.
@@
   1. (i)
-    If $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$, then all $i\in[-1/2,1/2]$ accept the recommendation.
+    If $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$, then all $i\in[-1/2,1/2]$ accept all recommendations.
   2. (ii)
-    If $\Delta_{S}^{B}<-2\Delta_{O}^{B}$, then all $i\geq\tilde{i}$ accept the recommendation.
+    If $\Delta_{S}^{B}<-2\Delta_{O}^{B}$, then all $i\geq\tilde{i}$ accept all recommendations.
   3. (iii)
-    If $\Delta_{S}^{B}>2\Delta_{O}^{B}$, then all $i\leq\tilde{i}$ accept the recommendation.
+    If $\Delta_{S}^{B}>2\Delta_{O}^{B}$, then all $i\leq\tilde{i}$ accept all recommendations.
 The intuition for Corollary 1, the proof of which is omitted, is straightforward.
+Without loss, consider a buy recommendation.
 For $\Delta_{S}^{B}=0$, the recommendation does not shift probability between the controversial product versions.
@@
 [^15] The effect on the expected payoffs in cases (ii) and (iii) is similar to the effect of a demand rotation as discussed in Johnson and Myatt (2006): For the types who accept the recommendation, the expected payoffs increase; for those who do not, they fall.
+Corollary 1 may potentially be useful to draw inferences about the relative importance of the objective and subjective effects of recommendations: If an observer sees a large share of receivers accepting recommendations, this can be interpreted as evidence for a relatively high objective effect of the recommendation.
 Finally, we analyze receiver behavior for extreme recommendation thresholds.
 Proposition 2.
-Suppose $R\rightarrow 1$ or $R\rightarrow 0$. Then, all types accept the recommendation irrespective of the product distribution.
+Suppose $R\rightarrow 1$ or $R\rightarrow 0$. Then, all types accept all recommendations irrespective of the product distribution.
 In the limits, both types of recommendations essentially contain only objective information.
@@
 Instead, it provides objective information that is equally valuable for all types, irrespective of their preferences for controversial products.
-Thus, all types accept the recommendation.
+Thus, all types accept all recommendations.
 5 The Value of the Recommendation System
@@
 Hence, the expected payoff of a receiver $i$ is
-$V_{0}(i):=U_{i}^{0}.$ (10)
+$V^{0}(i):=U^{0}(i).$ (5)
 Now suppose one of the products comes with a buy or don’t-buy recommendation, so the receiver has to update his belief about the quality distribution.
-A receiver who accepts the recommendation chooses this product (rather than one of the alternatives) only if it comes with a buy recommendation.
-With a recommendation system, the expected payoff of a receiver $i$ who accepts a recommendation is thus
-$V_{A}(i):=\pi^{B}U_{i}^{B}+(1-\pi^{B})U_{i}^{0}.$ (11)
+A receiver who accepts all recommendations chooses this product (rather than one of the alternatives) only if it comes with a buy recommendation.
+With a recommendation system, the expected payoff of a receiver $i$ who accepts all recommendations is thus
+$V^{A}(i):=\pi^{B}U^{B}(i)+(1-\pi^{B})U^{0}(i).$ (6)
 The first term is the probability of receiving a buy recommendation multiplied by the expected payoff from accepting it (and thus buying the recommended product).
 The second term is the probability of a don’t-buy recommendation multiplied by the expected payoff from buying the product without any recommendation, that is, based on priors.
-[^15] The notation $V_{A}(i)$ emphasizes the dependence of this expected payoff on the type of the receiver.
+[^16] The notation $V^{A}(i)$ emphasizes the dependence of this expected payoff on the type of the receiver.
 Clearly, $V_{A}$ also depends on $R$ and on the decision environment; similarly for the expressions below.
-A receiver who does not accept a recommendation will buy the recommended product if it comes with a don’t-buy recommendation.
+A receiver who never accepts recommendations will buy the recommended product if it comes with a don’t-buy recommendation.
 Otherwise, he will buy one of the other products.
 Hence, the expected payoff of such a receiver is
-$V_{N}(i):=$ $\pi^{B}U_{i}^{0}+(1-\pi^{B})U_{i}^{D}.$ (12)
-A buy recommendation does not create value for receivers who do not follow it, as such types choose the alternative for which no recommendation is available.
+$V^{N}(i):=$ $\pi^{B}U^{0}(i)+(1-\pi^{B})U^{D}(i).$ (7)
+A buy recommendation does not create value for receivers who do not accept it, as such types choose the alternative for which no recommendation is available.
 Instead, the don’t-buy recommendation, which is given with probability $(1-\pi^{B})$, does influence the behavior of the types who do not accept it: Instead of buying an alternative based on prior probabilities, updating leads them to choose the product for which they received the explicit recommendation not to buy.
@@
 (i) $\pi^{B}[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i]]$ if $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$; (ii) $(1-\pi^{B})F(\tilde{i})\left[\Delta_{O}^{D}-\Delta_{S}^{D}\mathbb{E}[i\mid i\leq\tilde{i}]\right]+\pi^{B}(1-F(\tilde{i}))\left[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i\mid i\geq\tilde{i}]\right]$ if $\Delta_{S}^{B}<-2\Delta_{O}^{B}$; (iii) $\pi^{B}F(\tilde{i})\left[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i\mid i\leq\tilde{i}]\right]+(1-\pi^{B})(1-F(\tilde{i}))\left[\Delta_{O}^{D}-\Delta_{S}^{D}\mathbb{E}[i\mid i\geq\tilde{i}]\right]$ if $\Delta_{S}^{B}>2\Delta_{O}^{B}$.
-Intuitively, when $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$, Corollary 1 implies that all receivers accept the recommendation as the objective effect dominates.
+Intuitively, when $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$, Corollary 1 implies that all receivers accept all recommendations as the objective effect dominates.
 Hence, by equation (6), the recommendation system creates value only in case of a buy recommendation, as a don’t-buy recommendation leads to the same expected payoff as without a recommendation system.
@@
 Finally, all of that only materializes with the probability of a buy recommendation, $\pi_{B}$.
-As the absolute value of $\Delta_{S}^{B}$ increases, the analysis becomes more complex because some types do not follow the recommendation.
+As the absolute value of $\Delta_{S}^{B}$ increases, the analysis becomes more complex because some types never accept recommendations.
 We consider the case $\Delta_{S}^{B}>2\Delta_{O}^{B}$ and reproduce the expression for $V(\mathcal{R})$ as follows: [^18] The argument for $\Delta_{S}^{B}<-2\Delta_{O}^{B}$ is analogous.
-$\underbrace{\pi^{B}F(\tilde{i})\left[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i\mid i\leq\tilde{i}]\right]}_{\text{types accept recommendation}}+\underbrace{(1-\pi^{B})(1-F(\tilde{i}))\left[\Delta_{O}^{D}-\Delta_{S}^{D}\mathbb{E}[i\mid i\geq\tilde{i}]\right]}_{\text{types reject recommendation}}.$ (13)
+$\underbrace{\pi^{B}F(\tilde{i})\left[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i\mid i\leq\tilde{i}]\right]}_{\text{types accept all recommendations}}+\underbrace{(1-\pi^{B})(1-F(\tilde{i}))\left[\Delta_{O}^{D}-\Delta_{S}^{D}\mathbb{E}[i\mid i\geq\tilde{i}]\right]}_{\text{types never accept recommendations}}.$ (8)
 In this case, the buy recommendation shifts probability from $(1,0)$ to $(0,1)$, which is unfavorable for receivers $i>0$, reducing their gains from accepting and possibly making them negative.
-Thus, for $\Delta_{S}^{B}>2\Delta_{O}^{B}$, only types below $\tilde{i}$ (a fraction $F(\tilde{i})$ of the receivers) accept the buy recommendation, which comes with probability $\pi^{B}$.
+Thus, for $\Delta_{S}^{B}>2\Delta_{O}^{B}$, only types below $\tilde{i}$ (a fraction $F(\tilde{i})$ of the receivers) accept all recommendations, which comes with probability $\pi^{B}$.
 In addition to the objective value of the recommendation ($\Delta_{O}^{B}$), these receivers obtain a subjective value.
@@
 It is positive if the negative contribution of types $i<0$ dominates the positive contribution of types $i>0$.
-The second component in (13), corresponding to the types who do not accept the recommendation, is arguably more surprising.
+The second component in (8), corresponding to the types who never accept recommendations, is arguably more surprising.
 Given their lack of alignment with the rest of the population, these types treat the recommendation not to buy as good news about the product and thus do not accept the recommendation.
@@
 Nonetheless, its informational content creates value for them.
-In the following, we shall often invoke the following symmetry assumption on the preference distribution.
+Going forward, we shall often invoke the following symmetry assumption on the preference distribution.
 Assumption 1.
 $F(-i)=1-F(i)$ for all $i\in[-1/2,1/2]$.
-With this assumption in place, the quantity
-$\beta:=F(1/2-R)$ (14)
-has a simple interpretation.
-Given a threshold $R$, the parameter $\beta$ captures the fraction of the population that is willing to recommend product versions $(1,0)$ or, equivalently (given the symmetry of the distribution), $(0,1)$.
+We now define the quantity
+$\beta:=F(1/2-R).$
+Due to Assumption 1, it has a simple interpretation: Given a threshold $R$, the parameter $\beta$ captures the fraction of the population that is willing to recommend product versions $(1,0)$ or, equivalently (given the symmetry of the distribution), $(0,1)$.
 For a fixed type distribution $F$, the variables $\beta$ and $R$ are inversely related.
@@
 Proposition 4.
-Suppose Assumption 1 holds. Then, all types accept the recommendation and
+Suppose Assumption 1 holds. Then, all types accept all recommendations and
 $V(\beta)=\pi^{B}\Delta_{O}^{B}\text{.}$
-To understand why all types accept a recommendation, the belief-updating decomposition in Section 4.1 is helpful.
-Its third component captures the potential change in the odds of the controversial products, which induces a purely type-dependent payoff effect.
+To understand why all types accept all recommendations, the belief-updating decomposition in Section 4.1 is helpful.
+Its third component captures the potential shift of probabilities between controversial products, which induces a purely type-dependent payoff effect.
 Given the symmetry assumption, both controversial product versions lead to a buy recommendation with probability $\beta$.
-Hence, a recommendation does not change the odds of the controversial products.
+Hence, a recommendation does not shift probabilities between the controversial products.
 As the other two effects of a buy recommendation on payoffs are positive for all types, the buy recommendation increases expected payoffs for any type, yielding the result.
@@
 Otherwise, one effect dominates the other.
-Based purely on uncontroversial products, a buy recommendation is more likely than a don’t-buy recommendation when the good product is relatively more likely than the bad one ($\sigma>1$).
+Based purely on uncontroversial products, a buy recommendation would be more likely than a don’t-buy recommendation when the good product is relatively more likely than the bad one ($\sigma>1$).
 It is thus worthwhile to increase the expected average payoff conditional on a buy recommendation by increasing $R$ at the expense of the probability of receiving such a recommendation.
@@
 We can now analyze the extent to which the result in Proposition 5 that, generically, extreme thresholds are desirable is an artifact of the symmetry assumption.
-This is challenging because, contrary to the symmetric case, the set of types that accept a recommendation depends on the threshold $R$.
+This is challenging because, contrary to the symmetric case, the set of types that accept all recommendations depends on the threshold $R$.
 Reflecting this complication, there is no direct counterpart to Proposition 5.
@@
 It shows that an interior optimum exists when (i) controversial products are likely ($Q$ is large) and the population is very asymmetric ($a$ is far away from $1$) or otherwise when (ii) the uncontroversial product versions are similarly likely ($\sigma$ is close to $1$).
+[^25] The figure might suggest that the parameter region where $R^{\ast}\rightarrow 1$ is in some sense larger than the area where $R^{\ast}\rightarrow 0$.
+This is an artefact of the otherwise quite convenient parameterization of the preference distribution but does not have any significant economic meaning.
 The result starkly contrasts the symmetric case, where an interior solution is only possible in the knife-edge case $\sigma=1$.
@@
 Thus, the choice of the optimal threshold no longer only depends on $\sigma$ but also on the asymmetry parameter $a$ and the prevalence of controversial products $Q$.
-For a wide range of parameters, the conditions given in Proposition 6 are necessary as well as sufficient:
+We explore under which circumstances the conditions given in Proposition 6 are necessary as well as sufficient:
 Proposition 7.
@@
 The proof essentially proceeds in two steps.
-First, all types accept the recommendation when $Q$ is sufficiently small, since then the subjective content in the recommendation vanishes.242424 $Q\rightarrow 0$ implies that $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$, ensuring that all types follow by Corollary 1.
+First, all types accept all recommendations when $Q$ is sufficiently small, since then the subjective content in the recommendation vanishes.262626 $Q\rightarrow 0$ implies that $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$, ensuring that all types accept all recommendations by Corollary 1.
 The key step is that $\Delta_{S}^{B}\rightarrow 0$ as $Q\rightarrow 0$, while $\Delta_{O}^{B}\rightarrow 1-q_{H}>0$ when $Q\rightarrow 0$.
@@
 [^27] The figure is drawn for $Q=0.1$, in which case the if-and-only-if statement is correct for $\sigma\in[1/60,60]$ and $a\in[1/10000,10000]$.
-In a nutshell, the result shows that an interior solution exists if (i) controversial products are sufficiently prevalent, (ii) the population asymmetry is substantial, and (iii) the odds ratio of the uncontroversial products is sufficiently small or large.
+In a nutshell, the result shows that an interior solution exists if (i) controversial products are sufficiently prevalent, (ii) the population asymmetry is substantial, and (iii) the odds ratio of the uncontroversial products is sufficiently close to 1.
 To understand this intuitively and to contrast it with the symmetric case, suppose that $\sigma>1$.
@@
 5.3.1 Preference Heterogeneity
+In this section, we analyze how changes in the (exogenously given) preference distribution affect the value of the recommendation system.
+[^28] When we consider distinct sender and receiver distributions in Section 6.1, we will build further on this and take a design perspective instead of the comparative statics approach here.
 Under Assumption 1, we can analyze increases in the polarization of the distribution of preferences by keeping the threshold $R$ fixed and varying $\beta$ [^29] For $R=1/2$, we must have $\beta=1/2$ because the symmetry of $F$ requires $F(0)=1/2$, so this exercise is only meaningful for thresholds $R\neq 1/2$.
@@
 5.3.2 Prevalence of Controversial Products
-Without controversial products, preference heterogeneity would not matter: All receivers would accept recommendations, as they would reveal whether the product is good or bad.
+Without controversial products, preference heterogeneity would not matter: All receivers would accept all recommendations, as they would reveal whether the product is good or bad.
 Hence, the value of the recommendation system would be independent of the preference distribution.
@@
 $\beta$ and $\sigma$ determine which of these effects dominates.
-When $\beta$ is high, the increase induced by controversial products leads to more buy recommendations, strengthening the positive effect.
+When $\beta$ is high, the greater likelihood that controversial products receive buy recommendations increases the total probability of buy recommendations, strengthening the positive effect.
 When $\sigma$ is high, the lower share of good products dominates, strengthening the negative effect.
@@
 We show that the results of the benchmark case easily generalize to this setting.
+More importantly, the general setting allows us to treat the choice of the sender distribution as a potentially interesting design variable.
 The sender distribution $F$ fully determines the probabilities of buy and don’t-buy recommendations, the posterior probabilities, and thus the objective and subjective effects and the expected payoffs, $V_{A}$ and $V_{N}$.
@@
 Suppose the sender distribution $F$ satisfies Assumption 1.
-Then, all types accept the recommendation.
+Then, all types accept all recommendations.
 Further, $V(R)=\pi^{B}_{F}[\Delta_{O,F}^{B}-\Delta_{S,F}^{B}\mathbb{E}_{G}[i]]$ is
@@
     constant in $R$ if $\sigma=\frac{Q-(q_{1}-q_{2})\mathbb{E}_{G}[i]}{Q-(q_{2}-q_{1})\mathbb{E}_{G}[i]}$.
-Recall that in the benchmark model symmetry of $F$ implies that all receivers follow the recommendation and the subjective effect of the recommendation cancels out.
+Recall that in the benchmark model symmetry of $F$ implies that all receivers accept all recommendations and the subjective effect of the recommendation cancels out.
 This is still true if $F$ is distinct from $G$ as long as $G$ is also symmetric and hence $\mathbb{E}_{G}[i]=0$.
-For more general distributions $G$, all receivers still accept the recommendation for the same reason as above.
+For more general distributions $G$, all receivers still accept all recommendations for the same reason as above.
 However, as the receiver distribution is not symmetric, the subjective effect does not cancel out anymore.
@@
 With asymmetrically distributed populations, Propositions 6 and 7 hold for arbitrary receiver distributions as long as we impose Assumption 2 on the sender distribution.
+Allowing for differences in sender and receiver distributions makes it possible to address another important design issue.
+In some cases, it is natural to assume that a designer, for instance a platform, can influence the sender distribution or at least the receiver’s perception of it.
+Assuming that the platform has better knowledge about sender types, it may choose to share some of this information with the receivers or even restrict the option of sending recommendations to certain types of senders (and inform the receivers accordingly).
+For instance, many platforms such as Amazon inform browsing customers that “customers who bought this item also bought” another item.
+There are different ways to model this.
+First, suppose that $F$ is symmetric and that the designer can render the sender distribution more or less precise using mean-preserving contractions or mean-preserving spreads, respectively.
+Intuitively, one would expect more precise knowledge about the sender to increase the value of the recommendation system.
+However, the logic of Proposition 9 shows that this is not necessarily the case.
+As in Section 5.3.1, the effects of mean-preserving contractions and spreads can be captured as changes in $\beta=F(1/2-R)$, just like changes in $R$.
+With mild adaptions, the proof of Proposition 9 can thus be used to show that whether more or less precise information about the sender increases the value of recommendations depends on the parameters.
+Second, suppose that the sender and receiver distributions are fixed but that the designer can provide the receivers with a signal about the sender that allows for inference about the sender’s type.
+To make the point as starkly as possible, suppose the signal fully reveals the sender’s type.
+If the sender is revealed to be type $i=0$, all types accept all recommendations, as both types of recommendations contain only objective information.
+In contrast to Proposition 2—where this behavior is shown to ensue with extreme thresholds—this would arise for all thresholds $R\in(0,1)$.
+Intuitively, as the type $i=0$ is indifferent between controversial products, her recommendation never reveals subjective information.
+In contrast, if the signal reveals a sender of type $i\neq 0$, the receiver will be able to learn about controversial products, too.
+[^33] For instance, if $i=1$ and $R\in(0,1)$, a buy recommendation reveals that the product is either of version $(1,1)$ or $(1,0)$ Thus, having a sender who is partial to some controversial product is not necessarily a bad thing.
+Finally, suppose that the sender and receiver distributions are not independent but somehow correlated.
+Recommendations would then be “tailored” to the receiver $\bar{i}$, as they would be drawn from a sender distribution $F(i|\bar{i})$, resulting in type-specific posteriors $p^{r}_{j}(i)$ for $j\in\{H,1,2,L\}$.
+In such a setting, an analogous result to Proposition 1 still holds, establishing that either all all or no recommendations are accepted.
+Further analysis would give rise to similar complications as in Section 5.2.2 with asymmetric distributions, as the the correlation between sender and receiver distribution would affect the set of types who accept all recommendations.
 6.2 Multiple Recommendation Thresholds
 We now analyze the effects of reducing the coarseness of the recommendation system.
-We maintain all assumptions of the benchmark case, except that there are two threshold levels $R_{1},R_{2}\in(0,1)$ with $R_{2}\geq R_{1}$ and three types of recommendations $r\in\{D,N,B\}$ which we interpret as “don’t-buy,” “neither buy nor don’t-buy,” and “buy.” The sender gives a buy recommendation if the payoff is at least $R_{2}$, a don’t-buy recommendation if it is below $R_{1}$, and neither if it is in $[R_{1},R_{2})$.
+We maintain all assumptions of the benchmark case, except that there are two threshold levels $R_{1},R_{2}\in(0,1)$ with $R_{2}\geq R_{1}$ and three types of recommendations $r\in\{D,N,B\}$ which we interpret as “don’t-buy,” “neither buy nor don’t-buy,” and “buy.” [^34] Such a two-threshold structure is found to be optimal in models of strategic communication in the context of social learning by Smirnov and Starkov (2024) and Bénabou and Vellodi (2024).
+The sender gives a buy recommendation if the payoff is at least $R_{2}$, a don’t-buy recommendation if it is below $R_{1}$, and neither if it is in $[R_{1},R_{2})$.
 We maintain Assumption 1.
@@
 The intermediate recommendation $r=N$, however, rules out both of these versions, so the receiver knows that the recommended product is controversial.
-We say that a receiver accepts a recommendation if he buys if $r=B$ and never buys if $r=D$, allowing for either behavior if $r=N$.
+We say that a receiver accepts all recommendations if he buys if $r=B$ and does not buy if $r=D$, allowing for either behavior if $r=N$.
 Assumption 1 again implies that the receiver always buys a product with $r=B$ and never buys if $r=D$.
@@
 First, as in the benchmark, all types accept $r=B$, resulting in the component $\pi^{B}\Delta_{O}^{B}$.
-[^30] As before, following don’t-buy recommendations does not create value.
+[^35] As before, accepting don’t-buy recommendations does not create value.
 Moreover, $r=N$ can create value for the types that buy following that neutral recommendation.
@@
 7 Conclusion
-We have set out to study the role of preference heterogeneity in recommendations, asking questions such as: Under what circumstances do people follow a recommendation?
+We have set out to study the role of preference heterogeneity in recommendations, asking questions such as: Under what circumstances do people accept a recommendation?
 What determines the value of a recommendation system and when is it maximized?
@@
 $p_{H}^{B}+p_{1}^{B}(1/2+i)+p_{2}^{B}(1/2-i)\geq q_{H}+q_{1}(1/2+i)+q_{2}(1/2-i)\text{.}$ (11)
-We shall show that conditions (16) and (17) are equivalent by proving the following two steps.
+We shall show that conditions (10) and (11) are equivalent, thus proving the claim that receivers either accept all recommendations or never accept recommendations.
+To do so, we prove the following two steps.
 Step 1: (10) is equivalent with
@@
 (ii) If $\Delta_{S}^{B}<-2\Delta_{O}^{B}$, inserting (5), (6) and (7) yields
-[…]
-$=\int_{-1/2}^{\tilde{i}}\pi^{B}\left(q_{H}+q_{1}(\frac{1}{2}+i)+q_{2}(\frac{1}{2}-i)\right)+(1-\pi^{B})\left(p_{H}^{D}+p_{1}^{D}(\frac{1}{2}+i)+p_{2}^{D}(\frac{1}{2}-i)\right)dF(i)$
-$+\int_{\tilde{i}}^{1/2}\pi^{B}\left(p_{H}^{B}+p_{1}^{B}(\frac{1}{2}+i)+p_{2}^{B}(\frac{1}{2}-i)\right)+(1-\pi^{B})\left(q_{H}+q_{1}(\frac{1}{2}+i)+q_{2}(\frac{1}{2}-i)\right)dF(i)$
-$-\int_{-1/2}^{1/2}q_{H}+q_{1}(\frac{1}{2}+i)+q_{2}(\frac{1}{2}-i)dF(i)$
 […]
@@
 Therefore, receiver $i$ buys the product with mixed recommendations if and only if
-$\frac{1}{2}+i\frac{q_{1}-q_{2}}{q_{1}+q_{2}}\geq V_{0}(i)=q_{H}+q_{1}(1/2+i)+q_{2}(1/2-i).$ (26)
+$\frac{1}{2}+i\frac{q_{1}-q_{2}}{q_{1}+q_{2}}\geq V^{0}(i)=q_{H}+q_{1}(1/2+i)+q_{2}(1/2-i).$ (20)
 The indifferent consumer satisfies

### v3

@@
-A Theory of Recommendations
+A Theory of “Likes”
 […]
 [linecolor=black,backgroundcolor=gray!10,linewidth=1pt]mybox
-A Theory of Recommendations
+A Theory of “Likes”
 Jean-Michel Benkert and Armin Schmutzler Benkert: Department of Economics, University of Bern, jean-michel.benkert@unibe.ch.
 Schmutzler: Department of Economics, University of Zurich, and CEPR, armin.schmutzler@econ.uzh.ch.
-We are grateful for very helpful discussions with Piotr Dworczak, Johannes Johnen, Igor Letina, Marc Möller, Nick Netzer, Volker Nocke, Egor Starkov, Jakub Steiner and seminar participants in Bern, Fribourg, Zurich, at SAET (2023), Swiss Society of Economics of Statistics (2024), Swiss IO Day (2024) and Verein für Socialpolitik (2024).
-(October 11, 2024)
+This paper was previously circulated under the title “A Theory of Recommendations”.
+We are grateful for very helpful discussions with Piotr Dworczak, Johannes Johnen, Igor Letina, Matt Mitchell, Marc Möller, Nick Netzer, Volker Nocke, Marcus Reisinger, Egor Starkov, Jakub Steiner and seminar participants in Bern, Fribourg, Zurich, at MaCCI (2025), SAET (2023), Swiss Society of Economics of Statistics (2024), Swiss IO Day (2024) and Verein für Socialpolitik (2024).
+(May 1, 2025)
 This paper investigates the value of recommendations for disseminating economic information, with a focus on frictions resulting from preference heterogeneity.
 We consider Bayesian expected-payoff maximizers who receive non-strategic recommendations by other consumers.
-The paper provides conditions under which different consumer types accept these recommendations.
-Moreover, we assess the overall value of a recommendation system and the determinants of that value.
-Our analysis highlights the importance of disentangling objective information from subjective preferences when designing value-maximizing recommendation systems.
+We show by which channels a recommendation system influences consumer demand and welfare.
+Our analysis reveals how the welfare-maximizing design of a recommendation system may differ from what a profit-maximizing designer would choose to do.
 Keywords: recommendations, preference heterogeneity, optimal design
@@
 1 Introduction
-Recommendations play an essential role for the diffusion of economic information.
-Job applicants frequently rely on assessments of potential employers by their friends.
-Conversely, job candidates often have to provide letters of recommendation from former employers.
-Consumers may base their choices of travel destinations, restaurants, or doctors on what they have heard from their acquaintances.
-In the digital economy, product ratings, typically reflecting anonymous consumer reports, are omnipresent.
-Despite their relevance, recommendations are unlikely to function smoothly.
-For instance, too few recommendations may be given (Kremer et al., 2014; Che and Hörner, 2018), recommenders may be positively selected (Acemoglu et al., 2022), and interested parties might interfere with the process by choosing which recommendations to publish (Bolton et al., 2013; Tadelis, 2016).
-In the literature, it is common to make the simplifying assumption that product qualities can be objectively ranked.
-[^1] See also Che and Hörner (2018) and Hui et al.
-(2023).
-Instead, our paper assumes that preferences are heterogeneous, in which case product qualities cannot be objectively ranked and recommendations may be hard to interpret, even without strategic biases.
-We ask whether and how recommendations can be useful in spite of preference heterogeneity.
-Even well-intended recommendations by non-strategic players may reflect the senders’ preferences rather than the objective truth.
-A receiver who does not question the informational content of a recommendation may therefore make biased decisions.
-To motivate our approach, consider a consumer’s choice between medical doctors.
-This is a complex problem, as the perceived quality of medical care will typically depend on many aspects.
-How carefully does a doctor listen to the patient?
-Is the technical equipment up to date?
-Is the staff friendly?
-How crowded is the waiting room?
-Moreover, different people may weigh different aspects differently.
-Despite such preference heterogeneity, rather than randomly choosing a doctor, people often base their decision on recommendations.
-Importantly, senders have limited stakes in the process and therefore lack incentives to think through the effects of their recommendation carefully.
-By contrast, adequately using the information can be crucial for the receiver.
-Finally, even if the sender is willing to think through the recommendation carefully, it will typically not be possible to convey all relevant information.
-For all these reasons, it is not obvious why or when recommendation can be useful.
-This example illustrates three key features of our theory of recommendations:
-Heterogeneity of Products and Preferences: Products come in different versions and consumers value them differently.
-Asymmetry of Stakes: Making sure that the recommendation is not misunderstood is more important for the receiver than for the sender.
-Coarseness of Recommendations: The sender is unable to provide enough detail for all relevant aspects of the service to become fully transparent to the receiver.
-These features will be even more salient for ratings provided on internet platforms, as such ratings usually do not come from close acquaintances.
-Hence, the receivers should consider that senders have preferences which differ from their own and are likely to give an incomplete account of their consumption experience without exerting much effort.
-Backing out the implications of the recommendation for the receiver’s own decision is therefore a complex task.
-Relevant considerations for the interpretation are these: To what extent does the recommendation contain objective truth?
-And when it does not, how likely is it that the preferences of the sender and receiver are aligned?
-The answers to these questions will determine whether it is in the receiver’s interest to accept the recommendation and how valuable it is.
-Our model captures the above three features of recommendations in the simplest possible way.
-An agent must decide between several options, referred to as products for ease of exposition.
-Products are two-dimensional objects, differing only in whether they have high or low quality in each dimension.
-Thus, four product versions exist.
+Recommendations play an essential role for the diffusion of economic information in diverse domains.
+For instance, consumers may base their choices of travel destinations, restaurants, or doctors on what they have heard from their acquaintances.
+In the digital economy, the supply of recommendations is increasingly mediated through simple mechanisms such as “likes,” star ratings, or thumbs up/down.
+These coarse signals compress personal evaluations into binary or low-dimensional feedback that guides others’ decisions.
+In this paper, we ask how recommendations can function when preferences are heterogeneous and product qualities cannot be objectively ranked: is a “like” a reflection of inherent quality or merely an expression of individual taste?
+Consider a user of a streaming platform deciding whether to “like” a documentary.
+Perceived quality depends on many aspects—informativeness, emotional impact, cinematography, pacing—and different viewers weigh these aspects differently.
+Moreover, consumers who send recommendations often have little incentive to consider how their like will be interpreted, while the receiver may depend on it to avoid poor matches.
+And even a thoughtful sender cannot convey all relevant nuances through a coarse signal.
+Yet, recommendation systems are pervasive and people routinely rely on them.
+This raises fundamental questions: When are such systems useful?
+What shapes the value of the recommendations they generate for consumers?
+And how might the design incentives of platforms diverge from the interests of consumers?
+The streaming example highlights three core features of recommendations that motivate our modeling approach.
+First, products and preferences are heterogeneous: Products come in different versions, and consumers value them differently.
+Second, stakes are asymmetric: Avoiding misunderstanding is more important for the receiver than for the sender.
+Third, recommendations are inherently coarse: The sender cannot fully convey all relevant aspects of the product.
+[^1] These features arise in virtually all recommendation contexts, including personal interactions with a friend.
+However, they are particularly pronounced on online platforms, where recommendations are typically stripped of context, reduced to simple signals, and come from (largely) anonymous senders.
+Our model captures these features in a parsimonious way.
+An agent must decide between two products, which are two-dimensional objects differing in whether they have high or low quality in each dimension.
+This gives rise to four distinct product versions.
 There is a continuum of agent types.
-All types agree that a product is worth choosing if it has high quality in both dimensions and is thus (objectively) good.
-Conversely, they agree that a product is not worth choosing if it has low quality in both dimensions and is thus (objectively) bad.
-Accordingly, we call the products uncontroversial in both of these cases.
-By contrast, we refer to products as controversial when they have high quality in one dimension and low quality in the other.
-An agent prefers a controversial product that has high quality in the dimension that is more relevant for them, and is thus subjectively good, to a product that has high quality only in the less-relevant dimension (thus subjectively bad).
-Moreover, even agents who agree about which controversial product they prefer may differ in the intensity of these preferences (as captured by the continuous type distribution).
-We assume that products are experience goods so that each agent only has stochastic knowledge of the quality distribution before consumption: They know the probability of each product version.
-The agent obtains a recommendation from a previous consumer of one of the products.
+All types agree that a product with high quality in both dimensions is worth choosing (objectively good), and that one with low quality in both is not (objectively bad)—we refer to such products as uncontroversial.
+By contrast, controversial products have high quality in one dimension and low quality in the other.
+An agent prefers the controversial product that scores highly in the dimension he cares more about (subjectively good) to the one that scores highly in the less important dimension (subjectively bad).
+We assume that products are experience goods so that, before consumption, each agent only knows the probability of each product version.
+However, he obtains a recommendation from a previous consumer of one of the products.
 The recommendation reveals whether the sender had a consumption experience that was sufficiently good (a buy recommendation), with payoffs above an exogenous threshold level $R$, or not (a don’t-buy recommendation).
-Independent of $R$, the sender (she) always gives a buy recommendation for a good product and a don’t-buy recommendation for a bad product.
+Therefore, the sender (she) always gives a buy recommendation for a good product and a don’t-buy recommendation for a bad product, independently of the threshold $R$.
 In addition, she sends a buy recommendation for a controversial product if her preferences for the product are strong enough.
-In line with her low stakes in the issue, the sender thus only provides a coarse description of her own consumption experience.
+This approach combines the features of recommendations discussed above in a stylized way: In line with her low stakes in the issue, the sender only provides a non-strategic and coarse description of her own consumption experience.
+[^2] In reality, recommendations can convey more content than whether a buyer liked a product or not, for instance, by including verbal descriptions.
+However, given the typical complexity of products, even relatively detailed recommendations will usually only transmit coarse information.
 By contrast, the receiver (he) of the recommendation, with his high stakes in the issue, will carefully evaluate the informational content of the recommendation using Bayesian inference, taking into account the distribution of product qualities, his own preferences, and those of the population.
-In this framework, we address a variety of questions: How do recommendations create value for receivers, how can we interpret this value, and when is it maximal?
-If one could influence the payoff threshold beyond which a sender gives a buy recommendation, how high should it be?
-Are recommendations valuable even when uncontroversial products are unlikely?
-How does the value depend on how polarized the population is?
-To answer these questions, we need to derive intermediate results of independent interest.
-Most importantly, we investigate under which conditions receivers accept all recommendations.
-We find that for a Bayesian expected-payoff maximizer, two key quantities measuring the objective and subjective content of the recommendation are decisive for receiver behavior.
-These quantities are simple functions of the model primitives.
-Intuitively, the objective informational content of a recommendation is high if it makes the product sufficiently more likely to be (objectively) good and less likely to be bad.
-The subjective content captures the effect of the recommendation on the relative probability of the two controversial product versions.
-The evaluation of the subjective content thus depends on the receiver’s type.
-When the recommendation bears little or no subjective content, all types accept all recommendations, thus buying after a buy recommendation and choosing the outside option (one of the other alternatives) after a don’t-buy recommendation.
-As the subjective content increases, revealing that one of the two controversial product versions becomes more likely, only receivers without strong preferences for the other version accept all recommendations.
-When the recommendation threshold $R$ approaches the maximal payoff, a buy recommendation essentially becomes objective, and similarly for a don’t-buy recommendation when $R$ approaches the minimal payoff.
-As receivers are aware of this, they accept all recommendations in each of these polar cases since the subjective component vanishes.
-Based on this analysis, we then study the value of recommendation systems.
-We think of this value as the expected-payoff increase of a randomly chosen receiver resulting from a recommendation given by a randomly chosen sender.
-When the recommendation is objective, or the subjective content is small, the analysis is comparatively simple: As all receivers accept all recommendations, they choose the outside option after a don’t-buy recommendation, just as if they had not received any recommendation at all.
-Therefore, only buy recommendations create value—they do so by inducing the receiver to purchase the product if it generates a higher expected payoff for him than the outside option.
-When the recommendation contains more subjective content, only a subset of the receivers accept it, rendering the analysis more complex.
-The value of the recommendation system then has two sources.
-First, receivers who accept all recommendations benefit because they expect the recommended product to be better than the outside option.
-Second, receivers who never accept recommendations benefit from buying the product in spite of a don’t-buy recommendation: These receivers have preferences that are not well aligned with the general population so that the don’t-buy recommendation provides good news about the chances of obtaining a subjectively good product rather than one they like less.
-Next, we turn to the design of recommendation systems.
-Motivated by the practice of digital platforms, we consider the recommendation threshold $R$ as a design variable.
-When the type distribution is symmetric, so that there is no preference bias on average, the value of the recommendation system is a monotone function of the threshold.
-If the value is increasing, the product optimally receives a buy recommendation only if it is good; if the value is decreasing, the product optimally receives a don’t-buy recommendation only if it is bad.
-The former case arises if good products are more likely ex ante than bad products are, and conversely for the latter case.
-Either way, the content of the recommendation essentially becomes objective.
-In contrast, with an asymmetrically distributed population, value maximization does not necessarily require an extreme threshold.
-Rather than revealing only whether a product is good or bad, it may be optimal for the sender to give a buy recommendation even if the product is controversial but sufficiently aligned with her preferences.
-Intuitively, giving up the exclusive focus on objective content makes sense when both controversial product versions are common and there is broad agreement on which of them is preferable.
-By contrast, extreme recommendation thresholds that only reveal objective content are still desirable if the product is much more likely to be objectively good than objectively bad or vice versa.
-We then discuss further determinants of the value of the recommendation system, assuming that the population is symmetric.
-First, we find that an increase in the probability of controversial products lowers the value only if the probabilities of the two controversial versions are similar.
-Second, whether a mean-preserving spread in the population distribution (corresponding to a more polarized population) increases the value of a given recommendation system depends on how strict the recommendation threshold is and on how common good products are relative to bad ones.
-Finally, we show the robustness of our results.
-With mild adjustments, the main insights continue to hold when sender and receiver distributions are distinct—a situation that could arise because, for instance, recommendations stem from early adopters who do not represent the population at large.
-This also has non-trivial design implications: Our results on the relation between mean-preserving spreads can be adapted to show how a designer would want to choose the sender distribution, which may be at least partly under his control.
-Second, we allow for more than two types of recommendations.
-Again, extreme recommendation thresholds increase the probability that a recommendation has objective content and thus increase the value of the system when the population is symmetrically distributed.
-Third, we show that having access to many recommendations does not necessarily change receiver behavior compared to the case of a single recommendation with an optimal threshold.
-Then, having access to multiple recommendations does not increase the value.
-Section 2 relates the paper to the literature.
-Section 3 introduces the model and discusses its key assumptions.
-In Section 4, we describe the conditions under which different receiver types accept all recommendations.
-Then, in Section 5 we first characterize the value of a recommendation system before investigating the determinants of the value and the design of value-maximizing systems.
+Further, our decision to focus only on buy and don’t-buy recommendations is a particularly simple way of capturing the ordinal nature of most (if not all) recommendation systems, which typically ask for giving “likes” or “dislikes”, rankings from 1 to 5 stars, and so forth.
+[^3] We do consider an extension to more than two types of recommendations in Section 6.
+Recommendations contain objective and subjective content.
+Intuitively, from a buy recommendation a receiver learns that the product is not objectively bad and the chances that it is objectively good have increased.
+All receivers agree that this information makes it more attractive to buy the product.
+Moreover, the buy recommendation potentially influences the relative probabilities of controversial products, making it more likely that the product is the one that most consumers prefer.
+For consumers whose tastes are aligned with the majority of the population, this subjective content reinforces the positive objective content conveyed by the message.
+Thus, such receivers will definitely accept the recommendation.
+Contrarians, however, who understand that their tastes are not aligned with those of the rest of the population may decide not to accept the recommendation if the negative subjective content dominates the positive objective content.
+Building on these simple insights into receiver behavior, we characterize both the value created by the recommendation system and the resulting demand for the recommended product.
+A recommendation system influences two groups of buyers: those who follow a buy recommendation (i.e., types who accept recommendations) and those who buy despite receiving a don’t-buy recommendation (i.e., types who reject recommendations).
+Each of these two groups may benefit from the recommendation system.
+For types whose preferences are sufficiently aligned with the general population, accepting buy recommendations provides value by successfully guiding them toward a product that outperforms the outside option in expectation.
+By contrast, for types whose preferences are misaligned with the general population, don’t-buy recommendations can provide a valuable positive signal about the fit of the product with their subjective preference.
+We then turn to the question of optimal design.
+In practice, platforms may exert influence over how users give recommendations.
+For example, Uber employs a five-star system but informs riders that “5 stars means the ride went without issues,” while “4 stars or less” prompt requests for feedback or suggested improvements (Uber, 2024).
+[^4] In the case of Airbnb, a service provider notes that “A 5-star rating on Airbnb means that everything was as expected and that guests are happy with their stay,” while “4 stars is also considered a positive rating, but it’s not enough,” and that “3 stars or less are considered the negative feedback” (Hospitable, 2024).
+This suggest that platforms do not merely choose a recommendation format but also influence how users map experiences into ratings.
+In our framework, this mapping is governed by the threshold $R$: by shaping what counts as a positive signal, the platform effectively controls how selectively buy recommendations are given.
+It is natural to assume that the platform’s choice of $R$ is guided by profit-maximization motives.
+As a result, the platform may aim to maximize demand, either because it receives fixed transaction fees or because higher traffic yields greater advertising revenue.
+We investigate the resulting choice of $R$ and compare it to what would be optimal from a consumer perspective.
+We begin with a setting in which preference heterogeneity is purely idiosyncratic noise, so that the population is symmetric with respect to their valuations for the controversial products.
+In this case, there is no systematic preference bias in the population, so recommendations have no subjective content.
+We find that the value for consumers generated by the recommendation system is monotone in the threshold $R$.
+If good products are more likely than bad ones, the value increases with $R$, and the optimal threshold induces consumers to issue buy recommendations only when the product is objectively good.
+If bad products are more likely than good ones, the value decreases with $R$, and, from a consumer perspective, buy recommendations should be given unless the product is objectively bad.
+Either way, the designer optimally renders recommendations essentially objective in content.
+In contrast with the value of the recommendation system for consumers, the demand for the recommended product is always decreasing in the threshold.
+Thus, a platform that seeks to maximize sales favors a low threshold, so that buy recommendations are issued except for objectively bad products.
+As a result, the objectives of maximizing value and maximizing demand are aligned in some environments but stand in sharp conflict in others.
+We conclude the analysis of a symmetric population by examining how both the value of the recommendation system and the demand it induces depend on the economic environment.
+We find nuanced effects of changes in the prevalence of controversial products and in the degree of polarization in the population.
+While the symmetric benchmark is analytically clean and provides useful intuition, it leaves no room for subjective content of recommendations.
+We therefore extend the analysis to asymmetric populations, where the controversial product versions are not valued equally on average.
+In this setting, value maximization no longer necessarily requires an extreme threshold which conveys the message whether a product is objectively good or objectively bad.
+Instead, it may be desirable to use interior thresholds which can effectively communicate whether a product is in line with majority tastes.
+Next, we turn to demand maximization in the asymmetric setting.
+Using a combination of theoretical and numerical analysis, we find that a demand-maximizing platform usually prefers a low threshold.
+While this behavior may coincide with value maximization in some environments, the two objectives generally diverge in the presence of asymmetries.
+Finally, various extensions of the model lead to further insights.
+First, we allow the sender and receiver populations to differ, as in cases where recommendations come from, say, early adopters.
+Our main results remain robust, and we discuss how the designer may want to shape the sender distribution.
+Second, we consider richer recommendation systems with more than two levels.
+While this added granularity might seem beneficial, we find that extreme thresholds remain optimal for value maximization in symmetric populations, and that additional levels do not help a demand-maximizing platform either.
+Third, we analyze the case where receivers observe multiple independent recommendations from a symmetric population.
+Perhaps surprisingly, access to more signals does not systematically increase value or demand: receiver behavior with more than one signal often mirrors that under a well-calibrated single-recommendation system.
+Research on recommendations in the presence of preference heterogeneity is surprisingly scarce.
+An exception is Acemoglu et al.
+(2022), which argues that consumers with strong initial preferences for a product are more likely to purchase it so that the selection of recommenders is biased.
+Receivers should recognize this bias rather than accepting recommendations at face value.
+Contrary to Acemoglu et al.
+(2022), we focus on the determinants of the value and induced demand of recommendation systems and the design of those systems rather than the possibility and speed of learning.
+More broadly related, various papers deal with public goods problems in the provision of recommendations, which lead to insufficient exploration and reviewing, in particular for lesser-known niche products (e.g., Kremer et al., 2014, and Che and Hörner, 2018).
+[^5] Expanding on these papers, Vellodi (2022) considers the impact on market structure, showing how unbiased reviews might disadvantage entrants due to the “cold start” problem, thereby reinforcing the monopoly position of incumbents.
+These studies also investigate how appropriate information management can mitigate these public goods problems.
+While truthful recommendations are plausible in our setting, there are contexts where recommenders, firms, and platforms may have strategic incentives to bias recommendations.
+[^6] Bolton et al.
+(2013) have shown that eBay users often refrained from leaving negative feedback to avoid retaliation.
+Building on such concerns, Tadelis (2016) discusses how these strategic incentives shaped eBay’s platform design—for instance, the decision to disable seller evaluations of buyers—and explores broader implications for the credibility of recommendation and reputation systems.
+Like us, Chakraborty and Harbaugh (2010) assume that the sender possesses multidimensional information.
+They show how this helps her influence the behavior of the receiver in spite of a common bias.
+Bourreau and Gaudin (2022) consider incentives for firms to manipulate recommendations in a setting with heterogeneous products.
+They focus on how a platform wants to use recommendations to shift consumers to more desirable suppliers and how this affects pricing.
+Abstracting from product heterogeneity, Peitz and Sobolev (forthcoming) consider profit-maximizing firms that engage in sophisticated strategies to inflate their ratings.
+[^7] Johnen and Ng (2024) analyze a related phenomenon in a model without preference heterogeneity.
+They find that firms initially lower prices to attract positive ratings, later leveraging these ratings to justify price increases—effectively reducing the reliability of ratings as a quality indicator.
+This leads to a reduction in welfare due to inefficient product–consumer matches.
+The rating inflation is somewhat akin to our result that a low recommendation threshold can maximize demand and thus profits.
+However, in our model, this need not come at the expense of consumer welfare: when products and preferences are heterogeneous, the objectives of value and demand maximization can align, making low thresholds potentially optimal from both perspectives.
+Section 2 introduces the model.
+In Section 3, we describe the conditions under which different receiver types accept recommendations and use this to characterize the value and demand induced by a recommendation system, respectively.
+Then, we in turn consider symmetric (Section 4) and asymmetric (Section 5) populations to study the optimal design of recommendation systems, both in terms of creating value and inducing demand.
 Section 6 discusses our results and considers extensions.
@@
 All proofs are relegated to the appendix.
-2 Relation to the Literature
-Our paper investigates frictions in recommendation systems arising from preference heterogeneity, emphasizing the behavior of receivers.
-By contrast, the literature has mainly focused on frictions on the sender side.
-Public Goods Problems
-Research on recommendation systems often deals with the exploration-exploitation dilemma and its implications for providing recommendations, interpreted as the provision of a public good.
-Several authors argue that consumers usually neglect the positive externalities generated by their reviews, leading to insufficient exploration and reviewing, in particular for lesser-known niche products (e.g., Kremer et al.
-(2014) and Che and Hörner (2018)).
-Expanding on these papers, Vellodi (2022) considers the impact on market structure, showing how unbiased reviews might disadvantage entrants due to the “cold start” problem, thereby reinforcing the monopoly position of incumbents.
-These studies view the design of the recommendation system as the solution of a public goods problem by appropriate information management.
-[^2] For example, Kremer et al.
-(2014) and Che and Hörner (2018) suggest reducing transparency or endorsing yet-to-be-proven products, while Vellodi (2022) supports the idea of withholding reviews from incumbents to encourage new participants.
-Sender Selection
-The selection of recommenders is another source of frictions in recommendation systems.
-In their analysis of the learning dynamics resulting from recommendations, Acemoglu et al.
-(2022) argue that consumers with strong initial preferences for a product are more likely to purchase it.
-[^3] In contrast to our paper, their question is essentially whether and how fast consumers find out if the product is good or not, an assessment with which all consumers agree.
-Beyond that, we model preference heterogeneity differently.
-Recommendations thus inevitably carry an inherent bias reflecting the predispositions of the recommender pool.
-[^4] In Section 6.1, we allow for distinct sender and receiver populations and thus for the possibility of a bias among the recommenders relative to the overall population.
-Receivers should recognize this bias rather than accepting recommendations at face value.
-Our paper intersects with theirs by highlighting the significance of preference heterogeneity for interpreting recommendations.
-However, contrary to Acemoglu et al.
-(2022), we focus on the determinants of the value of recommendation systems and the design of those systems rather than the possibility and speed of learning.
-[^5] In Section 6.3, we introduce multiple recommendations, as do Acemoglu et al.
-(2022), but, in contrast to them, we ask under which conditions multiple recommendations enhance system value.
-Non-truthful Recommendations and Strategic Manipulations
-Strategic actions of recommenders, firms, and platforms may further reduce the value of recommendations for consumers.
-Bolton et al.
-(2013) have shown that eBay users often refrain from negative feedback to avoid retaliation.
-[^6] Tadelis (2016) discusses the design implications of these issues—using the example of eBay deciding to refrain from allowing seller evaluations of buyers—and also touches on strategic motives in recommendations more generally.
-Like us, Chakraborty and Harbaugh (2010) assume that the sender possesses multidimensional information.
-They show how this helps her influence the behavior of the receiver in spite of a common bias.
-Instead, we emphasize the effects of preference heterogeneity without strategic behavior, assuming that senders mechanically provide truthful recommendations.
-In contrast with our model with consumer recommendations, in Peitz and Sobolev (forthcoming) and Johnen and Ng (2024) firms engage in sophisticated strategies to inflate their ratings.
-[^7] Peitz and Sobolev (forthcoming) investigate firm-driven recommendations and the choice of sales channel: targeted (with recommendation) or non-targeted.
-They identify motives behind “inflated recommendations,” where the firm also recommends welfare-reducing bad product matches.
-Johnen and Ng (2024) analyze a related phenomenon, where firms initially lower prices to attract positive ratings, later leveraging these ratings to justify price increases—effectively reducing the reliability of ratings as a quality indicator.
-More closely related to our paper, Bourreau and Gaudin (2022) consider incentives for firms to manipulate recommendations in a setting with heterogeneous preferences.
 2 Model
-We begin by introducing the assumptions of our benchmark model in Section 3.1. Then, in Section 3.2, we discuss and motivate these assumptions.
-3.1 Assumptions
 Products and Payoffs
-We consider a consumer’s choice of a product from a set of alternatives $N\in\mathbb{N}$ with $N>1$.
+We consider a consumer’s choice between two products.
 We assume that each product is fully characterized by a two-dimensional quality vector, where $Q_{d}=1$ and $Q_{d}=0$ capture high and low quality, respectively, in dimension $d\in\{1,2\}$.
 Accordingly, four different product versions exist.
-We assume that the products are experience goods.
-Thus, the consumer cannot observe product quality before consumption, so the alternatives are identical for him ex ante.
+The consumer cannot observe product quality before consumption, so the alternatives are identical for him ex ante.
 However, he knows the prior distribution of the quality vectors, referred to as quality distribution for brevity.
@@
 $\Pr[(Q_{1},Q_{2})=(0,0)]$ $=q_{L},$ $\Pr[(Q_{1},Q_{2})=(0,1)]$ $=q_{2}.$
-We assume that $\mathbf{q}>0$ unless mentioned otherwise.
+We assume that $\mathbf{q}>0$ unless mentioned otherwise. We sometimes use the following self-explanatory terminology:
+Definition 1.
+The prevalence of controversial products is given as $Q:=\frac{q_{1}+q_{2}}{2}$. The odds of a good product are $\sigma:=q_{H}/q_{L}$.
 Recommendations and Updating
 We assume that one of the products comes with a recommendation.
-The receiver has to choose between this product and an alternative without a recommendation.
+The receiver has to choose between this product and the alternative without a recommendation.
 We distinguish between buy recommendations ($r=B$) and don’t-buy recommendations ($r=D$).
@@
 We assume that receivers are Bayesian, updating beliefs about product quality after having obtained a recommendation while taking into account the uncertainty about its informational content.
-We define $\pi^{B}(R)$ and $\pi^{D}(R)$ as the probability that a randomly chosen sender gives a buy or don’t-buy recommendation, respectively.
-Clearly,
+The probability that a randomly chosen sender gives a buy or don’t-buy recommendation, respectively is given as
 $\pi^{B}(R)$ $:=q_{H}+q_{1}\phi_{1}(R)+q_{2}\phi_{2}(R),$
-$\pi^{D}(R)$ $:=q_{1}(1-\phi_{1}(R))+q_{2}(1-\phi_{2}(R))+q_{L},$
-where $\pi^{D}(R)=1-\pi^{B}(R)$, and $\phi_{1}(R):=1-F(R-1/2)$ and $\phi_{2}(R):=F(1/2-R)$ denote the probability of a buy recommendation conditional on the product version being $(1,0)$ or $(0,1)$, respectively.
+$\pi^{D}(R)$ $:=1-\pi^{B}(R)=q_{1}(1-\phi_{1}(R))+q_{2}(1-\phi_{2}(R))+q_{L}.$
+Here $\phi_{1}(R):=1-F(R-1/2)$ and $\phi_{2}(R):=F(1/2-R)$ denote the probability of a buy recommendation conditional on the product version being $(1,0)$ or $(0,1)$, respectively.
 In words, the probability of receiving a buy recommendation from a randomly chosen sender is given by the sum of the probability that a product is objectively good and the probability that it is subjectively good enough from a randomly chosen sender’s perspective.
-Thus, accepting a buy recommendation the receiver’s posterior $\mathbf{p}^{B}(R):=(p_{H}^{B}(R),p_{1}^{B}(R),p_{2}^{B}(R),p_{L}^{B}(R))$ becomes
-$p_{H}^{B}(R):=\frac{q_{H}}{q_{H}+q_{1}\phi_{1}(R)+q_{2}\phi_{2}(R)},\quad\quad$ $p_{2}^{B}(R):=\frac{q_{2}\phi_{2}(R)}{q_{H}+q_{1}\phi_{1}(R)+q_{2}\phi_{2}(R)},$
-$p_{1}^{B}(R):=\frac{q_{1}\phi_{1}(R)}{q_{H}+q_{1}\phi_{1}(R)+q_{2}\phi_{2}(R)},\quad\quad$ $p_{L}^{B}(R):=0.$
-A don’t-buy recommendation gives rise to $\mathbf{p}^{D}(R):=(p_{H}^{D}(R),p_{1}^{D}(R),p_{2}^{D}(R),p_{L}^{D}(R))$, where
-$p_{H}^{D}(R):=0,\,$ $p_{2}^{D}(R):=\frac{q_{2}(1-\phi_{2}(R))}{q_{1}(1-\phi_{1}(R))+q_{2}(1-\phi_{2}(R))+q_{L}},$
-$p_{1}^{D}(R):=\frac{q_{1}(1-\phi_{1}(R))}{q_{1}(1-\phi_{1}(R))+q_{2}(1-\phi_{2}(R))+q_{L}},\,$ $p_{L}^{D}(R):=\frac{q_{L}}{q_{1}(1-\phi_{1}(R))+q_{2}(1-\phi_{2}(R)))+q_{L}}.$
-In the above expressions $\pi^{B}(R)$, $p^{B}(R)$, etc., we will typically drop the dependence on $R$ to simplify notation. Finally, we use the following summary terminology:
+In the appendix, we present the receiver’s posterior $\mathbf{p}^{r}(R):=(p_{H}^{r}(R),p_{1}^{r}(R),p_{2}^{r}(R),p_{L}^{r}(R))$ after a buy recommendation $r=B$ or a don’t-buy recommendation $r=D$ [^10] To ease notation, we will typically drop the dependence of these posteriors and other variables on $R$.
+Finally, we use the following summary terminology:
 Definition 2.
 $\mathcal{E}:=(\textbf{q},F)$ defines a decision environment. A recommendation system $\mathcal{R}$ consists of a decision environment $\mathcal{E}$ and a recommendation threshold $R\in(0,1)$.
-3.2 Discussion of the Framework
-Our theory of recommendations incorporates three key features: heterogeneity of products and preferences, coarseness of recommendations, and asymmetry of stakes between senders and receivers.
-We now discuss how our model addresses these and further points.
-Heterogeneity of Products and Preferences
-Without controversial products or preference heterogeneity, the analysis would become trivial.
-The case without controversial products would correspond to $q_{1}=q_{2}=0$.
-Then, for arbitrary $R\in(0,1),$ buy recommendations would reveal that the product is good, while don’t-buy recommendations would reveal that it is bad.
-Thus, there would be no friction in the recommendation system and no role for design considerations.
-With controversial products but absent preference heterogeneity, all consumers would have the same type $i^{*}$.
-Assuming w.l.o.g.
-that $i^{*}\geq 0$, senders would send a buy recommendation (i) only after consuming a good product if $R>\frac{1}{2}+i^{*}$, (ii) after consuming a good product or a controversial product of version $(1,0)$ if $\frac{1}{2}+i^{*}\geq R>\frac{1}{2}-i^{*}$, and (iii) after consuming the good product or any of the controversial products if $\frac{1}{2}-i^{*}\geq R$.
-The first-best could be induced by the appropriate choice of the recommendation threshold, taking into account how the payoffs from controversial products compare to the expected payoff from the outside option, In sum, a meaningful analysis requires controversial products as well as preference heterogeneity.
-Coarse Information Transmission
-With a single threshold above and below which senders give buy and don’t-buy recommendations, respectively, it is impossible to always reveal the product version.
-Thus, as desired, recommendations provide coarse information.
-If, in contrast, the sender could give four different ratings depending on the consumption experience, full revelation would be possible in our parsimonious model with only four product versions.
-In reality, recommendations can convey more content than whether a buyer liked a product or not, for instance, by including verbal descriptions.
-However, given the typical complexity of products, even relatively detailed recommendations will usually only transmit coarse information.
-Asymmetry of Stakes
-To capture the sender’s limited stakes in the process, we assume that she behaves truthfully, mechanically reporting whether her consumption experience was good enough or not.
-More sophisticated and potentially strategic recommendations seem at odds with the lack of incentives to invest in this reporting activity, which results in the underprovision of recommendations (see the literature review).
-Recommendation Threshold
-Our preferred way of thinking about the recommendation threshold is as a design parameter.
-To motivate this, we rely on Decker (2022) who argues that Airbnb advises its customers to give a five-star rating to every host that provides reasonably good services, thus reserving lower ratings exclusively to very bad hosts.
-[^10] In our setting, this approach corresponds to the use of a coarse recommendation system, with the vast majority of hosts receiving a buy recommendation and a small subset receiving a don’t-buy recommendation.
-The recommendation threshold would then be close to $R=0$ so that products with a don’t-buy recommendation are typically objectively bad.
-In this spirit, we think of the threshold as reflecting the guidance of a platform such as Airbnb or as a social norm that is being actively shaped.
-Other Recommendation Systems
-Assuming that senders mechanically provide a binary recommendation is adequate in view of their limited willingness to invest more time into giving a recommendation.
-Nonetheless, one could imagine more sophisticated recommendation systems.
-For instance, the sender might recommend a product when it is either good or when it is controversial but preferred by most of the population.
-This, however, presupposes a sender with detailed information about the receiver’s decision environment.
-Instead, the sender could give a buy recommendation when the average payoff she expects receivers to obtain from consuming the product in question is higher than for the outside option.
-In that case, the expected value of the outside option would determine the recommendation threshold, which would then no longer be a design parameter.
-Product Ranking
-When $q_{H}>0$ and $q_{L}>0$, all types agree on which product version is the best and which is the worst because their utility is fully determined by the qualities in each dimension and is increasing in each of them.
-In other environments, no products are objectively the best or the worst.
-For instance, there would be no such products in a modified version of our setting where $q_{H}=0$ and $q_{L}=0$, as the valuations for $(1,0)$ and $(0,1)$ are perfectly negatively correlated.
-We nonetheless keep the specific feature of the model because we believe it covers a wide range of plausible scenarios.
-4 Optimal Receiver Behavior
-In this section, we ask under which circumstances a receiver will accept all recommendations, depending on his type $i$ and the recommendation system $\mathcal{R}$, as introduced in Definition 1.
-In Section 4.1, we discuss through which channels recommendations affect beliefs, distinguishing between the objective and subjective content of the recommendations.
-Section 4.2 then characterizes the optimal behavior of the receiver by identifying thresholds which divide the type set into receivers that accept all recommendations and those that never do.
-4.1 The Effects of Recommendations on Beliefs
-Recommendations contain coarse information about the product under consideration, leading the receiver to update his beliefs about its quality.
-We first focus on buy recommendations, the total effect of which can be decomposed as follows: [^11] In Appendix A.1, we provide a more formal analysis, decomposing the difference between posteriors and priors into three corresponding components.
-  1. (I)
-    The probability of a bad product falls to zero.
-  2. (II)
-    A good product becomes more likely relative to the controversial products.
-  3. (III)
-    The odds between the controversial products change.
-The first two effects in this belief-updating decomposition are objectively positive: They correspond to increases of the expected payoff for every receiver type, albeit of different sizes.
-By contrast, the third effect is fully type-dependent and thus subjective, resulting in an increase in the payoff for some receivers and a decrease for others.
-The effects of don’t-buy recommendations are analogous to those of buy recommendations: First, the probability of a good product falls to zero; second, bad products become more likely relative to controversial products, and third, the odds between the controversial products change.
-Thus, again, the recommendation has two objective effects (but these are both bad now) and a subjective effect (which can be good or bad, depending on the receiver’s type).
-4.2 Characterizing Optimal Receiver Behavior
-We are now in a position to understand whether it is optimal for the receiver to accept recommendations.
-To this end, note that, after receiving a recommendation, the receiver can either buy the recommended product or not buy it.
+3 Preliminaries: Receiver Behavior, Value and Demand
+In this section, we begin by asking under which circumstances a receiver will accept recommendations, depending on his type $i$ and the recommendation system $\mathcal{R}$.
+Then, we will use this result to characterize the value of a recommendation system and determine the demand induced by the recommendation system.
+3.1 Receiver Behavior
 Denote by $U^{B}(i)$ and $U^{D}(i)$ player $i$’s expected payoff of buying a product with a buy and don’t-buy recommendation, respectively, and by $U^{0}(i)$ the expected payoff of not buying, i.e., of buying the alternative product with no recommendation.
@@
 $U^{0}(i)$ $:=q_{H}+(1/2+i)q_{1}+(1/2-i)q_{2}.$
-Now, the difference between $U^{r}(i)$ and $U^{0}(i)$ determines the behavior of an expected-payoff maximizing receiver as follows.
-Definition 2.
-We say that a receiver $i$
+Receiver $i$ will accept a buy recommendation if $U^{B}(i)\geq U^{0}(i)$ or equivalently
+$p_{H}^{B}-q_{H}+\frac{p_{1}^{B}-q_{1}}{2}+\frac{p_{2}^{B}-q_{2}}{2}\geq i[(p_{2}^{B}-q_{2})-(p_{1}^{B}-q_{1})].$ (1)
+Intuitively, a buy recommendation conveys the following information to the receiver:
   1. (I)
-    accepts a buy recommendation if $U^{B}(i)\geq U^{0}(i)$; otherwise we say that he does not accept a buy recommendation;
+    The probability of a bad product has fallen to zero.
   2. (II)
-    accepts a don’t-buy recommendation if $U^{0}(i)\geq U^{D}(i)$; otherwise we say that he does not accept a don’t-buy recommendation
+    The likelihood of a good product relative to the controversial products has increased.
   3. (III)
-    accepts all recommendations if he accepts both buy and don’t-buy recommendations;
-  4. (iv)
-    never accepts recommendations if he does not accept either buy and don’t-buy recommendations.
-Hence, a receiver who accepts all recommendations “does as he is told”, whereas a receiver who never accepts recommendations is a “contrarian” who does the opposite of what he is told.
-[^12] The use of the term aligns with Drehmann et al.
-(2005), who document contrarian behavior, i.e., investors going against the behavior of the predecessors, in an financial-markets experiment.
-Note that the above definition does not cover all conceivable types of behavior, as a receiver $i$ could accept buy recommendations and not accept don’t-buy recommendations, which is tantamount to always buying the recommended product and thus ignoring the content of the recommendations.
-Analogously, a receiver $i$ could ignore the content of the recommendations and never buy the recommended product.
-We will return to these cases shortly.
-The condition from Definition 2(i) for receiver $i$ to accept a buy recommendation can equivalently be written as
-$p_{H}^{B}-q_{H}+\frac{p_{1}^{B}-q_{1}}{2}+\frac{p_{2}^{B}-q_{2}}{2}\geq i[(p_{2}^{B}-q_{2})-(p_{1}^{B}-q_{1})].$ (1)
+    The odds between the controversial products has changed.
+Similarly, the receiver will accept a don’t-buy recommendation if $U^{0}(i)\geq U^{D}(i)$, and the effects of these recommendations are analogous.
 For more compact formulations, we introduce the following terminology.
@@
 […]
+For the unbiased receiver $i=0$, equation (4) simplifies to $\Delta_{O}^{B}\geq 0$.
+As this receiver does not care about the shift in probabilities between the two controversial versions of the products, $\Delta_{O}^{B}\geq 0$ reflects the two objective effects (I) and (II).
+For $i\neq 0$, the interpretation of $\Delta_{O}^{B}\geq 0$ is more nuanced.
+If $q_{1}=q_{2}$, the expected payoff of a controversial product is type-independent, ensuring that $\Delta_{O}^{B}$ fully captures the objective benefits of a buy recommendation.
+If $q_{1}\neq q_{2}$, the size of these benefits depends on the receiver’s type.
+[^12] Intuitively, if one controversial version (e.g., $(1,0)$) is ex ante more likely, learning that a product is controversial rather than bad is more valuable for those who prefer $(1,0)$ over $(0,1)$, and vice versa for the opposite preference.
+Similarly, learning that a product is good rather than controversial is more valuable for those who favor the less likely version.
+Nonetheless, the sign of $\Delta_{O}^{B}$ remains a reliable indicator of the objective effect, as all types agree that (I) and (II) are beneficial.
+We can now state our first result.
 Proposition 1.
@@
   1. (i)
-    either accepts all recommendations or never accepts recommendations;
+    accepts buy recommendations if and only if he accepts don’t-buy recommendations;
   2. (ii)
-    accepts all recommendations if and only if $\Delta_{O}^{B}\geq i\Delta_{S}^{B}$.
-The first part of the result states that Definition 2 indeed captures all relevant forms of behavior.
-That is, no type of receiver will ignore the informational content of recommendations by always or never buying the recommended product.
-Intuitively, a receiver accepts all recommendations if he thinks that they have sufficient objective content and/or that his preferences are sufficiently aligned with the sender’s.
-Conversely, a receiver never accepts recommendations if he thinks that they do not have sufficient objective content and/or that his preferences are not sufficiently aligned with the sender’s.
+    accepts (buy and don’t-buy) recommendations if and only if $\Delta_{O}^{B}\geq i\Delta_{S}^{B}$.
+Intuitively, a receiver accepts recommendations if he thinks that they have sufficient objective content and/or that his preferences are sufficiently aligned with the sender’s.
+Conversely, a receiver does not accept recommendations if he thinks that they do not have sufficient objective content and/or that his preferences are not sufficiently aligned with the sender’s.
 Whether the recommendation is positive or negative does not matter for this assessment.
-Rather, as laid out in the decomposition in Section 4.1, the objective effect $\Delta_{O}^{B}$ of a buy recommendation is positive, as the latter makes a good product more likely relative to the controversial products and rules out a bad one.
-By contrast, the sign of the subjective effect $\Delta_{S}^{B}$ can be positive or negative.
-The second part of the result shows which types accept all recommendations, which happens unless the subjective effect is sufficiently large and the receiver’s preferences are not sufficiently well aligned with those of the rest of the population.
-We emphasize that receivers who never accept recommendations still make use of them, by doing the opposite of what they are told.
-Proposition 1 tells us that we can decompose the population into at most two different groups: those who accept all and those who never accept recommendations.
-Denoting the type who is indifferent between always accepting and never accepting recommendations as
+Denoting the type who is indifferent between accepting and not accepting recommendations as
 $\tilde{i}:=$ $\frac{\Delta_{O}^{B}}{\Delta_{S}^{B}}\text{ if }\Delta_{S}^{B}\neq 0\text{,}$
@@
   1. (i)
-    If $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$, then all $i\in[-1/2,1/2]$ accept all recommendations.
+    If $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$, then all $i\in[-1/2,1/2]$ accept recommendations.
   2. (ii)
-    If $\Delta_{S}^{B}<-2\Delta_{O}^{B}$, then all $i\geq\tilde{i}$ accept all recommendations.
+    If $\Delta_{S}^{B}<-2\Delta_{O}^{B}$, then all $i\geq\tilde{i}$ accept recommendations.
   3. (iii)
-    If $\Delta_{S}^{B}>2\Delta_{O}^{B}$, then all $i\leq\tilde{i}$ accept all recommendations.
+    If $\Delta_{S}^{B}>2\Delta_{O}^{B}$, then all $i\leq\tilde{i}$ accept recommendations.
 The intuition for Corollary 1, the proof of which is omitted, is straightforward.
-Without loss, consider a buy recommendation.
-For $\Delta_{S}^{B}=0$, the recommendation does not shift probability between the controversial product versions.
-Accordingly, accepting the recommendation is optimal for all players because of its positive objective effect $\Delta_{O}^{B}\geq 0$.
-For $\Delta_{S}^{B}\neq 0$, the conclusion is the same as long as $|\Delta_{S}^{B}|$ remains small enough—as required by (i).
-In that case, even the receivers who are most adversely affected by the subjective content of the recommendation (those at $i=1/2$ or $i=-1/2$) are convinced by its relatively objective character.
-As $|\Delta_{S}^{B}|$ increases, the role of the probability shift becomes apparent.
-As types $i>0$ prefer (1,0) over (0,1), a shift of the probability from the latter to the former induced by a recommendation ($\Delta_{S}^{B}<0$) increases such receivers’ expected payoff, reinforcing the objective effect.
-By contrast, such a shift reduces the expected payoff of types $i<0$.
-The condition in (ii) always holds for $i>0$, as the objective and subjective effects reinforce each other.
-It also holds for $i<0$ as long as $|i|<|\tilde{i}|$ so that the adverse subjective effect does not dominate the objective effect.
-Similarly, when the recommendation shifts the probability from (1,0) to (0,1) (when $\Delta_{S}^{B}>0$), all types $i<0$ accept.
-By contrast, types $i>0$ experience a reduction in their value, which can outweigh the positive objective effect.
-The condition in (iii) makes sure that the latter case does not arise.
-[^15] The effect on the expected payoffs in cases (ii) and (iii) is similar to the effect of a demand rotation as discussed in Johnson and Myatt (2006): For the types who accept the recommendation, the expected payoffs increase; for those who do not, they fall.
-Corollary 1 may potentially be useful to draw inferences about the relative importance of the objective and subjective effects of recommendations: If an observer sees a large share of receivers accepting recommendations, this can be interpreted as evidence for a relatively high objective effect of the recommendation.
-Finally, we analyze receiver behavior for extreme recommendation thresholds.
+As long as the subjective content of the recommendation is small (case (i)), even the most extreme receivers satisfy the cut-off condition of Proposition 1(ii).
+As the subjective content increases, consumers who are not sufficiently aligned with the population at large will reject recommendations (cases (ii) and (iii)).
+The interplay between objective and subjective effect leads to a stark result for extreme recommendation thresholds:
 Proposition 2.
-Suppose $R\rightarrow 1$ or $R\rightarrow 0$. Then, all types accept all recommendations irrespective of the product distribution.
-In the limits, both types of recommendations essentially contain only objective information.
-For instance, when $R\rightarrow 1$, buy recommendations require that the product is objectively good, while don’t-buy recommendations require that it is not.
-Either way, the recommendation does not reveal anything about the relative probabilities of the two controversial product versions.
-Instead, it provides objective information that is equally valuable for all types, irrespective of their preferences for controversial products.
-Thus, all types accept all recommendations.
-5 The Value of the Recommendation System
-We now deal with the value of recommendation systems.
-In Section 5.1, we provide a characterization result.
-Section 5.2 asks how to use the recommendation threshold to influence the value.
-Finally, we investigate the effects of the preference and quality distributions on the value in Section 5.3.
-5.1 Characterizing the Value of the Recommendation System
-Before defining the value of a recommendation system, we describe the expected payoffs of a fixed type $i$ receiver, depending on whether he has access to a recommendation system and whether he accepts its recommendations.
+Suppose $R\rightarrow 1$ or $R\rightarrow 0$. Then, all types accept recommendations irrespective of the product distribution.
+In the limits, recommendations contain only objective information.
+As $R\to 1$, a buy recommendation guarantees the product is good, while a don’t-buy recommendation implies it is not.
+Since no information is revealed about the relative likelihood of controversial products, all types value the recommendation equally and accept it.
+Next, we discuss how the analysis would become trivial in the absence of either controversial products or preference heterogeneity.
+If there were no controversial products ($q_{1}=q_{2}=0$), recommendations would simply reveal whether the product is good or bad.
+In this case, recommendations would contain only objective information, eliminating any friction in the system.
+If there was no preference heterogeneity, the sender’s recommendations would convey objective information about whether the product is good, bad, or controversial.
+Without heterogeneity in receiver preferences, the subjective effect of a recommendation—how it shifts beliefs between different controversial product versions—would be irrelevant, as all receivers would interpret the information in the same way.
+As a result, the sender could fully align recommendations with the receivers’ interests by adjusting the recommendation threshold, effectively eliminating any tension.
+Hence, it is the interplay between controversial products or preference heterogeneity that makes the design of the recommendation system nontrivial and interesting.
+3.2 Value of the Recommendation System
+Having determined receiver behavior, we can now characterize the value of the recommendation system.
 Without a recommendation system, all products are identical ex ante, with an expected payoff based on the prior distribution of product qualities.
@@
 Now suppose one of the products comes with a buy or don’t-buy recommendation, so the receiver has to update his belief about the quality distribution.
-A receiver who accepts all recommendations chooses this product (rather than one of the alternatives) only if it comes with a buy recommendation.
-With a recommendation system, the expected payoff of a receiver $i$ who accepts all recommendations is thus
+A receiver who accepts recommendations chooses this product (rather than one of the alternatives) only if it comes with a buy recommendation.
+With a recommendation system, the expected payoff of a receiver $i$ who accepts recommendations is thus
 $V^{A}(i):=\pi^{B}U^{B}(i)+(1-\pi^{B})U^{0}(i).$ (6)
-The first term is the probability of receiving a buy recommendation multiplied by the expected payoff from accepting it (and thus buying the recommended product).
-The second term is the probability of a don’t-buy recommendation multiplied by the expected payoff from buying the product without any recommendation, that is, based on priors.
-[^16] The notation $V^{A}(i)$ emphasizes the dependence of this expected payoff on the type of the receiver.
-Clearly, $V_{A}$ also depends on $R$ and on the decision environment; similarly for the expressions below.
-A receiver who never accepts recommendations will buy the recommended product if it comes with a don’t-buy recommendation.
+The first term on the right-hand side comes from buy recommendations, the second one from don’t-buy recommendations.
+A receiver who does not accept recommendations will buy the recommended product if it comes with a don’t-buy recommendation.
 Otherwise, he will buy one of the other products.
@@
 A buy recommendation does not create value for receivers who do not accept it, as such types choose the alternative for which no recommendation is available.
-Instead, the don’t-buy recommendation, which is given with probability $(1-\pi^{B})$, does influence the behavior of the types who do not accept it: Instead of buying an alternative based on prior probabilities, updating leads them to choose the product for which they received the explicit recommendation not to buy.
-We now define the value of a recommendation system. [^17] This definition is in line with the definition of a valuable signal in Kamenica and Gentzkow (2011).
+Instead, the don’t-buy recommendation, which is given with probability $(1-\pi^{B})$, does influence the behavior of the types who do not accept it: Rather than buying an alternative based on prior probabilities, updating leads them to choose the product for which they received the explicit recommendation not to buy.
+We now define the value of a recommendation system in terms of the expected welfare increase it creates for receivers.
+[^13] This definition is in line with the definition of a valuable signal in Kamenica and Gentzkow (2011).
 Definition 4.
@@
 vertical space
-The simple structure of $V(\mathcal{R})$ for $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$ in (i) results because, by Corollary 1, all types accept a recommendation.
-The difference between the expressions for $\Delta_{S}^{B}<-2\Delta_{O}^{B}$ in (ii) and $\Delta_{S}^{B}>2\Delta_{O}^{B}$ in (iii) arises because only types $i>\tilde{i}$ accept the recommendation in the former case, whereas only types $i<\tilde{i}$ do so in the latter.
-We now express the value $V(\mathcal{R})$ in terms of its objective and subjective content.
+Importantly, the definition captures the idea that the increase in expected utility is measured before drawing the sender-receiver pair.
+Part (i) relies on Corollary 1, which implies that all receivers accept recommendations when $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$, as the objective effect dominates.
+Hence, by equation (6), the recommendation system creates value only in case of a buy recommendation, as a don’t-buy recommendation leads to the same expected payoff as without a recommendation system.
+Parts (ii) and (iii) refer to cases where the subjective content of the messages matters more.
+Then, not all receivers accept the recommendation, and the system may create value for both receivers who accept recommendations as well as those who do not.
+We restate the value $V(\mathcal{R})$ to better understand how it reflects the role of the objective and subjective content of recommendations:
 Proposition 3.
-The value $V(\mathcal{R})$ of the recommendation system $\mathcal{R}$ is:
+The value $V(R)$ of the recommendation system $\mathcal{R}$ is: [^14] As we focus on $R$ as the design parameter, we often write $V(R)$ instead of $V(\mathcal{R})$.
 vertical space
 (i) $\pi^{B}[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i]]$ if $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$; (ii) $(1-\pi^{B})F(\tilde{i})\left[\Delta_{O}^{D}-\Delta_{S}^{D}\mathbb{E}[i\mid i\leq\tilde{i}]\right]+\pi^{B}(1-F(\tilde{i}))\left[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i\mid i\geq\tilde{i}]\right]$ if $\Delta_{S}^{B}<-2\Delta_{O}^{B}$; (iii) $\pi^{B}F(\tilde{i})\left[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i\mid i\leq\tilde{i}]\right]+(1-\pi^{B})(1-F(\tilde{i}))\left[\Delta_{O}^{D}-\Delta_{S}^{D}\mathbb{E}[i\mid i\geq\tilde{i}]\right]$ if $\Delta_{S}^{B}>2\Delta_{O}^{B}$.
-Intuitively, when $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$, Corollary 1 implies that all receivers accept all recommendations as the objective effect dominates.
-Hence, by equation (6), the recommendation system creates value only in case of a buy recommendation, as a don’t-buy recommendation leads to the same expected payoff as without a recommendation system.
-The positive payoff effect of the recommendation reflects the type-independent objective part $\Delta_{O}^{B}$ and the subjective part $\Delta_{S}^{B}$, which is weighted by the average type.
-Finally, all of that only materializes with the probability of a buy recommendation, $\pi_{B}$.
-As the absolute value of $\Delta_{S}^{B}$ increases, the analysis becomes more complex because some types never accept recommendations.
-We consider the case $\Delta_{S}^{B}>2\Delta_{O}^{B}$ and reproduce the expression for $V(\mathcal{R})$ as follows: [^18] The argument for $\Delta_{S}^{B}<-2\Delta_{O}^{B}$ is analogous.
-$\underbrace{\pi^{B}F(\tilde{i})\left[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i\mid i\leq\tilde{i}]\right]}_{\text{types accept all recommendations}}+\underbrace{(1-\pi^{B})(1-F(\tilde{i}))\left[\Delta_{O}^{D}-\Delta_{S}^{D}\mathbb{E}[i\mid i\geq\tilde{i}]\right]}_{\text{types never accept recommendations}}.$ (8)
-In this case, the buy recommendation shifts probability from $(1,0)$ to $(0,1)$, which is unfavorable for receivers $i>0$, reducing their gains from accepting and possibly making them negative.
-Thus, for $\Delta_{S}^{B}>2\Delta_{O}^{B}$, only types below $\tilde{i}$ (a fraction $F(\tilde{i})$ of the receivers) accept all recommendations, which comes with probability $\pi^{B}$.
-In addition to the objective value of the recommendation ($\Delta_{O}^{B}$), these receivers obtain a subjective value.
-In (8), the subjective value is reflected in $-\Delta_{S}^{B}\mathbb{E}[i\mid i\leq\tilde{i}]$, the average effect of the shift in the relative probabilities of $(1,0)$ and $(0,1)$ on the value of accepting the recommendation for these buyers.
-The sign of this term depends on the properties of $F$ captured in $\mathbb{E}[i\mid i\leq\tilde{i}]$.
-It is positive if the negative contribution of types $i<0$ dominates the positive contribution of types $i>0$.
-The second component in (8), corresponding to the types who never accept recommendations, is arguably more surprising.
-Given their lack of alignment with the rest of the population, these types treat the recommendation not to buy as good news about the product and thus do not accept the recommendation.
-[^19] In detail, a don’t-buy recommendation (which happens with probability $1-\pi^{B}$) carries an adverse objective content (summarized by $\Delta_{O}^{D}$ in the term in squared brackets).
-However, for receivers with $i>0$ it also carries the good news that the product is more likely to be of the preferred version $(1,0)$ rather than $(0,1)$.
-Those with $i>\tilde{i}(>0)$ then buy the product despite the adverse objective content, valuing the beneficial subjective effect with $-\Delta_{S}^{D}\mathbb{E}[i\mid i\geq\tilde{i}]>0$ on average.
-Such receivers exist only if $\Delta_{S}^{B}$ is sufficiently large (and hence $\Delta_{S}^{D}$ is sufficiently negative relative to $\Delta_{O}^{D}$).
-Nonetheless, its informational content creates value for them.
-Going forward, we shall often invoke the following symmetry assumption on the preference distribution.
+The positive payoff effect of the recommendation in case (i) reflects the type-independent objective part $\Delta_{O}^{B}$ and the subjective part $\Delta_{S}^{B}$, which is weighted by the average type.
+As the subjective component $\Delta_{S}^{B}$ becomes stronger in absolute terms, not all receiver types will follow the recommendation anymore, and the analysis becomes more nuanced.
+To illustrate this, consider case (iii) where $\Delta_{S}^{B}>2\Delta_{O}^{B}$.
+In this case, a buy recommendation shifts beliefs toward the controversial product version that is less preferred by receivers $i>0$, which implies that they may decide not to follow the recommendation.
+Only those whose preferences are sufficiently aligned with the sender’s—specifically, those with types below a certain threshold $\tilde{i}$—still choose to buy.
+These types represent a fraction $F(\tilde{i})$ of the population.
+For them, the recommendation carries both objective value (because the product is likely better than the outside option) and subjective value (because the controversial version it signals is relatively well aligned with their preferences).
+This subjective value is captured by the average type among those who follow the recommendation, $\mathbb{E}[i\mid i\leq\tilde{i}]$.
+Even those types who reject the buy recommendation may benefit.
+Types above $\tilde{i}$, who choose not to follow the recommendation, infer that the product is likely of the version they personally dislike—yet this very inference tells them something useful.
+When these same types receive a don’t-buy recommendation, they update in favor of the controversial product version they prefer, and may end up buying despite, or, rather, because of the negative signal.
+So, while the objective signal says “don’t buy”, the subjective shift in beliefs may outweigh it for some types.
+As a result, even rejected recommendations can create value, precisely because they help those whose preferences diverge from the majority make more informed choices.
+3.3 Demand for the Recommended Product
+A platform does not necessarily design a recommendation system to maximize the value it creates for consumers.
+Rather, the objective may be to maximize profits, which in some circumstances amounts to maximizing demand.
+[^15] For instance, assuming that a platform has committed to a fixed transaction fee, the profit-maximizing platform will want to to adjust its recommendation system to maximize demand.
+Whether a receiver buys the recommended product depends on the content of the recommendation and whether the receiver accepts or rejects the recommendation.
+Therefore, demand may comprise receivers who accept and receivers who reject recommendations.
+Using Corollary 1, the demand for the recommended product $D(R)$ can be expressed analogously to the value of the recommendation system in Proposition 3 as
+$D(R)=\begin{cases}\pi_{B}(R)&\text{if }|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}\\ \pi_{B}(R)F(\tilde{i}(R))+(1-\pi_{B}(R))(1-F(\tilde{i}(R)))&\text{if }\Delta_{S}^{B}>2\Delta_{O}^{B},\\ \pi_{B}(R)(1-F(\tilde{i}(R)))+(1-\pi_{B}(R))F(\tilde{i}(R))&\text{if }\Delta_{S}^{B}<-2\Delta_{O}^{B},\end{cases}$ (8)
+where we have made explicit that the indifferent receiver $\tilde{i}$ and the probability of a buy recommendation $\pi_{B}$ are functions of the threshold $R$.
+As for the value above, the relation between $\Delta_{S}^{B}$ and $\Delta_{O}^{B}$ determines which types accept a recommendation.
+In the first case, all types accept a recommendation so that demand is given by the probability of a buy recommendation $\pi_{B}(R)$.
+In the other cases, demand comprises both types that accept a recommendation and types that do not.
+Importantly, types that do not accept recommendations buy the recommended product following a don’t-buy recommendation, thus contributing to demand.
+4 Symmetric Populations
+In this section, we consider populations that are symmetrically distributed.
+Hence, there is no systematic preference bias for either of the controversial products in the population, so the preference heterogeneity reflects pure noise.
+We will first consider value-maximizing recommendation systems.
+Then, we turn to demand-maximizing recommendation systems.
+Throughout this section, we invoke the following symmetry assumption on the preference distribution.
 Assumption 1.
 $F(-i)=1-F(i)$ for all $i\in[-1/2,1/2]$.
-We now define the quantity
-$\beta:=F(1/2-R).$
-Due to Assumption 1, it has a simple interpretation: Given a threshold $R$, the parameter $\beta$ captures the fraction of the population that is willing to recommend product versions $(1,0)$ or, equivalently (given the symmetry of the distribution), $(0,1)$.
-For a fixed type distribution $F$, the variables $\beta$ and $R$ are inversely related.
-We then obtain the following result:
+4.1 Value-Maximizing Recommendation Systems
+Under Assumption 1, the value of the recommendation system can be expressed particularly concisely, as evidenced by the following result:
 Proposition 4.
-Suppose Assumption 1 holds. Then, all types accept all recommendations and
-$V(\beta)=\pi^{B}\Delta_{O}^{B}\text{.}$
-To understand why all types accept all recommendations, the belief-updating decomposition in Section 4.1 is helpful.
-Its third component captures the potential shift of probabilities between controversial products, which induces a purely type-dependent payoff effect.
-Given the symmetry assumption, both controversial product versions lead to a buy recommendation with probability $\beta$.
+If Assumption 1 holds, all types accept recommendations and
+$V(R)=\pi^{B}\Delta_{O}^{B}\text{.}$
+Intuitively, given the symmetry assumption, both controversial product versions lead to a buy recommendation with the same probability.
 Hence, a recommendation does not shift probabilities between the controversial products.
-As the other two effects of a buy recommendation on payoffs are positive for all types, the buy recommendation increases expected payoffs for any type, yielding the result.
-[^20] An analogous argument applies for don’t-buy recommendations.
+As the other two (objective) effects of a buy recommendation on payoffs are positive for all types, the buy recommendation increases expected payoffs for any type, so that all types accept recommendations.
 Proposition 3 thus implies $V(R)=\pi^{B}\left[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i]\right]$.
 Further, because of the symmetry of the population, $\mathbb{E}[i]=0$, so $V(R)=\pi^{B}\Delta_{O}^{B}$ as required.
-5.2 Optimal Design of Recommendation Systems
 We now use our characterization of the value of a recommendation system to address design issues, focusing on the optimal choice of the threshold.
-We start with the case of a symmetric population, then we treat asymmetric populations.
-5.2.1 Symmetric Type Distribution
-In the symmetric setting, we obtain a simple expression for $V(R)$, yielding transparent comparative statics results.
-[^21] As we focus on $R$ as the design parameter, we often write $V(R)$ instead of $V(\mathcal{R})$.
-To this end, a reparametrization is helpful:
-Definition 5.
-The prevalence of controversial products is given as $Q:=\frac{q_{1}+q_{2}}{2}$. The odds of a good product are $\sigma:=q_{H}/q_{L}$.
-Using these parameters, we immediately obtain:
-Lemma 1.
-Suppose Assumption 1 holds. Then,
-$\pi^{B}$ $=\frac{(1-2Q)\sigma}{1+\sigma}+2Q\beta;$
-$\Delta_{O}^{B}$ $=\frac{\beta Q(\sigma+1)-2Q\sigma+\sigma}{2\beta Q(\sigma+1)-2Q\sigma+\sigma}-\frac{(1-2Q)\sigma}{1+\sigma}-Q;$
-$V(\beta)$ $=\pi^{B}\Delta_{O}^{B}=(1-2Q)\frac{\sigma+Q\left(\beta-\sigma+\sigma^{2}(1-\beta)\right)}{(\sigma+1)^{2}}.$
-The value of the recommendation system depends on $\beta$ as well as on the prevalence of controversial products $Q$ and on the odds ratio $\sigma$ between good and bad products.
-Recall that, for each of the two controversial product versions, $\beta$ measures the fraction of the population willing to recommend it.
-Therefore, the parameter captures the influence of the distribution and the recommendation threshold on the value.
-[^22] Observe that by varying $R$ we can obtain any $\beta\in(0,1)$.
-We can interpret changes of $\beta$ as either changes of the threshold $R$ for a given distribution $F$ or as changes of the distribution $F$ for a given threshold $R$.
-The following result abstracts from this distinction, allowing for either interpretation.
-It shows that countervailing effects of $\beta$ on the probability of a buy recommendation and the objective value $\Delta_{O}^{B}$ of accepting it result in a non-trivial effect on $V(\beta)=\pi^{B}\Delta_{O}^{B}$.
-Lemma 2.
-Suppose Assumption 1 holds. Then,
-  • (i)
-    $\pi^{B}$ is strictly increasing and $\Delta_{O}^{B}$ strictly decreasing in $\beta$;
-  • (ii)
-    $\frac{\partial^{2}V}{\partial\beta\partial\sigma}<0$ and $V^{\prime}(\beta)=\begin{cases}>0&\text{if }\sigma<1;\\ =0&\text{if }\sigma=1;\\ <0&\text{if }\sigma>1.\end{cases}$
-The first statement in (i) holds because as $\beta$ increases a greater share of the population is willing to recommend a controversial product.
-Moreover, conditional on receiving a buy recommendation, the recommended product is thus more likely to be controversial.
-Therefore, the objective payoff component $\Delta_{O}^{B}$ decreases as $\beta$ increases, implying the second statement in (i).
-The effect of $\beta$ on $V(\beta)$ in (ii) reflects the effect on its two components.
-As $\sigma$ increases (good products become relatively more likely than bad ones), a buy recommendation becomes more likely.
-Hence, a high $\Delta_{O}^{B}$ is particularly valuable when $\sigma$ is high so that increasing $\sigma$ and lowering $\beta$ are complementary ($\frac{\partial^{2}V}{\partial\beta\partial\sigma}<0$).
-[^23] More technically, repeated application of the product rule shows that $\frac{\partial^{2}V}{\partial\beta\partial\sigma}=\frac{\partial\pi^{B}}{\partial\beta}\frac{\partial\Delta_{O}^{B}}{\partial\sigma}+\frac{\partial\pi^{B}}{\partial\sigma}\frac{\partial\Delta_{O}^{B}}{\partial\beta}+\frac{\partial^{2}\pi^{B}}{\partial\beta\partial\sigma}\Delta_{O}^{B}+\frac{\partial^{2}\Delta_{O}^{B}}{\partial\beta\partial\sigma}\pi^{B}$.
-The (incomplete) intuition just given corresponds to the first term being positive.
-The value-maximizing level of $\beta$ must thus be weakly decreasing in $\sigma$.
-The symmetry assumption turns out to render the problem linear in $\beta$, so the second result in (ii) follows.
-Using $\beta=F(1/2-R)$, Lemma 2 directly implies the following result.
-[^24] Result (iii) follows because $V(\beta)$ is independent of $\beta$ in this case.
 Proposition 5.
@@
     constant in $R$ if $\sigma=1$.
-By Lemma 2, an increase in $R$ reduces the probability of a buy recommendation, while it increases $\Delta_{O}^{B}$, the expected average payoff conditional on a buy recommendation.
-These effects exactly offset each other if and only if $\sigma=1$.
-Otherwise, one effect dominates the other.
-Based purely on uncontroversial products, a buy recommendation would be more likely than a don’t-buy recommendation when the good product is relatively more likely than the bad one ($\sigma>1$).
-It is thus worthwhile to increase the expected average payoff conditional on a buy recommendation by increasing $R$ at the expense of the probability of receiving such a recommendation.
-Put differently, the designer chooses between revealing whether the product is good or whether it is bad.
-He moves the threshold in the direction which yields full revelation with higher probability, which hinges only on the odds ratio of the uncontroversial product versions.
-5.2.2 Asymmetric Type Distribution
-For a symmetric population, the subjective component cancels out in the value of the recommendation system, allowing for a focus on the role of the objective content.
-Moreover, symmetry simplifies the analysis by ensuring that all types accept the recommendation.
-However, the assumption of a symmetric population may prove restrictive in some contexts, and it does not allow us to study the role of the subjective component for the optimal design of the recommendation system.
-To make headway in that direction, we introduce the following assumption.
+Intuitively, increasing $R$ makes the system more selective: fewer buy recommendations are issued, but those that are issued are more informative—i.e., they imply a higher expected payoff ($\Delta_{O}^{B}$).
+Conversely, lowering $R$ makes the system less selective, increasing the likelihood of buy recommendations at the cost of informativeness.
+When good products are more likely than bad ones ($\sigma>1$), it is optimal to be selective with buy recommendations, using a higher threshold to ensure that buy recommendations are more trustworthy.
+When bad products are more likely ($\sigma<1$), the opposite is true: being selective with don’t-buy recommendations is more valuable, and the platform should lower $R$.
+These forces exactly balance out when $\sigma=1$, so that the value of the system becomes independent of the threshold.
+Put differently, the designer chooses whether the recommendation system should focus on identifying good or bad products.
+The threshold is set to reveal the more likely of these with greater certainty, aligning the selectiveness of the system with the underlying distribution of product quality.
+4.2 Demand-Maximizing Recommendation Systems
+We now turn to the analysis of demand-maximizing recommendation systems. Demand maximization requires a lenient recommendation threshold:
+Proposition 6.
+Suppose Assumption 1 holds. Then, demand is decreasing in $R$.
+By Proposition 4 all types accept recommendations at all thresholds $R$ under Assumption 1.
+Thus, we can ignore the components of demand in equation (8) that come from receivers who do not accept don’t-buy recommendations.
+As a result, demand simplifies to $D(R)=\pi_{B}(R)$, and the result follows because the probability of buy recommendations is decreasing in $R$ [^16] See Lemma A.2 in the Appendix.
+This result contrasts with Proposition 5 on value maximization.
+When $\sigma<1$, both value and demand are decreasing in $R$, so the objectives align.
+But when $\sigma>1$, they are diametrically opposed: while value increases with $R$, demand continues to fall.
+The underlying tradeoff is that increasing $R$ makes the system more selective—it raises the bar for a buy recommendation.
+From a value-maximization perspective, this can be desirable: fewer but more informative recommendations help receivers avoid bad matches, increasing their expected payoff.
+In this sense, selectivity enhances informativeness and thereby the value of the system.
+By contrast, from a demand perspective, greater selectivity is never beneficial.
+The goal is simply to maximize the number of purchases, which is achieved by issuing buy recommendations as frequently as possible.
+A lenient system with a low threshold maximizes demand.
+Crucially, the symmetry assumption ensures that this leniency does not backfire: all types interpret recommendations in the same way and accept them.
+Thus, a low threshold increases demand without causing any mismatch between recommendation content and receiver preferences.
+4.3 Preference Heterogeneity and Controversial Products
+As discussed at the end of Section 3.1, preference heterogeneity only matters when controversial products exist.
+This naturally raises the question: does increasing heterogeneity—or increasing the share of controversial products—make recommendation systems less valuable?
+One might expect so, but the answer turns out to be more nuanced.
+To streamline the exposition, we focus only on value rather than demand in the two following subsections.
+4.3.1 Preference Heterogeneity
+To study the effect of changes in the preference heterogeneity, we consider mean-preserving spreads of the type distribution $F$, which preserve symmetry but increase the share of extreme types.
+It is useful to define the quantity $\beta:=F(1/2-R)$, which captures the fraction of the population willing to recommend a controversial product—either $(1,0)$ or $(0,1)$—at threshold $R$.
+For a fixed type distribution $F$, $\beta$ and $R$ are inversely related.
+Hence, we can interpret changes in $\beta$ as resulting either from changes in $R$ for a fixed distribution or from changes in the distribution itself at a fixed threshold.
+In what follows, we treat $\beta$ as a measure of polarization: for $R>1/2$, an increase in $\beta$ reflects a shift toward more extreme preferences, whereas for $R<1/2$, polarization corresponds to a decrease in $\beta$.
+Proposition A.1 in Appendix A.2 shows formally how the value of the recommendation system depends on the threshold $R$ and the odds ratio $\sigma$.
+In the following, we sketch the main insights.
+Recall that $\beta$ represents the share of senders who are willing to recommend a controversial product.
+When $R>1/2$, a buy recommendation for a controversial product requires relatively strong preferences.
+[^17] An analogous argument applies to $R<1/2$.
+A higher $\beta$ then corresponds to an increase in polarization, making buy recommendations more frequent but less informative.
+Whether it is desirable to resolve this trade-off by increasing $\beta$ depends on $\sigma$: If good products are more likely than bad ones ($\sigma>1$), the system gains value from having buy recommendations that are highly informative, even if they are infrequent.
+In that case, polarization is harmful because it makes buy recommendations more common but more noisy.
+Conversely, when bad products are more likely ($\sigma<1$), the value of the system comes from being able to rule out bad products through lenient buy recommendations, and polarization helps by increasing their frequency.
+4.3.2 Prevalence of Controversial Products
+Without controversial products, preference heterogeneity would not matter: all receivers would accept recommendations, as these would reveal whether the product is good or bad.
+Beyond this extreme case, a natural conjecture is that recommendations become less valuable as controversial products become more common.
+We find that this conjecture does not hold without qualification.
+When good and bad products are similarly likely, a higher prevalence of controversial products indeed reduces the value of the recommendation system.
+But when one type of product is much more likely than the other, the value of the system may actually be maximized at an interior share of controversial products.
+A formal statement of this result is given in Corollary 4 in Appendix A.2.
+The reason lies in the structure of the value expression, which is the product of the buying probability and the objective informativeness of a buy recommendation (see Proposition 4).
+As controversial products become more prevalent, good products become less likely, and fewer buy recommendations are triggered by objectively good products.
+At the same time, more buy recommendations stem from controversial products.
+Thus, both the frequency and the informativeness of buy recommendations change.
+Whether the overall value increases or decreases depends on which of these effects dominates and on the odds-ratio $\sigma$.
+When good products are relatively rare ($\sigma<1$), identifying and avoiding bad products becomes especially valuable, which can raise the value of the system despite a lower share of objectively good recommendations.
+Conversely, when bad products are rare ($\sigma>1$), the value can increase as the recommendation system becomes more effective at identifying good ones, even if some recommendations are noisy.
+5 Asymmetric Populations
+While providing a useful benchmark, the assumption of a symmetric population does not allow us to study the role of the subjective component for the optimal design of the recommendation system.
+To make headway in that direction, we study asymmetric populations in this section.
+For the sake of tractability, we make the following assumption.
 Assumption 2.
@@
 We thus capture asymmetry by means of a simple but flexible parameterization.
-By assuming that the controversial products are equally likely a priori, we ensure that any effect of a recommendation exclusively reflects the population asymmetry.
-Corollary 1 implies the following result:
-Corollary 2.
-Suppose Assumption 2 holds. Then, we have:
-[…]
-Thus, whether and how a recommendation shifts the probabilities of the controversial product versions depends entirely on the type distribution.
-When it is symmetric ($a=1$), there is no shift.
-When it is convex ($a>1$), so that more people prefer version $(1,0)$ to $(0,1)$, a buy recommendation shifts probability from $(0,1)$ to $(1,0)$, as it is more likely to come from a person who prefers $(1,0)$ to $(0,1)$.
-Similarly, when the type distribution is concave ($a<1$), the probability is shifted to $(0,1)$.
+By assuming that the controversial products are equally likely a priori, we ensure that the subjective effect of a buy recommendation depends entirely on the asymmetry parameter $a$.
+When the type distribution is symmetric ($a=1$), there is no systematic bias in how controversial products are evaluated.
+As a result, buy recommendations do not shift beliefs about which controversial version the product might be—there is no subjective effect.
+In contrast, when the population is asymmetric (i.e., $a\neq 1$), one controversial version is systematically preferred over the other.
+In that case, buy recommendations tend to come from people who favor the majority-preferred version, and this introduces a systematic bias: The recommendation carries subjective information, shifting beliefs towards the more popular version.
+5.1 Value-Maximizing Recommendation Systems
 We can now analyze the extent to which the result in Proposition 5 that, generically, extreme thresholds are desirable is an artifact of the symmetry assumption.
-This is challenging because, contrary to the symmetric case, the set of types that accept all recommendations depends on the threshold $R$.
-Reflecting this complication, there is no direct counterpart to Proposition 5.
-Nevertheless, the following two results provide some insights into the optimal recommendation system in the case of an asymmetrically distributed population.
+This is challenging because, contrary to the symmetric case, the set of types that accept recommendations depends on the threshold $R$.
+The following two results provide some insights into the optimal recommendation system in the case of an asymmetrically distributed population.
 Proposition 7.
@@
 […]
-Figure 2: Illustration of Propositions 6 and 7
-As illustrated in Figure 2, Proposition 6 provides sufficient conditions for interior solutions.
+Figure 2: Illustration of Propositions 7 and 8
+[…]
+Proposition 7, illustrated in Figure 2, provides sufficient conditions for interior solutions.
 It shows that an interior optimum exists when (i) controversial products are likely ($Q$ is large) and the population is very asymmetric ($a$ is far away from $1$) or otherwise when (ii) the uncontroversial product versions are similarly likely ($\sigma$ is close to $1$).
@@
 In contrast, the average subjective payoff effect does not cancel out in the asymmetric case.
-Thus, the choice of the optimal threshold no longer only depends on $\sigma$ but also on the asymmetry parameter $a$ and the prevalence of controversial products $Q$.
-We explore under which circumstances the conditions given in Proposition 6 are necessary as well as sufficient:
+Thus, the optimal threshold no longer only depends on $\sigma$ but also on the asymmetry parameter $a$ and the prevalence of controversial products $Q$.
+We explore under which circumstances the conditions given in Proposition 7 are not only sufficient but also necessary.
 Proposition 8.
-Suppose Assumption 2 holds and fix $\sigma$. Then, as $Q\rightarrow 0$, $V(R)$ is
+Suppose Assumption 2 holds. If $Q$ is small enough, then
   1. (i)
-    decreasing in $R$ if $\sigma\leq\min\left\{\frac{1-Q(a+1)}{a-Q(a+1)},\frac{a-Q(a+1)}{1-Q(a+1)}\right\}$;
+    $V(R)$ is decreasing in $R$ if $\sigma\leq\min\left\{\frac{1-Q(a+1)}{a-Q(a+1)},\frac{a-Q(a+1)}{1-Q(a+1)}\right\}$;
   2. (ii)
-    increasing in $R$ if $\sigma\geq\max\left\{\frac{1-Q(a+1)}{a-Q(a+1)},\frac{a-Q(a+1)}{1-Q(a+1)}\right\}$;
+    $V(R)$ is increasing in $R$ if $\sigma\geq\max\left\{\frac{1-Q(a+1)}{a-Q(a+1)},\frac{a-Q(a+1)}{1-Q(a+1)}\right\}$;
   3. (iii)
-    maximized for some $R^{\ast}\in(0,1)$ otherwise.
-The proof essentially proceeds in two steps.
-First, all types accept all recommendations when $Q$ is sufficiently small, since then the subjective content in the recommendation vanishes.262626 $Q\rightarrow 0$ implies that $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$, ensuring that all types accept all recommendations by Corollary 1.
-The key step is that $\Delta_{S}^{B}\rightarrow 0$ as $Q\rightarrow 0$, while $\Delta_{O}^{B}\rightarrow 1-q_{H}>0$ when $Q\rightarrow 0$.
-The proof shows that, even for larger values of $Q$, $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$ often holds, explaining why the conclusion of Proposition 7 holds not only in the limit.
+    $V(R)$ is maximized for some $R^{\ast}\in(0,1)$ otherwise.
+The proof proceeds in two steps.
+First, all types accept recommendations when $Q$ is sufficiently small, since then the subjective content in the recommendation is dominated by the objective content.
 This simplifies $V(R)$ by eliminating the case distinctions.
-Leveraging this, we can prove the quasiconcavity of the function, ensuring that the sufficient conditions in Proposition 6 are necessary, too.
-The requirement that $Q\rightarrow 0$ is much stronger than necessary for this result.
-Even for higher values of $Q$, Figure 2 quite accurately separates the parameter range under which an interior solution exists from the area where it does not.
-[^27] The figure is drawn for $Q=0.1$, in which case the if-and-only-if statement is correct for $\sigma\in[1/60,60]$ and $a\in[1/10000,10000]$.
-In a nutshell, the result shows that an interior solution exists if (i) controversial products are sufficiently prevalent, (ii) the population asymmetry is substantial, and (iii) the odds ratio of the uncontroversial products is sufficiently close to 1.
-To understand this intuitively and to contrast it with the symmetric case, suppose that $\sigma>1$.
-Hence, the value of the recommendation system is increasing in $R$ in the symmetric case, as the gain in the expected payoff conditional on receiving a buy recommendation outweighs the loss from the lower probability of a buy recommendation.
-This follows because the subjective effect in $V(R)=\pi^{B}[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i]]$ cancels out with symmetry.
-With an asymmetric population, the term $-\Delta_{S}^{B}\mathbb{E}[i]$ does not vanish.
-To illustrate this, consider a very asymmetric preference distribution such as $F(x)=(x+1/2)^{10}$, for which the probability of a type $i\leq 0$ is less than 0.1%.
-Then, $\mathbb{E}[i]>0$ and, by Corollary 2, $\Delta_{S}^{B}<0$ so that $-\Delta_{S}^{B}\mathbb{E}[i]>0$.
-Hence, the subjective effect is positive on average.
-Intuitively, though this population is distributed very asymmetrically around zero, it is essentially homogeneous in the sense that almost all types agree on the ranking of the controversial products: If such a product results in a buy recommendation, it almost surely has quality vector (1,0), as almost all types prefer this to $(0,1)$.
-As such, these products are not that controversial anymore, and even the subjective effect is “fairly objective.” Now, as $R$ increases, the probability of a buy recommendation still falls, and the “purely objective” payoff part $\Delta_{O}^{B}$ increases.
-However, for $R$ high enough, this “fairly objective” payoff part $-\Delta_{S}^{B}\mathbb{E}[i]$ also decreases, as an extreme threshold $R$ impedes the ability to distinguish between a good product and a widely preferred controversial product.
-While consumers are homogeneous regarding which controversial product version they prefer, they are still heterogeneous with respect to the intensity with which they do so.
-Hence, disentangling the good and the preferred controversial product is worthwhile.
-Therefore, an interior optimal threshold results.
-5.3 Preference Heterogeneity and Controversial Products
-As discussed after Proposition 7, preference heterogeneity only matters when controversial products exist.
-Moreover, the less homogeneous the population, the more likely it is that extreme thresholds are desirable so that receivers learn solely about uncontroversial products.
-Hence, one might conjecture that greater preference heterogeneity and a higher prevalence of controversial products reduce the value of recommendation systems.
-In the following, we address both issues in turn, demonstrating that these conjectures do not hold without qualification.
-For tractability, we focus on symmetric populations.
-5.3.1 Preference Heterogeneity
-In this section, we analyze how changes in the (exogenously given) preference distribution affect the value of the recommendation system.
-[^28] When we consider distinct sender and receiver distributions in Section 6.1, we will build further on this and take a design perspective instead of the comparative statics approach here.
-Under Assumption 1, we can analyze increases in the polarization of the distribution of preferences by keeping the threshold $R$ fixed and varying $\beta$ [^29] For $R=1/2$, we must have $\beta=1/2$ because the symmetry of $F$ requires $F(0)=1/2$, so this exercise is only meaningful for thresholds $R\neq 1/2$.
-Further, for $R>1/2$ we must have $\beta\leq 1/2$ because $\beta=F(1/2-R)\leq F(0)=1/2$.
-Accordingly, for $R<1/2$ we must have $\beta\geq 1/2$.
-This corresponds to a mean-preserving spread of $F$ (see Figure 3), coinciding with an increase in $\beta$ for $R>1/2$ and a decrease for $R<1/2$.
-We obtain the following result.
+We can use this information to prove the quasiconcavity of the function, ensuring that the sufficient conditions in Proposition 7 are necessary, too.
+The result clarifies when interior thresholds emerge despite the recommendation system being inherently coarse.
+In the symmetric case, where subjective preferences cancel out on average, value maximization requires choosing an extreme threshold—either very lenient or very selective—depending on which uncontroversial product is more likely.
+The goal is to make recommendations essentially objective.
+However, in an asymmetric population, recommendations systematically shift beliefs about controversial products in favor of the majority-preferred version.
+This subjective component can be informative, especially when controversial products are common and the population is largely aligned in which one they prefer.
+In that case, even a buy recommendation for a controversial product is not very controversial—it likely signals the version that most people want.
+As a result, it can be optimal to back off from extreme thresholds and instead choose an intermediate threshold that allows drawing inferences about the controversial products, too.
+When preferences are less aligned or controversial products are rare, this logic breaks down, and more extreme thresholds become optimal again.
+We next discuss how the interior threshold $R^{\ast}$—when it exists—responds to changes in the key parameters of the model.
+[^19] Due to the intractability of the problem, all of these statements are based on numerical simulations, see Appendix B.
+First, $R^{\ast}$ is increasing in $\sigma$: as good products become more likely relative to bad ones, the recommendation system should be more selective, and a higher threshold becomes optimal.
+Second, the further the preference distribution departs from symmetry (i.e., the further $a$ is from 1), the more $R^{\ast}$ moves to the interior.
+Intuitively, as $a$ moves further away from 1, the population becomes more homogeneous in its ranking of controversial products, so that including subjective content into the recommendation is informative and not too controversial.
+Third, the effect of the prevalence of controversial products $Q$ is more subtle.
+As $Q$ decreases, the setting effectively approaches that of a symmetric population, since asymmetry becomes irrelevant in the absence of controversial products.
+Accordingly, $R^{\ast}$ moves closer to the extreme threshold that would be optimal in the symmetric case: it approaches 1 if $\sigma>1$ and 0 if $\sigma<1$.
+5.2 Demand-Maximizing Recommendation Systems
+Like value maximization, demand maximization becomes significantly more complex when the population is asymmetrically distributed.
+Nonetheless, we establish a sufficient condition under which a low recommendation threshold is optimal, mirroring the result in the symmetric case.
 Proposition 9.
-Suppose Assumption 1 holds.
-  1. 1.
-    Suppose $\sigma<1$. Then, a mean-preserving spread of the type distribution increases the value of the recommendation system for $R>1/2$ and decreases it for $R<1/2$.
-  2. 2.
-    Suppose $\sigma>1$. Then, a mean-preserving spread of the type distribution decreases the value of the recommendation system for $R>1/2$ and increases it for $R<1/2$.
-To gain some intuition, suppose $\sigma<1$ and $R>1/2$.
-Figure 1 shows that with $R>1/2$, only extreme types give buy recommendations for controversial products.
-[^30] Conversely, for $R<1/2$ only relatively extreme types give don’t-buy recommendations for controversial products.
-This can be seen by flipping the colors in Figure 1 so that it corresponds to $R=1/4$.
-Increasing $\beta$ corresponds to a mean-preserving spread of $F$ given $R$.
-Consequently, there are more such extreme senders and, hence, buy recommendations are increasingly likely to result from controversial products.
-This increases the probability of buy recommendations while lowering their objective value.
-The resolution of this trade-off, and thus whether the value of the recommendation system goes up or down, is determined by the odds ratio $\sigma$ (see Proposition 5).
-For instance, if $\sigma<1$, good products are rare compared to bad ones, so the value of the recommendation system is increasing in $\beta$.
-Therefore, the positive effect of revealing bad products by means of don’t-buy recommendations outweighs the negative effect of a reduced value conditional on receiving a buy recommendation.
-Thus, given a high recommendation threshold ($R>1/2$), a more polarized population benefits more from the recommendation system, as this means only bad products generate don’t-buy recommendations.
-Conversely, when the recommendation threshold is relatively low ($R<1/2$), a less polarized population centered around the middle gives don’t-buy recommendations only to bad products.
-Figure 3: Population polarization
-[…]
-5.3.2 Prevalence of Controversial Products
-Without controversial products, preference heterogeneity would not matter: All receivers would accept all recommendations, as they would reveal whether the product is good or bad.
-Hence, the value of the recommendation system would be independent of the preference distribution.
-More broadly, we now ask how an increasing prevalence of controversial products (an increase in $Q$) affects the value of a recommendation system.
-A natural conjecture is that recommendations become less valuable when controversial products are more likely.
-We find that this conjecture needs to be qualified.
-Corollary 3.
-Suppose Assumption 1 holds.
-  • (i)
-    If $03\sigma-\beta-\sigma^{2}+\sigma^{2}\beta>$ 0, then the value of the recommendation system is decreasing in $Q$. In particular, this is the case when $\sigma=1$.
-  • (ii)
-    If $03\sigma-\beta-\sigma^{2}+\sigma^{2}\beta<$ 0, then $Q^{*}:=\frac{3\sigma-\beta-\sigma^{2}+\sigma^{2}\beta}{4\sigma-4\beta-4\sigma^{2}+4\sigma^{2}\beta}\in(0,\frac{1}{2})$ maximizes $V(R)$.
-Figure 4: Illustration of Corollary 3
-(a) Panel A
-$\sigma$$\beta$111133335555777799990.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81.01.01.01.0
-(b) Panel B
-$\sigma$$\beta$111133335555777799990.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81.01.01.01.0
-(c) Panel C
-$\sigma$$\beta$111133335555777799990.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81.01.01.01.0
-Figure 4 illustrates the result.
-In the white area in Panel A, $V(\beta)$ is decreasing in $\beta$; in the shaded area, an interior maximizer exists.
-When good and bad products are similarly likely ($\sigma$ is close to $1$), a higher prevalence of controversial products indeed reduces $V(\beta)$.
-For very high or low values of $\sigma$, however, this need not be true.
-This surprising result reflects the multiplicative structure of $V(\beta)$.
-Under Assumption 1, $V(\beta)$ is the product of the buying probability $\pi^{B}$ and the objective value of the recommendation per buyer, $\Delta_{O}^{B}$.
-The effect of $Q$ on $\pi^{B}$ is negative only if $\sigma$ is sufficiently large (see Panel B).
-Intuitively, as $Q$ increases, good products are less likely, reducing the probability of buy recommendations.
-However, there are now more buy recommendations stemming from controversial products, as these have become more likely.
-$\beta$ and $\sigma$ determine which of these effects dominates.
-When $\beta$ is high, the greater likelihood that controversial products receive buy recommendations increases the total probability of buy recommendations, strengthening the positive effect.
-When $\sigma$ is high, the lower share of good products dominates, strengthening the negative effect.
-Similarly, one can show that the effect of $Q$ on $\Delta_{O}^{B}$ becomes negative only for small $\sigma$ (see Panel C).
-Again, this reflects countervailing forces.
-The negative effect results from the increased probability that a product with a buy recommendation is controversial.
-The positive effect reflects a change in the value of the outside option.
-The values of $\sigma$ and $\beta$ determine which effect dominates.
-[^31] When $\beta$ is small, the negative effect is weak, as only few types give buy recommendations for controversial products.
-When $\sigma$ is small, the outside option becomes more valuable, as bad products become less likely, weakening the positive effect.
-6 Extensions
-Our benchmark model relies on several assumptions.
-First, the preference distribution of senders and receivers is identical.
-Second, there are only two levels of recommendations (“buy” and “don’t-buy”).
-Third, each receiver only has access to a single recommendation.
-We now relax each of these assumptions in turn.
+Suppose $q_{H}\geq 1/2$. Then, demand is maximized for $R\rightarrow 0$.
+While this result holds without Assumption 1 or Assumption 2, it restricts the analysis to the case that high quality goods are quite common.
+This complication arises because, contrary to the symmetric case, raising $R$ does not simply reduce the probability of buy recommendations—it also changes who follows them: As $R$ increases, controversial products are less likely to be recommended, and the composition of receivers who accept recommendations shifts toward those whose preferences are more aligned with the sender.
+To understand the intuition for Proposition 9, recall that demand can come from two sources: types who accept buy recommendations and types who reject don’t-buy recommendations and choose to buy the product anyway.
+This suggests two potential strategies for maximizing demand.
+One option is to set $R$ low, ensuring that many buy recommendations are issued and that all types follow them (see Proposition 2).
+Another option would be to set an intermediate $R$ where not all types accept recommendations, thus relying on many people rejecting don’t-buy recommendations so that demand remains high even when buy recommendations are rarer.
+However, when $q_{H}\geq 1/2$, buy recommendations are always more frequent than don’t-buy recommendations.
+Consequently, the second approach is always dominated by the first approach.
+Thus, the optimal strategy is to maximize the frequency of buy recommendations by setting $R\to 0$.
+This result stands in stark contrast with value maximization under asymmetry.
+According to Proposition 8, choosing as low a threshold as possible is optimal if and only if $\sigma\leq\min\left\{\frac{1-Q(a+1)}{a-Q(a+1)},\frac{a-Q(a+1)}{1-Q(a+1)}\right\}$ holds.
+However, given $q_{H}\geq 1/2$, it can be shown that this condition is never satisfied.
+This means that the objectives of demand and value maximization are never aligned in this setting: while demand is always maximized at $R\to 0$, value maximization typically favors a higher threshold.
+Intuitively, note that good products are relatively frequent when $q_{H}\geq 1/2$, meaning that buy recommendations are already quite likely.
+In this case, the emphasis in value maximization shifts from frequency to informativeness: Since buy recommendations are common, their informativeness matters more than their frequency.
+Increasing $R$ filters out more controversial products, making buy recommendations more selective and thus more valuable.
+In contrast, demand maximization focuses exclusively on making buy recommendations as frequent as possible.
+As the requirement that $q_{H}\geq 1/2$ is quite strong, we carry out a complementary numerical analysis (see Appendix B) for the case where it does not hold.
+For this analysis, we need to invoke Assumption 2.
+The results suggest that demand is monotonically decreasing in $R$, just as in the symmetric case.
+Therefore, as in the case of a symmetric population, demand maximization will still be aligned with value maximization for some parameters but not for others.
+6 Robustness and Further Design Considerations
+We now relax the assumptions that the preference distribution of senders and receivers is identical, that there are only two levels of recommendations (“buy” and “don’t-buy”) and that each receiver only has access to a single recommendation.
 6.1 Distinct Sender and Receiver Distributions
-We modify the benchmark model by assuming that the sender and the receiver are drawn from different distributions $F$ and $G$, respectively.
-For instance, this enables us to take into account that the “early adopters” of a new product may have different preferences than the general population (similar to Acemoglu et al.
+Allowing for distinct sender and receiver distributions is interesting for two reasons.
+First, it captures situations such as when early adopters differ systematically from the broader population (similar to Acemoglu et al.
 (2022)).
-We show that the results of the benchmark case easily generalize to this setting.
-More importantly, the general setting allows us to treat the choice of the sender distribution as a potentially interesting design variable.
-The sender distribution $F$ fully determines the probabilities of buy and don’t-buy recommendations, the posterior probabilities, and thus the objective and subjective effects and the expected payoffs, $V_{A}$ and $V_{N}$.
-Hence, all results on receiver behavior in Section 4 apply verbatim if $F$ is taken to be the sender distribution, as the receiver’s behavior is independent of the receiver distribution.
-The value of the recommendation system, however, depends on the receiver distribution $G$ as well, as it involves integration over this distribution.
-The next result generalizes Proposition 3.323232The notation for the quantities in Proposition 9 reflects their dependence on $F$ and $G$, respectively.
-Proposition 9.
-Consider the model with distinct preference distributions of senders and receivers.
-Suppose the sender distribution $F$ satisfies Assumption 1.
-Then, all types accept all recommendations.
-Further, $V(R)=\pi^{B}_{F}[\Delta_{O,F}^{B}-\Delta_{S,F}^{B}\mathbb{E}_{G}[i]]$ is
-  1. (i)
-    decreasing in $R$ if $\sigma<\frac{Q-(q_{1}-q_{2})\mathbb{E}_{G}[i]}{Q-(q_{2}-q_{1})\mathbb{E}_{G}[i]}$;
-  2. (ii)
-    increasing in $R$ if $\sigma>\frac{Q-(q_{1}-q_{2})\mathbb{E}_{G}[i]}{Q-(q_{2}-q_{1})\mathbb{E}_{G}[i]}$;
-  3. (iii)
-    constant in $R$ if $\sigma=\frac{Q-(q_{1}-q_{2})\mathbb{E}_{G}[i]}{Q-(q_{2}-q_{1})\mathbb{E}_{G}[i]}$.
-Recall that in the benchmark model symmetry of $F$ implies that all receivers accept all recommendations and the subjective effect of the recommendation cancels out.
-This is still true if $F$ is distinct from $G$ as long as $G$ is also symmetric and hence $\mathbb{E}_{G}[i]=0$.
-For more general distributions $G$, all receivers still accept all recommendations for the same reason as above.
-However, as the receiver distribution is not symmetric, the subjective effect does not cancel out anymore.
-Even so, the result qualitatively mirrors Proposition 5, as changing the threshold still leads to the same trade-off: Reducing the threshold increases the probability of a buy recommendation but decreases its value conditional on that recommendation.
-However, in the latter effect, the value is no longer purely objective but consists of the objective part $\Delta_{O}^{B}$ and a subjective part given by $-\Delta_{S}^{B}\mathbb{E}_{G}[i]$, reflecting the receiver distribution $G$.
-Further, as the subjective effect now influences the value of the system, so does the prevalence of controversial products by affecting the tipping point at which the designer opts for a high or a low threshold.
-In any case, when $\mathbb{E}_{G}[i]=0$, we obtain the result from the baseline model as a special case.
-Similarly, the other results in the benchmark model remain (qualitatively) valid.
-With asymmetrically distributed populations, Propositions 6 and 7 hold for arbitrary receiver distributions as long as we impose Assumption 2 on the sender distribution.
-Allowing for differences in sender and receiver distributions makes it possible to address another important design issue.
-In some cases, it is natural to assume that a designer, for instance a platform, can influence the sender distribution or at least the receiver’s perception of it.
-Assuming that the platform has better knowledge about sender types, it may choose to share some of this information with the receivers or even restrict the option of sending recommendations to certain types of senders (and inform the receivers accordingly).
-For instance, many platforms such as Amazon inform browsing customers that “customers who bought this item also bought” another item.
-There are different ways to model this.
-First, suppose that $F$ is symmetric and that the designer can render the sender distribution more or less precise using mean-preserving contractions or mean-preserving spreads, respectively.
-Intuitively, one would expect more precise knowledge about the sender to increase the value of the recommendation system.
-However, the logic of Proposition 9 shows that this is not necessarily the case.
-As in Section 5.3.1, the effects of mean-preserving contractions and spreads can be captured as changes in $\beta=F(1/2-R)$, just like changes in $R$.
-With mild adaptions, the proof of Proposition 9 can thus be used to show that whether more or less precise information about the sender increases the value of recommendations depends on the parameters.
-Second, suppose that the sender and receiver distributions are fixed but that the designer can provide the receivers with a signal about the sender that allows for inference about the sender’s type.
-To make the point as starkly as possible, suppose the signal fully reveals the sender’s type.
-If the sender is revealed to be type $i=0$, all types accept all recommendations, as both types of recommendations contain only objective information.
-In contrast to Proposition 2—where this behavior is shown to ensue with extreme thresholds—this would arise for all thresholds $R\in(0,1)$.
-Intuitively, as the type $i=0$ is indifferent between controversial products, her recommendation never reveals subjective information.
+Second, and more importantly, it may be possible to shape the sender population or influence how receivers perceive it.
+We first show that our earlier results extend naturally to the case of distinct sender and receiver distributions.
+Let the sender be drawn from $F$ and the receiver from $G$.
+If $F$ satisfies Assumption 1 or 2, the results on value- and demand-maximizing thresholds from Sections 4 and 5 remain valid.
+For example, Proposition A.2 in the appendix generalizes Proposition 5, showing that all receiver types accept recommendations under symmetry of $F$, and that $V(R)$ remains monotonic in $R$ with the slope depending on $\sigma$.
+Similarly, the potential conflict between value and demand objectives persists.
+Beyond such robustness results, differing sender and receiver distributions give rise to new design considerations.
+In many cases, platforms may have partial control over the sender population or the information receivers have about it.
+For instance, Amazon’s phrasing “customers who bought this item also bought…” implicitly informs the receiver about who the senders are.
+We outline different ways to model this.
+First, suppose the designer can manipulate the informativeness of a symmetric sender distribution $F$ through mean-preserving spreads or contractions.
+Though it may seem that making the sender distribution more precise would raise the value of the system, this is not always true.
+As in Section 4.3.1, changes in precision translate into changes in $\beta=F(1/2-R)$.
+With mild adaptations, the proof of Proposition A.2 can be used to show how the effect on value depends on the overall environment.
+Second, suppose that the sender and receiver distributions are fixed, but that the designer can provide the receivers with a signal about the sender’s type.
+To make the point as starkly as possible, suppose that the signal fully reveals the type.
+If the sender is revealed to be $i=0$, all types accept recommendations, as the latter contain only objective information.
+In contrast to Proposition 2, where this behavior requires extreme thresholds, this would arise for all thresholds $R\in(0,1)$.
+Intuitively, as $i=0$ is indifferent between controversial products, her recommendation does not reveal subjective information.
 In contrast, if the signal reveals a sender of type $i\neq 0$, the receiver will be able to learn about controversial products, too.
 [^20] For instance, if $i=1$ and $R\in(0,1)$, a buy recommendation reveals that the product is either of version $(1,1)$ or $(1,0)$ Thus, having a sender who is partial to some controversial product is not necessarily a bad thing.
-Finally, suppose that the sender and receiver distributions are not independent but somehow correlated.
-Recommendations would then be “tailored” to the receiver $\bar{i}$, as they would be drawn from a sender distribution $F(i|\bar{i})$, resulting in type-specific posteriors $p^{r}_{j}(i)$ for $j\in\{H,1,2,L\}$.
-In such a setting, an analogous result to Proposition 1 still holds, establishing that either all all or no recommendations are accepted.
-Further analysis would give rise to similar complications as in Section 5.2.2 with asymmetric distributions, as the the correlation between sender and receiver distribution would affect the set of types who accept all recommendations.
+Third, suppose sender and receiver types are correlated so that recommendations are “tailored” to the receiver.
+In this case, the sender distribution becomes type-dependent, $F(i|\bar{i})$, and induces posterior beliefs that differ across receivers.
+While a suitably modified version of Proposition 1 still ensures that either all or no types follow recommendations, the set of types who do so now depends on the joint distribution.
+This leads to complications in further analyses similar to those discussed in Section 5 for the case of asymmetric populations.
 6.2 Multiple Recommendation Thresholds
-We now analyze the effects of reducing the coarseness of the recommendation system.
-We maintain all assumptions of the benchmark case, except that there are two threshold levels $R_{1},R_{2}\in(0,1)$ with $R_{2}\geq R_{1}$ and three types of recommendations $r\in\{D,N,B\}$ which we interpret as “don’t-buy,” “neither buy nor don’t-buy,” and “buy.” [^34] Such a two-threshold structure is found to be optimal in models of strategic communication in the context of social learning by Smirnov and Starkov (2024) and Bénabou and Vellodi (2024).
-The sender gives a buy recommendation if the payoff is at least $R_{2}$, a don’t-buy recommendation if it is below $R_{1}$, and neither if it is in $[R_{1},R_{2})$.
-We maintain Assumption 1.
-The proof of Proposition 10 in Appendix A.3 provides the receiver’s posteriors.
-For buy and don’t-buy recommendations, these posteriors are analogous to the benchmark, except that the different thresholds in each case have to be taken into account.
-As in the benchmark model, the highest and lowest recommendations rule out the bad and good product versions, respectively.
-The intermediate recommendation $r=N$, however, rules out both of these versions, so the receiver knows that the recommended product is controversial.
-We say that a receiver accepts all recommendations if he buys if $r=B$ and does not buy if $r=D$, allowing for either behavior if $r=N$.
-Assumption 1 again implies that the receiver always buys a product with $r=B$ and never buys if $r=D$.
-As noted above, $r=N$ rules out both uncontroversial products.
-Given the symmetric population, $r=N$ leaves the relative likelihood of the controversial products unchanged.
-Hence, whether a type $i$ wants to buy a product with $r=N$ depends on (i) how likely a good product is and (ii) how much he (dis-)likes the more likely controversial product.
-As in the benchmark model, we can derive the indifferent type so that, depending on parameters, either all, only sufficiently aligned, or no types buy the product with an intermediate recommendation.
-$V(R)$ now consists of two components.
-First, as in the benchmark, all types accept $r=B$, resulting in the component $\pi^{B}\Delta_{O}^{B}$.
-[^35] As before, accepting don’t-buy recommendations does not create value.
-Moreover, $r=N$ can create value for the types that buy following that neutral recommendation.
-As noted above, parameter values determine which types do this.
-Despite this ambiguity, we can formulate the following result.
+We now consider a recommendation system with finer granularity.
+Specifically, we relax the benchmark assumption of binary recommendations by introducing a third, intermediate option.
+Such richer recommendation structures appear in real-world platforms and are also shown to be optimal in models of strategic communication under social learning (e.g., Smirnov and Starkov, 2024; Bénabou and Vellodi, 2024).
+Formally, we introduce two threshold levels $R_{1},R_{2}\in(0,1)$ with $R_{2}\geq R_{1}$, and allow for three recommendation types $r\in\{D,N,B\}$: “don’t-buy,” “neutral,” and “buy.” The sender issues a buy recommendation if her payoff is at least $R_{2}$, a don’t-buy if it is below $R_{1}$, and a neutral recommendation otherwise.
+We maintain Assumption 1 throughout.
+As in the benchmark model, the highest and lowest recommendations convey objective information: $r=B$ rules out bad products, and $r=D$ rules out good ones.
+The neutral recommendation, by contrast, carries no objective content; it implies that the product is controversial, but does not distinguish between versions $(1,0)$ and $(0,1)$.
+This leads to the following result:
 Proposition 10.
-Consider the model with multiple recommendation thresholds. Suppose Assumption 1 holds. $V(R)$ is (i) decreasing in $R_{1}$ if $\sigma<1$ and (ii) increasing in $R_{2}$ if $\sigma>1$.
-The result provides an incomplete characterization of the optimal recommendation system: It establishes that one of the two thresholds should be extreme so that one recommendation essentially becomes objective.
-Thus, as in Proposition 5, the designer maximizes the probability of an objective recommendation, considering which uncontroversial product version is more likely a priori.
-The optimal value of the threshold that is not pinned down by Proposition 10 depends on the model parameters unless $q_{1}=q_{2}$.
-In this degenerate case, any value for the other threshold is optimal, but it is also possible that both the high and the low recommendations are essentially turned objective.
-[^36] This arises for instance for $F(i)=1/2+i$ with parameters $q_{1}=0.7,\ q_{2}=0.1,$ and $q_{H}=0.03.$
+Suppose there are recommendation thresholds $R_{2}\geq R_{1}$ and Assumption 1 holds. Then $V(\mathcal{R})$ is (i) decreasing in $R_{1}$ if $\sigma<1$, and (ii) increasing in $R_{2}$ if $\sigma>1$.
+This result mirrors the logic of Proposition 5: the designer maximizes the value of the recommendation system by ensuring that one of the thresholds produces recommendations that are essentially objective.
+The optimal threshold depends on which uncontroversial product version is more likely.
+If, for instance, bad products are more prevalent ($\sigma<1$), the low threshold $R_{1}$ should be set close to zero to ensure that don’t-buy recommendations are informative.
+The optimal value of the remaining threshold depends on the model parameters, unless $q_{1}=q_{2}$, in which case any choice yields the same outcome.
+In some cases, both thresholds may be set at extreme values, rendering both $r=B$ and $r=D$ essentially objective.
+[^21] This arises, for instance, under $F(i)=1/2+i$ with $q_{1}=0.7$, $q_{2}=0.1$, and $q_{H}=0.03$.
+For a demand-maximizing platform, however, richer recommendation systems offer no benefit under Assumption 1.
+Since receivers of the neutral recommendation are only partially informed, they may or may not buy the product, depending on the environment.
+As shown in the proof of Proposition 10, this ambiguity leads to weakly lower demand than under a binary system with a very low threshold $R\to 0$, which maximizes the number of buy recommendations.
+Hence, the intermediate recommendation dampens demand without offsetting gains in platform revenue.
 6.3 Multiple Recommendations
 We now show that the value of a recommendation system does not always increase if receivers obtain more than one recommendation: The additional recommendations do not add value if the ex ante probabilities of the good and bad products are similar while those of the two controversial product versions are not.
-We assume that receivers have access to a number $b$ of buy recommendations and a number $d$ of don’t-buy recommendations.
-As in the benchmark model, recommendations come from randomly drawn senders.
-We add the assumption that they are drawn independently.
-[^37] In Appendix A.4, we provide the posteriors, depending on $b$ and $d$.
-Just as before, a single buy recommendation is enough to rule out a bad product, and a single don’t-buy recommendation is enough to rule out a good product.
-Consequently, when the receiver obtains mixed recommendations, that is, both buy and don’t-buy recommendations, he rules out both uncontroversial product versions.
-[^38] In Sun (2011), a higher variance in the recommendations corresponds to a niche product that some consumers like and some don’t.
-Similarly, in our paper, mixed recommendations (and thus a higher variance) result from controversial products that, indeed, some consumers like more than others.
-To make our main point that multiple recommendations need not increase the value of the recommendation system as starkly as possible, we focus on the case of “infinite learning” where the receiver has access to an unlimited number of recommendations.
+Formally, suppose the receiver observes $b$ buy and $d$ don’t-buy recommendations, each independently drawn from a population of senders as in the benchmark model.
+[^22] We derive the posteriors for general $(b,d)$ in Appendix A.5.
+A single buy recommendation is sufficient to rule out a bad product, and a single don’t-buy recommendation rules out a good product.
+Hence, when the receiver obtains a mix of both, he infers that the product is not uncontroversial—it must be one of the two controversial versions.
+[^23] See also Sun (2011), who relate recommendation variance to product “nicheness” in a similar spirit.
+To highlight the limits of informational gains from additional recommendations, we focus on the case of infinite learning, where the receiver observes an unbounded number of independent recommendations.
 Proposition 11.
 Consider the model with multiple recommendations.
-Suppose Assumption 1 holds.
-Let $\lambda:=q_{1}/q_{2}$.
-Compared with the optimal recommendation system in the benchmark model, infinite learning does not increase the value of the recommendation system if and only if either (i), (ii), or (iii) holds:
+Suppose Assumption 1 holds and let $\lambda:=q_{1}/q_{2}$.
+Then, infinite learning does not increase the value of the recommendation system relative to the optimally designed single-recommendation benchmark if and only if one of the following holds:
 […]
-To grasp the result intuitively, recall the receiver’s behavior in the single-recommendation system.
-With a threshold of $R\rightarrow 0$, the receiver buys the recommended product unless it is bad.
-For $R\rightarrow 1$, he does so only if it is good.
-In the benchmark model, V(R) is decreasing in $R$ if $\sigma<1$ and increasing if $\sigma>1$.
-Lemma 4 in Appendix A.4 describes optimal receiver behavior with infinite learning.
-It shows that behavior coincides with the optimally designed single-recommendation system if the product is uncontroversial: Good products are bought but bad ones are not.
-Thus, the infinite-learning system potentially differs from the benchmark only for controversial products, which generate mixed recommendations.
-When $\sigma$ is very low, bad products are relatively likely, so the outside option looks relatively bleak.
-A consumer facing a product with mixed recommendations will thus buy this (controversial) product.
-In contrast, when $\sigma$ is very high, any such consumer will buy the outside option, as it is likely to be good.
-Thus, when the probabilities of the uncontroversial product versions are not too similar, the behavior with a single recommendation and with multiple recommendations coincides, and multiple recommendations do not increase the value of the recommendation system.
-However, multiple recommendations do affect behavior if the probabilities of the good and bad products are similar.
-For instance, when $\lambda>1$, low types who value the less likely controversial product version $(0,1)$ more strongly than product version $(1,0)$ will opt for the outside option rather than the controversial product.
-Then, the system’s value with multiple recommendations is higher than that of the single-recommendation system, as it provides more granular information.
-Inspecting the conditions on $\sigma$ in (ii) and (iii), we find that the more extreme (further away from 1) the odds ratio of the controversial products is, the more extreme the odds ratio of the uncontroversial product versions needs to be for there to be no value in additional recommendations.
+The intuition is as follows.
+With infinitely many recommendations, the receiver can perfectly infer whether the product is good, bad, or controversial.
+As in the benchmark case, receivers buy good products and reject bad ones.
+Thus, any potential gain must come from improved decisions about controversial products.
+However, controversial products always generate mixed recommendations.
+Whether the receiver ultimately chooses to buy depends on his posterior belief and thus on the relative likelihood of the outside option being good or bad.
+If good products are relatively rare (that is, $\sigma$ is low), the outside option is unattractive, and the receiver may prefer a controversial product even if it is only an imperfect match.
+If good products are common (that is, $\sigma$ is high), the receiver is more willing to choose the outside option.
+In either extreme, the decision under infinite learning coincides with what could already be achieved by a single recommendation from an optimally calibrated recommendation system.
+Only when the uncontroversial versions are similarly likely and controversial products are sufficiently ambiguous does additional information add value.
+Finally, note that providing more recommendations also does not help a demand-maximizing platform under Assumption 1.
+If the platform’s goal is to induce as many purchases as possible, it suffices to choose a single threshold $R\rightarrow 0$, which leads all senders to issue buy recommendations.
+Since mixed recommendations can only reduce the likelihood of purchasing controversial products, adding them weakly reduces demand.
 7 Conclusion
-We have set out to study the role of preference heterogeneity in recommendations, asking questions such as: Under what circumstances do people accept a recommendation?
-What determines the value of a recommendation system and when is it maximized?
-Our analysis highlights the importance of disentangling objective information from subjective preferences.
-We have seen that, depending on the preference distribution, the subjective content can reinforce the positive news of a buy recommendation or reverse it so that a buy recommendation is treated as a bad signal.
-Conversely, the subjective content can turn a don’t-buy recommendation into good news for some people.
-The design of a recommendation system, as summarized in the threshold $R$, should thus take into account how it influences the relative objectivity of recommendations.
-We find that extreme recommendation systems may be valuable, as they convey objective information even when preferences are heterogeneous.
-However, when preferences are not distributed symmetrically and controversial products are sufficiently prevalent, with enough consensus on their ranking, the optimal recommendation threshold is set at an intermediate level.
-This allows receivers to not only draw inferences about the uncontroversial products but also about the controversial ones.
-For asymmetrically distributed populations, subjective content in the recommendation is thus valuable when controversial products are common and not too controversial.
-Throughout the paper, we have analyzed the value of the recommendation system from a consumer perspective.
-Arguably, online recommendation systems employed on an e-commerce platform are more likely to be geared towards the interests of the platform.
-Depending on the details of the fee system that it employs, the platform could, for instance, be interested in maximizing the probability of purchases.
-It is possible to look at such questions in the context of our model, too.
-A naïve conjecture would be that it is then optimal to set as low a recommendation threshold as possible so that all products except the bad one yield buy recommendations.
-However, recall that it is not necessarily the case that all receivers accept recommendations.
-The analysis of this problem is then non-trivial.
-[^39] In the case of a symmetric population, in which all recommendations are accepted, the problem is simpler, and a sales-maximizing recommendation system indeed sets as low a threshold as possible.
-Moreover, our results show that this may even be optimal for consumers.
-This latter part stands in contrast to, for instance, the welfare-reducing inflated ratings in Peitz and Sobolev (forthcoming).
+This paper has examined under which circumstances coarse, non-strategic recommendations can meaningfully guide decision-making in environments with heterogeneous preferences and subjective product quality.
+Our analysis highlights that even minimal feedback, such as a binary “likes”, can transmit valuable information.
+At the heart of our approach is the distinction between the objective and subjective content of a recommendation.
+While a buy or don’t-buy signal always reveals something about product quality, its usefulness depends on how much of that signal reflects shared preferences versus individual taste.
+We show that when recommendations are largely objective, all receivers are willing to follow them, either by construction (e.g., through extreme thresholds) or by aggregation (e.g., in symmetric populations).
+Systematic preference heterogeneity complicates this picture.
+In asymmetric populations, where controversial products are common and preferences are aligned in one direction, a recommendation typically contains both objective and subjective content.
+The same signal—say, a buy recommendation—may reflect a product’s broad appeal, or it may simply reveal that it aligns with the preferences of a particular sender.
+In such settings, subjectivity introduces friction: Receivers must infer how much of the recommendation reflects shared values rather than idiosyncratic taste, and in some cases, rejecting a recommendation may be the better choice.
+However, subjectivity need not be treated solely as a source of noise.
+When preference alignment is strong and controversial products are sufficiently prevalent, many controversial products are in fact not very controversial in the eyes of the majority.
+Then, an optimally designed recommendation system enables meaningful inference about such products by preserving some subjective content.
+In contrast, when this condition fails, it becomes optimal to adopt an extreme threshold that filters out all subjective noise.
+This stands in contrast to the symmetric population case, where the subjective content of the recommendation system should be suppressed by focusing entirely on extreme thresholds.
+Throughout, we contrast the objectives of maximizing user value with maximizing demand.
+These goals are sometimes aligned, but they often diverge.
+A platform seeking to drive demand may prefer lenient recommendation thresholds that generate more buy signals, even if this reduces informativeness and harms users.
 Finally, we have maintained a number of classical assumptions such as risk neutrality, linear probability weights, and Bayesian updating.
@@
 A.1 Proofs and Technical Details for the Benchmark Model
-The Effects of Recommendations on Beliefs
-To understand how beliefs are updated after a recommendation, it is useful to decompose the difference between beliefs with and without a recommendation into several components.
-We illustrate this for buy recommendations.
-Let $\textbf{p}(r)=(p_{H}^{r},p_{1}^{r},p_{2}^{r},p_{L}^{r})$ denote the vector of posterior probabilities following a recommendation $r\in\{B,D\}$.
-The total effect of the buy recommendation $B$ on the probability vectors is the sum of three parts,
-$\textbf{p}(B)-\textbf{q}=(\textbf{p}(B)-\textbf{q}^{\prime\prime})+(\textbf{q}^{\prime\prime}-\textbf{q}^{\prime})+(\textbf{q}^{\prime}-\textbf{q})\text{,}$ (9)
-where the three components of the right-hand side will be explained in Steps 1 to 3 below.
-Step 1
-The first effect of a buy recommendation is that the probability of an objectively bad product falls to zero.
-We isolate this effect by assuming that the odds ratios between the remaining events remain unchanged so that the effect corresponds to a change from the prior probability q to $\textbf{q}^{\prime}:=(\frac{q_{H}}{1-q_{L}},\frac{q_{1}}{1-q_{L}},\frac{q_{2}}{1-q_{L}},0)$.
-This effect of the buy recommendation states that all three quality profiles other than $(0,0)$ become more likely at the expense of quality profile $(0,0)$.
-Step 2
-Next, the buy recommendation means that the objectively good product becomes more likely relative to the controversial products.
-We isolate this effect by assuming that (i) the probability of $(1,1)$ increases to $p_{H}$ and (ii) the odds ratios between the remaining events remain unchanged so that the effect corresponds to a change from ${\textbf{q}}^{\prime}:=(\frac{q_{H}}{1-q_{L}},\frac{q_{1}}{1-q_{L}},\frac{q_{2}}{1-q_{L}},0)$ to ${\textbf{q}}^{\prime\prime}:=(p_{H},k\frac{q_{1}}{1-q_{L}},k\frac{q_{2}}{1-q_{L}},0)$, where $k=\frac{(1-p_{H})(1-q_{L})}{q_{1}+q_{2}}$ to guarantee that ${\textbf{q}}^{\prime\prime}$ is a probability vector.
-Clearly, $p_{H}>\frac{q_{H}}{1-q_{L}}$.
-Step 3
-Finally, the buy recommendation may change the odds between the controversial product versions.
-We isolate this effect as a change in the probability vector from ${\textbf{q}}^{\prime\prime}:=(p_{H},k\frac{q_{1}}{1-q_{L}},k\frac{q_{2}}{1-q_{L}},0)$ to $\textbf{p}(B)$.
+Posterior Distributions
+The receiver’s posterior $\mathbf{p}^{B}(R):=(p_{H}^{B}(R),p_{1}^{B}(R),p_{2}^{B}(R),p_{L}^{B}(R))$ after a buy recommendation is
+$p_{H}^{B}(R):=\frac{q_{H}}{q_{H}+q_{1}\phi_{1}(R)+q_{2}\phi_{2}(R)},\quad\quad$ $p_{2}^{B}(R):=\frac{q_{2}\phi_{2}(R)}{q_{H}+q_{1}\phi_{1}(R)+q_{2}\phi_{2}(R)},$
+$p_{1}^{B}(R):=\frac{q_{1}\phi_{1}(R)}{q_{H}+q_{1}\phi_{1}(R)+q_{2}\phi_{2}(R)},\quad\quad$ $p_{L}^{B}(R):=0.$
+A don’t-buy recommendation gives rise to $\mathbf{p}^{D}(R):=(p_{H}^{D}(R),p_{1}^{D}(R),p_{2}^{D}(R),p_{L}^{D}(R))$, where
+$p_{H}^{D}(R):=0,\,$ $p_{2}^{D}(R):=\frac{q_{2}(1-\phi_{2}(R))}{q_{1}(1-\phi_{1}(R))+q_{2}(1-\phi_{2}(R))+q_{L}},$
+$p_{1}^{D}(R):=\frac{q_{1}(1-\phi_{1}(R))}{q_{1}(1-\phi_{1}(R))+q_{2}(1-\phi_{2}(R))+q_{L}},\,$ $p_{L}^{D}(R):=\frac{q_{L}}{q_{1}(1-\phi_{1}(R))+q_{2}(1-\phi_{2}(R)))+q_{L}}.$
 Proof of Proposition 1
-Part (i): Accepting the don’t-buy recommendation is optimal if and only if $U(0)-U(D)\geq 0$ or, equivalently,
-$q_{H}+q_{1}(1/2+i)+q_{2}(1/2-i)\geq p_{1}^{D}(1/2+i)+p_{2}^{D}(1/2-i)\text{.}$ (10)
-Similarly, accepting the buy recommendation is optimal if and only if $U(B)-U(0)\geq 0$ or, equivalently,
-$p_{H}^{B}+p_{1}^{B}(1/2+i)+p_{2}^{B}(1/2-i)\geq q_{H}+q_{1}(1/2+i)+q_{2}(1/2-i)\text{.}$ (11)
-We shall show that conditions (10) and (11) are equivalent, thus proving the claim that receivers either accept all recommendations or never accept recommendations.
-To do so, we prove the following two steps.
-Step 1: (10) is equivalent with
-$1-p_{1}^{B}-p_{2}^{D}\geq(1/2+i)[p_{2}^{B}-p_{2}^{D}+p_{1}^{D}-p_{1}^{B}]\text{.}$ (12)
-Step 2: (11) is equivalent with (12).
-The proofs of both steps will rely on the identities
-$q_{s}=\pi^{B}p_{s}^{B}+\pi^{D}p_{s}^{D}=(1-\pi^{D})p_{s}^{B}+\pi^{D}p_{s}^{D}\text{ for }s\in\{H,1,2\}$ (13)
-and
-$p_{H}^{B}=1-p_{1}^{B}-p_{2}^{B}\text{,}$ (14)
-Proof of Step 1: Replacing $q_{H}$, $q_{1}$ and $q_{2}$ in (10) using (13), we obtain
-$\left(1-\pi^{D}\right)p_{H}^{B}+\left(\left(1-\pi^{D}\right)p_{1}^{B}+\pi^{D}p_{1}^{D}\right)\left(\frac{1}{2}+i\right)$
-$+\left(\left(1-\pi^{D}\right)p_{2}^{B}+\pi^{D}p_{2}^{D}\right)\left(\frac{1}{2}-i\right)\geq p_{1}^{D}\left(\frac{1}{2}+i\right)+p_{2}^{D}\left(\frac{1}{2}-i\right)\text{.}$
-This simplifies to
-$p_{H}^{B}+p_{1}^{B}\left(\frac{1}{2}+i\right)+p_{2}^{B}\left(\frac{1}{2}-i\right)\geq\left(p_{1}^{D}\left(\frac{1}{2}+i\right)+p_{2}^{D}\left(\frac{1}{2}-i\right)\right)\text{.}$
-Using (14) and then rearranging further gives
-$1-p_{1}^{B}-p_{2}^{B}+p_{1}^{B}\left(\frac{1}{2}+i\right)+p_{2}^{B}\left(\frac{1}{2}-i\right)\geq p_{1}^{D}\left(\frac{1}{2}+i\right)+p_{2}^{D}\left(\frac{1}{2}-i\right)\Leftrightarrow$
-$1-p_{1}^{B}-p_{2}^{D}-p_{2}^{B}+p_{1}^{B}\left(\frac{1}{2}+i\right)+p_{2}^{B}\left(\frac{1}{2}-i\right)\geq p_{1}^{D}\left(\frac{1}{2}+i\right)-p_{2}^{D}\left(\frac{1}{2}+i\right)\Leftrightarrow$
-$1-p_{1}^{B}-p_{2}^{D}+p_{1}^{B}\left(\frac{1}{2}+i\right)+p_{2}^{B}\left(-\frac{1}{2}-i\right)\geq\left(p_{1}^{D}-p_{2}^{D}\right)\left(\frac{1}{2}+i\right)\Leftrightarrow$
-$1-p_{1}^{B}-p_{2}^{D}\geq(1/2+i)[p_{2}^{B}-p_{2}^{D}+p_{1}^{D}-p_{1}^{B}]\text{.}$
-Proof of Step 2: Replacing $q_{H}$ and $q_{1}$ and $q_{2}$ in (11) using (13) gives
-$p_{H}^{B}+p_{1}^{B}\left(\frac{1}{2}+i\right)+p_{2}^{B}\left(\frac{1}{2}-i\right)\geq$
-$\left(1-\pi^{D}\right)p_{H}^{B}+\left(\left(1-\pi^{D}\right)p_{1}^{B}+\pi^{D}p_{1}^{D}\right)\left(\frac{1}{2}+i\right)+\left(\left(1-\pi^{D}\right)p_{2}^{B}+\pi^{D}p_{2}^{D}\right)\left(\frac{1}{2}-i\right)\text{.}$
-Rearranging this expression gives
-$p_{H}^{B}+p_{1}^{B}\left(\frac{1}{2}+i\right)+p_{2}^{B}\left(\frac{1}{2}-k\right)-p_{1}^{D}\left(\frac{1}{2}+i\right)-p_{2}^{D}\left(\frac{1}{2}-i\right)\geq 0$
-Adding $-p_{1}^{B}-p_{2}^{D}+p_{2}^{D}+p_{1}^{B}$ on the left-hand side gives:
-$p_{H}^{B}-p_{1}^{B}-p_{2}^{D}+p_{2}^{D}+p_{1}^{B}+p_{1}^{B}\left(\frac{1}{2}+i\right)+p_{2}^{B}\left(\frac{1}{2}-i\right)-p_{1}^{D}\left(\frac{1}{2}+i\right)-p_{2}^{D}\left(\frac{1}{2}-i\right)\geq 0$
-and thus
-$p_{H}^{B}-p_{1}^{B}-p_{2}^{D}\geq-p_{2}^{D}-p_{1}^{B}-p_{1}^{B}\left(\frac{1}{2}+i\right)-p_{2}^{B}\left(\frac{1}{2}-i\right)+p_{1}^{D}\left(\frac{1}{2}+i\right)+p_{2}^{D}\left(\frac{1}{2}-i\right)\text{.}$
-Using (14) gives
-$1-p_{1}^{B}-p_{2}^{B}-p_{1}^{B}-p_{2}^{D}$
-$\geq-p_{2}^{D}-p_{1}^{B}-p_{1}^{B}\left(\frac{1}{2}+i\right)-p_{2}^{B}\left(\frac{1}{2}-i\right)+p_{1}^{D}\left(\frac{1}{2}+i\right)+p_{2}^{D}\left(\frac{1}{2}-i\right)$
-or equivalently
-$1-p_{1}^{B}-p_{2}^{B}\geq-p_{1}^{B}\left(\frac{1}{2}+i\right)-p_{2}^{B}\left(\frac{1}{2}-i\right)+p_{1}^{D}\left(\frac{1}{2}+i\right)+p_{2}^{D}\left(\frac{1}{2}-i\right)$
-Adding $-p_{2}^{D}+p_{2}^{B}$ on both sides gives:
-$1-p_{1}^{B}-p_{2}^{D}-p_{2}^{B}+p_{2}^{B}$
-$\geq-p_{1}^{B}\left(\frac{1}{2}+i\right)-p_{2}^{B}\left(\frac{1}{2}-i\right)+p_{1}^{D}\left(\frac{1}{2}+i\right)+p_{2}^{D}\left(\frac{1}{2}-i\right)-p_{2}^{D}+p_{2}^{B}\text{,}$
-which is equivalent with (12).
+Part (i): Let $\Delta^{B}(i)\equiv U^{B}(i)-U^{0}(i)$ be the gain from following a buy–recommendation and $\Delta^{D}(i)\equiv U^{0}(i)-U^{D}(i)$ the gain from following a don’t-buy–recommendation.
+A receiver follows a recommendation iff the corresponding gain is non–negative.
+For every state $s\in\{H,1,2,L\}$, Bayes’ rule implies
+$\pi^{B}(p^{B}_{s}-q_{s})+\pi^{D}(p^{D}_{s}-q_{s})=0.$
+Multiplying by the state–contingent payoffs $\{1,1/2+i,1/2-i,0\}$ and summing over the states we obtain
+$\pi^{B}\Delta^{B}(i))=\pi^{D}\Delta^{D}(i).$
+Because $\pi^{B},\pi^{D}>0$, the differences $\Delta^{D}(i)>0$ and $\Delta^{D}(i)$ have the same sign, so that the receiver either follows both recommendations or rejects both, proving part (i).
 Part (ii): This follows directly from plugging in $\Delta^{r}_{O}$ and $\Delta^{r}_{S}$ into equation (1) for the buy recommendation the analogous equation for a don’t-buy recommendation.
@@
 Assumption 1 implies $\mathbb{E}[i]=0$, so that the result follows from Proposition 3.
-Proof of Lemma 1
+Proof of Proposition 5
+We first restate the probabilities of buy recommendations, their objective content and the value of the recommendation system in terms of $\sigma$, $\beta$ and $Q$:
+Lemma A.1.
+Suppose Assumption 1 holds. Then,
+$\pi^{B}$ $=\frac{(1-2Q)\sigma}{1+\sigma}+2Q\beta;$
+$\Delta_{O}^{B}$ $=\frac{\beta Q(\sigma+1)-2Q\sigma+\sigma}{2\beta Q(\sigma+1)-2Q\sigma+\sigma}-\frac{(1-2Q)\sigma}{1+\sigma}-Q;$
+$V(\beta)$ $=\pi^{B}\Delta_{O}^{B}=(1-2Q)\frac{\sigma+Q\left(\beta-\sigma+\sigma^{2}(1-\beta)\right)}{(\sigma+1)^{2}}.$
+Proof.
 The result on $\pi^{B}$ follows from inserting $q_{H}=(1-2Q)\sigma/(1+\sigma)$, $\phi_{1}(R)=\phi_{2}(R)=\beta$ and $q_{1}+q_{2}=2Q$ into $\pi^{B}=q_{H}+q_{1}\phi_{1}(R)+q_{2}\phi_{2}(R)$.
@@
 The result on $V(\beta)$ then follows immediately from Proposition 4.
-Proof of Lemma 2
+The following result shows that countervailing effects of $\beta$ on the probability of a buy recommendation and the objective value $\Delta_{O}^{B}$ of accepting it result in a non-trivial effect on $V(\beta)=\pi^{B}\Delta_{O}^{B}$.
+Lemma A.2.
+Suppose Assumption 1 holds. Then,
+  • (i)
+    $\pi^{B}$ is strictly increasing and $\Delta_{O}^{B}$ strictly decreasing in $\beta$;
+  • (ii)
+    $\frac{\partial^{2}V}{\partial\beta\partial\sigma}<0$ and $V^{\prime}(\beta)=\begin{cases}>0&\text{if }\sigma<1;\\ =0&\text{if }\sigma=1;\\ <0&\text{if }\sigma>1.\end{cases}$
+Proof.
 (i) Note that $\pi^{B}=(1-2Q)\sigma/(1+\sigma)+2Q\beta$, so that $\pi_{B}$ is strictly increasing. Moreover,
 $\frac{\partial{\Delta_{O}^{B}}}{\partial{\beta}}=-\frac{(Q(1-2Q)\sigma(1+\sigma))}{(\sigma-2Q\sigma+2\beta Q(1+\sigma))^{2}}<0.$
-(ii) Lemma 1 implies
+(ii) Lemma A.1 implies
 $V^{\prime}(\beta)=\frac{Q(1-2Q)(1-\sigma)}{1+\sigma},$
@@
 $V^{\prime}(\beta)$ is negative for $\sigma>1$, positive for $\sigma<1$ and zero for $\sigma=1$.
-Proof of Corollary 2
-Using (3), applying the $q_{H}$ notation for ease of exposition and imposing Assumption 2 and $q_{1}=q_{2}$ we obtain
-$\Delta_{S}^{B}=-\frac{Q(1-R^{a}-(1-R)^{a})}{q_{H}+Q(1-R^{a}+(1-R)^{a})}.$ (18)
-For $a=1$, this term simplifies to $\Delta_{S}^{B}=0$.
-Hence, it suffices to show that $\Delta_{S}^{B}$ is strictly increasing in $a$.
-To this end, note that the derivative of $\Delta_{S}^{B}$ with respect to $a$ is given by
-$\frac{\partial{\Delta_{S}^{B}}}{\partial{a}}$ $=\frac{Q\left[\log(R)R^{a}\left(2Q(1-R)^{a}+q_{H}\right)+\log(1-R)(1-R)^{a}\left(2Q\left(1-R^{a}\right)+q_{H}\right)\right]}{\left(Q\left(1-R^{a}+(1-R)^{a}\right)+q_{H}\right)^{2}}$
-Since $R\in(0,1)$ this expression is negative. Further, given the assumptions, the indifferent type can be expressed as (for $a\neq 1$)
-$\tilde{i}$ $=\frac{2Q^{2}\left(1-R^{a}+(1-R)^{a}\right)-2(1-q_{H})q_{H}}{2Q\left(1-R^{a}-(1-R)^{a}\right)}$
-$-\frac{Q\left(1-R^{a}+(1-R)^{a}-2q_{H}\left(2-R^{a}+(1-R)^{a}\right)\right)}{2Q\left(1-R^{a}-(1-R)^{a}\right)}$
-where
-$\frac{\partial{\tilde{i}}}{\partial{Q}}=\frac{Q^{2}\left(1-R^{a}+(1-R)^{a}\right)+(1-q_{H})q_{H}}{Q^{2}\left(1-R^{a}-(1-R)^{a}\right)}$
-so that $\tilde{i}$ is increasing in $Q$ when $a>1$ and decreasing in $Q$ when $a<1$.
+Then, using $\beta=F(1/2-R)$, Lemma A.2 directly implies Proposition 5.242424Result (iii) follows because $V(\beta)$ is independent of $\beta$ in this case.
 Proof of Proposition 7
@@
 Proof of Proposition 8
-Using (17) and (18), we obtain
+Observe that
+$\frac{\partial\tilde{i}}{\partial Q}=\frac{Q^{2}\left(R^{a}-(1-R)^{a}-1\right)+(q_{H}-1)q_{H}}{Q^{2}\left(R^{a}+(1-R)^{a}-1\right)}$ (13)
+so that $\tilde{i}$ is monotonically increasing in $Q$ for $a>1$ and monotonically decreasing for $a<1$. Moreover, we have
 $\lim_{Q\rightarrow 0}\Delta_{S}^{B}$ $=\lim_{Q\rightarrow 0}-\frac{Q(1-R^{a}-(1-R)^{a})}{q_{H}+Q(1-R^{a}+(1-R)^{a})}=0,$
@@
 $\lim_{Q\rightarrow 0}\Delta_{O}^{B}$ $=\lim_{Q\rightarrow 0}\frac{2q_{H}+Q(1-R^{a}+(1-R)^{a})}{2(q_{H}+Q(1-R^{a}+(1-R)^{a}))}-q_{H}-Q=1-q_{H}>0.$
-Therefore, $2\Delta_{O}^{B}\geq|\Delta_{S}^{B}|$ so that all types accept a recommendation according to Corollary 1. Thus, Proposition 3 implies
+Therefore, $\tilde{i}$ diverges as $Q\rightarrow 0$. Hence, by continuity and Corollary 1, there exists a cutoff $\bar{Q}$ below which all types accept recommendations. Thus, Proposition 3 implies
 $V(R)$ $=\pi^{B}[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i]]$
@@
 Proof of Proposition 9
+We need to show that $D(R)\leq D(0)\forall R>0$. This is equivalent with
+$\pi_{B}(R)\hat{F}(R)+(1-\pi_{B}(R))(1-\hat{F}(R))\leq q_{H}+q_{1}+q_{2}$ (14)
+To show (14), first note that $q_{H}\geq 1/2$ implies that $\pi_{B}(R)=q_{H}+q_{1}\phi_{1}(R)+q_{2}\phi_{2}(R)\geq 1/2$. Next, $\pi_{B}(R)\geq 1/2$ is equivalent with
+$\pi_{B}(R)\hat{F}(R)+(1-\pi_{B}(R))(1-\hat{F}(R))\leq\pi_{B}(R)$ (15)
+From $D(0)=q_{H}+q_{1}+q_{2}$, we further obtain $\pi_{B}(R)\leq D(0)$. Together, this gives (14).
+A.2 Results from Section 4.3
+Increasing Heterogeneity (Section 4.3.1)
+The following result captures the effects of a mean-preserving spread on the value of the recommendation system.
+Proposition A.1.
+Suppose Assumption 1 holds.
+  1. 1.
+    If $\sigma<1$, then a mean-preserving spread of the type distribution increases the value of the recommendation system when $R>1/2$ and decreases it when $R<1/2$.
+  2. 2.
+    If $\sigma>1$, then a mean-preserving spread of the type distribution decreases the value of the recommendation system when $R>1/2$ and increases it when $R<1/2$.
+Proof.
 As argued in the text, a mean-preserving spread corresponds to an increase in $\beta$ for $R>1/2$, while for $R<1/2$ it corresponds to a decrease in $\beta$.
-From the proof of Lemma 2, we can see that the value of the recommendation system is increasing in $\beta$ for $\sigma<1$ and decreasing in $\beta$ for $\sigma>1$.
+From the proof of Lemma A.2, we can see that the value of the recommendation system is increasing in $\beta$ for $\sigma<1$ and decreasing in $\beta$ for $\sigma>1$.
 Thus, the statement follows.
-Proof of Corollary 3
+To build intuition, suppose $\sigma<1$ and $R>1/2$.
+As shown in Figure 1, when $R>1/2$, only extreme types issue buy recommendations for controversial products.
+[^26] For $R<1/2$, only relatively extreme types issue don’t-buy recommendations for controversial products.
+This follows by reversing the colors in Figure 1 to correspond to $R=1/4$.
+A mean-preserving spread of $F$ (which corresponds to an increase in $\beta$ given $R$), thus leads to more extreme senders and thus a higher likelihood that buy recommendations stem from controversial products.
+This raises the probability of buy recommendations while lowering their objective value.
+The net effect, and whether the system’s value rises or falls, depends on the odds ratio $\sigma$ (see Proposition 5).
+When $\sigma<1$, good products are rarer than bad ones, making the system more valuable as $\beta$ increases.
+The benefit of revealing bad products by means of don’t-buy recommendations outweighs the reduced value of buy recommendations.
+Thus, for high $R$, a more polarized population benefits more from the system, as only bad products result in don’t-buy recommendations.
+Conversely, for low $R$, a less polarized population centered around the middle ensures that only bad products receive don’t-buy recommendations, thereby yielding a higher value.
+Figure 3: Population polarization
+[…]
+Turning to the impact of preference heterogeneity on demand, the analysis is simpler.
+Recall that under Assumption 1, all receiver types accept recommendations.
+Consequently, it follows from Lemma A.2 that demand is increasing in $\beta$.
+Finally, the same distinction depending on whether $R>1/2$ or $R<1/2$ in terms of the impact of a mean-preserving spread applies as above.
+The Prevalence of Controversial Products (Section 4.3.2)
+We prove the following result:
+Corollary A.1.
+Suppose Assumption 1 holds. Then for every $\beta\in(0,1)$, there exist $\underline{\sigma}(\beta)\in(0,1)$ and $\overline{\sigma}(\beta)>1$ such that
+  • (i)
+    the value of the recommendation system is decreasing in $Q$ on $(\underline{\sigma}(\beta),\overline{\sigma}(\beta))$.
+  • (ii)
+    For $\sigma<\underline{\sigma}$ and $\sigma>\underline{\sigma}$, the value of the recommendation system is maximized for some $Q^{*}>0$.
+Figure 4: Illustration of Corollary 4
+(a) Panel A
+$\sigma$$\beta$111133335555777799990.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81.01.01.01.0
+(b) Panel B
+$\sigma$$\beta$111133335555777799990.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81.01.01.01.0
+(c) Panel C
+$\sigma$$\beta$111133335555777799990.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81.01.01.01.0
+Figure 4 illustrates the result.
+In the white area in Panel A, $V(\beta)$ is decreasing in $\beta$; in the shaded area, an interior maximizer exists.
+When good and bad products are similarly likely ($\sigma$ is close to $1$), a higher prevalence of controversial products indeed reduces $V(\beta)$.
+For very high or low values of $\sigma$, however, this need not be true.
+This surprising result reflects the multiplicative structure of $V(\beta)$.
+Under Assumption 1, $V(\beta)$ is the product of the buying probability $\pi^{B}$ and the objective value of the recommendation per buyer, $\Delta_{O}^{B}$.
+The effect of $Q$ on $\pi^{B}$ is negative only if $\sigma$ is sufficiently large (see Panel B).
+Intuitively, as $Q$ increases, good products are less likely, reducing the probability of buy recommendations.
+However, there are now more buy recommendations stemming from controversial products, as these have become more likely.
+$\beta$ and $\sigma$ determine which of these effects dominates.
+When $\beta$ is high, the greater likelihood that controversial products receive buy recommendations increases the total probability of buy recommendations, strengthening the positive effect.
+When $\sigma$ is high, the lower share of good products dominates, strengthening the negative effect.
+Similarly, one can show that the effect of $Q$ on $\Delta_{O}^{B}$ becomes negative only for small $\sigma$ (see Panel C), reflecting the increased probability that a product with a buy recommendation is controversial.
+[^27] There is also a positive effect of higher $Q$, reflecting a change in the value of the outside option.
+The values of $\sigma$ and $\beta$ determine which effect dominates.
+Corollary is an immediate implication of the following statement:
+Lemma A.3.
+Suppose Assumption 1 holds.
+  • (i)
+    If $03\sigma-\beta-\sigma^{2}+\sigma^{2}\beta>$ 0, then the value of the recommendation system is decreasing in $Q$. In particular, this is the case when $\sigma=1$.
+  • (ii)
+    If $03\sigma-\beta-\sigma^{2}+\sigma^{2}\beta<$ 0, then $Q^{*}:=\frac{3\sigma-\beta-\sigma^{2}+\sigma^{2}\beta}{4\sigma-4\beta-4\sigma^{2}+4\sigma^{2}\beta}\in(0,\frac{1}{2})$ maximizes $V(R)$.
+Proof.
 We first take the derivative of $V$ with respect to $Q$ to obtain
@@
 which always holds for $\beta\leq 1$.
-A.2 Distinct Sender and Receiver Distributions
-Proof of Proposition 9
+A.3 Distinct Sender and Receiver Distributions
+We prove the following result:
+Proposition A.2.
+Consider the model with distinct preference distributions of senders and receivers.
+Suppose the sender distribution $F$ satisfies Assumption 1.
+Then, all types accept recommendations.
+Further, $V(R)=\pi^{B}_{F}[\Delta_{O,F}^{B}-\Delta_{S,F}^{B}\mathbb{E}_{G}[i]]$ is
+  1. (i)
+    decreasing in $R$ if $\sigma<\frac{Q-(q_{1}-q_{2})\mathbb{E}_{G}[i]}{Q-(q_{2}-q_{1})\mathbb{E}_{G}[i]}$;
+  2. (ii)
+    increasing in $R$ if $\sigma>\frac{Q-(q_{1}-q_{2})\mathbb{E}_{G}[i]}{Q-(q_{2}-q_{1})\mathbb{E}_{G}[i]}$;
+  3. (iii)
+    constant in $R$ if $\sigma=\frac{Q-(q_{1}-q_{2})\mathbb{E}_{G}[i]}{Q-(q_{2}-q_{1})\mathbb{E}_{G}[i]}$.
+Proof.
 The value of the recommendation system can be written as
@@
 Replacing $q_{H}=((1-2Q)\sigma)/(\sigma+1)$ and rearranging the inequalities $V^{\prime}(\beta)>0$ and $V^{\prime}(\beta)<0$ yields the statement in the proposition.
-A.3 Multiple Recommendation Levels
+A.4 Multiple Recommendation Levels
 Proof of Proposition 10
@@
 Thus, for $q_{L}<q_{H}$ any $\beta_{1}$ and thus any $R_{1}$ is optimal, whereas for $q_{L}>q_{H}$ any $\beta_{2}$ and thus any $R_{2}$ is optimal.
-A.4 Multiple Recommendations
+A.5 Multiple Recommendations
+We now derive some auxiliary results for the case of multiple recommendations, before we prove Proposition 11.
 After $b$ buy recommendations and $d$ don’t-buy recommendations, where $b+d>0$, the posterior reads
@@
 $p_{L}(b,d)$ =$=\begin{cases}\frac{q_{L}}{q_{1}(1-\phi_{1}(R))^{d}+q_{2}(1-\phi_{2}(R))^{d}+q_{L}}&\text{if }b=0\\ 0&\text{if }d=0\\ \end{cases}$
-Lemma 3.
+Lemma A.4.
 Consider the model with multiple recommendations. Suppose Assumption 1 holds and consider any threshold $R\in(0,1)$.
@@
 Optimal Receiver Behavior with Infinite Learning
-Lemma 4.
+Lemma A.5.
 Consider the “infinite learning” recommendation system with interior threshold $R\in(0,1)$. Then,
@@
 Proof.
-First, Lemma 3 implies that infinite learning fully reveals the good and bad products, whereas the posteriors of the controversial product versions are $p_{i}(b,d)=q_{i}/(q_{1}+q_{2})$ for $b,d>0$ Hence, the expected payoff of buying the recommended good with mixed recommendations reads
+First, Lemma A.4 implies that infinite learning fully reveals the good and bad products, whereas the posteriors of the controversial product versions are $p_{i}(b,d)=q_{i}/(q_{1}+q_{2})$ for $b,d>0$ Hence, the expected payoff of buying the recommended good with mixed recommendations reads
 $\frac{q_{1}}{q_{1}+q_{2}}(1/2+i)+\frac{q_{2}}{q_{1}+q_{2}}(1/2-i)=\frac{1}{2}+i\frac{q_{1}-q_{2}}{q_{1}+q_{2}}.$
@@
   2. (ii)
-    If $\lambda>1$, only $i\geq\tilde{i}_{\infty}$ buy the product by Lemma 4.
+    If $\lambda>1$, only $i\geq\tilde{i}_{\infty}$ buy the product by Lemma A.5.
 Hence, no type buys a controversial product in the infinite-learning recommendation system if $\tilde{i}_{\infty}\geq 1/2$, which is equivalent to $\sigma\geq\lambda$.
@@
     The proof is analogous to case (ii) and thus omitted.
+Appendix B Interactive Plots of $V(R)$ and $D(R)$
+In what follows, we describe how to generate interactive plots of $V(R)$ and $D(R)$ under Assumption 2.
+The numerical analysis is based on a self-contained R script using the shiny and ggplot2 packages.
+The script allows users to manipulate model parameters in real time and visualize the corresponding plots through a simple web-based interface.
+Requirements
+  • R (version 4.0 or higher recommended)
+  • RStudio (optional but recommended)
+The script requires the R packages shiny and ggplot2, both of which are automatically checked for and installed if necessary.
+No manual installation is required.
+Beyond that, you only need to download the file recommendations_simulations.R .
+This single R script contains the entire code for both value and demand plots.
+Instructions for Use
+  1. 1.
+    Open RStudio.
+  2. 2.
+    Open the file recommendations_simulations.R .
+  3. 3.
+    Run the script (e.g., select all lines and press Ctrl + Enter or click “Run App”).
+  4. 4.
+    A Shiny application will launch in your Viewer pane or in a web browser.
+  5. 5.
+    Use the sliders to manipulate the parameters $\sigma$, $a$ and $Q$.
+  6. 6.
+    Navigate between the two tabs:
+    • V(R): Value of the recommendation system
+    • D(R): Demand for the recommended product
+Technical Details
+  • The $\sigma$ and $a$ sliders operate on a logarithmic scale centered around $1$ to enable easy exploration.
+  • The $R$ thresholds are plotted from $0.001$ to $0.999$.
+  • The app will dynamically update plots based on the current input values.